### v1

# Time-domain study of coupled collective excitations in quantum materials

Abstract
Quantum materials hold immense promises for future applications due to their intriguing electronic, magnetic, thermal, and mechanical properties that are often traced to a complex interplay among different microscopic degrees of freedom. Important insights of such interactions come from studying the collective excitations of electrons, spins, orbitals, and lattice, whose cooperative motions play a crucial role in determining the novel behavior of these systems and offer us a key tuning knob to modify material properties on-demand through external perturbations. In this regard, ultrafast light-matter interaction has shown great potential in controlling the couplings of collective excitations, and rapid progress in a plethora of time-resolved techniques down to the attosecond regime has significantly advanced our understanding of the coupling mechanism and guided us in manipulating the dynamical property of quantum materials. This review aims to highlight recent experiments on visualizing collective excitations in the time domain, focusing on the coupling mechanism between different collective modes such as phonon-phonon, phonon-magnon, phonon-exciton, magnon-magnon, magnon-exciton, and various polaritons. We introduce how these collective modes are excited by an ultrashort laser pulse and probed by different ultrafast techniques, and we explain how the coupling between collective excitations governs the ensuing nonequilibrium dynamics. We also provide some perspectives on future studies that can lead to further discoveries regarding the emergent properties of quantum materials both in and out of equilibrium.
Contents
I Introduction
Quantum materials have garnered considerable interest due to their remarkable properties associated with, for instance, unconventional superconductivity 1, 2, 3, 4, multiferroicity 5, colossal magnetoresistance 6, and topological quantum states 7. These properties hold great promise for future applications 8 such as topological quantum computing 9 and spintronic devices 10, yet a precise understanding of the underlying mechanism remains elusive in many cases. At the microscopic level, these properties emerge as a result of complex interactions among the charge, spin, orbital, and lattice degrees of freedom. Take superconductivity as an example: the formation of Cooper pairs in the ground state is either aided by electron-phonon interaction in conventional superconductors 11, 12 or other types of electron-boson coupling, if any, in unconventional superconductors. Unraveling the interplay among these degrees of freedom is crucial for advancing our knowledge of the rich phase diagrams of quantum systems. However, the key challenge has been the coexistence of multi-way couplings of a myriad of quasiparticles and excitations, which makes it difficult to isolate a specific type of interaction we are interested in.
To tackle this challenge, recent advances in time-resolved techniques based on ultrafast light-matter interaction have identified time as a new dimension to investigate the couplings among the microscopic degrees of freedom. Leveraging the time resolution of ultrashort pulses down to the attosecond level, experiments based on the pump-probe protocol offer a unique way to disentangle the intertwined couplings at the intrinsic timescales of microscopic interactions 13, 14. Aside from studying the dynamics of individual quasiparticles, a cooperative and wave-like motion of different degrees of freedom — referred to as collective excitations — can be directly observed in the time domain. One of the most notable examples is photoinduced coherent phonons 15, 16, where a macroscopic number of atoms move in perfect synchronization. Such lattice dynamics show up in a variety of time-resolved probes, such as the periodic oscillations of diffraction peak intensity in time-resolved scattering 17, 18, 19, 20, 21, 22 or the oscillation of optical reflectivity and transmissivity in transient optical spectroscopy 23, 24, 25. Oftentimes, more than one collective excitations emerge after a photoexcitation event, and their coupling is governed by the underlying interaction among the constituent degrees of freedom. Therefore, investigating these collective modes and their couplings provides an important perspective to study the microscopic interactions in quantum materials, which in turn offers an opportunity to use light to modify dynamical material properties.
This review focuses on recent experimental progress in understanding the couplings between collective excitations thanks to the development of ultrafast techniques with ever better spatiotemporal resolution and the capability to generate intense laser pulses with highly tailored properties. We begin in Sec. II by explaining why we are interested in studying collective excitations and how time-domain probes of these excitations and their couplings bring out new insights into material properties that are otherwise missing in equilibrium probes. Given that a single probe cannot capture all the degrees of freedom and their collective modes, we subsequently introduce some select ultrafast techniques in Sec. III. In Sec. IV, we delve deeper into the couplings of collective excitations by introducing specific cases such as photon-phonon, phonon-phonon, phonon-magnon, phonon-exciton, among other coupling mechanisms. An outlook for future studies in this field is included in Sec. V.
As this review does not aim to cover the more general area of ultrafast science in condensed matter systems yet an understanding of the broader field is helpful for contextualizing the present article, we hereby refer readers to a more general survey of theoretical advances and experimental techniques in ref. 26 as well as a few other topical reviews. For techniques, see the following review articles on ultrafast optical spectroscopy 27, 28, 29, 30, 31, 25, 32, ultrafast electron diffraction (UED) and microscopy 33, 34, 35, time-resolved X-ray scattering 36, 37, 38, 39, 35, and time- and angle-resolved photoemission spectroscopy (tr-ARPES) 40, 41, 42, 43, 44. Regarding precision engineering of material properties through ultrafast light pulses, we refer readers to excellent reviews on nonequilibrium engineering such as nonlinear phononics 45, 46, 47 and Floquet engineering 48, 49. From the perspective of different material classes, readers can check out reviews on two-dimensional materials 50, magnetic materials 51, high-temperature superconductors 30, 46, complex oxides 52, and topological insulators 49. These lists are by no means exhaustive but provide sufficient conceptual background for appreciating the experiments highlighted in this review.
II Collective excitations viewed from an ultrafast lens
II.1 Collective excitations in and out of equilibrium
Collective excitations, as the name suggests, manifest as cooperative motions of individual degrees of freedom in solids 54. As illustrated in Fig. 1, some common examples include magnons, phonons, excitons, and plasmons. While these collective excitations are intrinsic to a material, others emerge due to hybridization with an external field. For instance, photons become “dressed” when they propagate inside a solid, forming various types of polaritons 55, 56. These distinct collective excitations are characterized by different time scales in their underlying periodic dynamics, where the associated energy scale ranges from eV (GHz frequency) to tens of eV (extreme ultraviolet frequency).
In equilibrium, collective excitations can reveal crucial information about the ground state orders 57 and the associated phase transitions due to a symmetry change. For example, structural or magnetic transitions are often accompanied by soft phonons or magnons, respectively, which signify an instability in the system. In addition, these collective modes are highly responsive to external stimuli, making it possible to achieve on-demand properties of quantum materials as one applies electromagnetic fields, pressure, strain, or ultrafast photoexcitation to modify the collective dynamics 58. While collective excitations are readily detected in an equilibrium setting and have hallmarks in a variety of measurements ranging from transport to scattering experiments, the underlying microscopic motions are incoherent and hence only time-averaged dynamics are obtained.
This situation changed with the development of ultrafast techniques that are capable of visualizing microscopic motion with a time resolution commensurate with the characteristic period of collective excitations. Under the paradigm of pump-probe measurements, coherent dynamics can be excited by an ultrashort laser pulse, offering a general framework for inducing and controlling coherent dynamics in solids. For example, coherent phonons or magnons are often observed following a photoexcitation event, where ionic displacement or spin precession occurs in complete synchrony with the same phase over a macroscopic sample volume. Such coherent motions are in stark contrast to the random movements that account for incoherent phonons and magnons in thermal equilibrium.
II.2 Photoinduced phase transitions
The ability to observe these coherent dynamics yields critical information about light-induced phase transitions. One extensively studied example is transitions that involves a symmetry change under the Ginzburg-Landau framework, such as those found in material systems hosting a charge density wave (CDW). One type of collective excitations of the CDW ground state is the amplitude mode, which has been widely observed following above-gap photoexcitation that transiently suppresses the order parameter amplitude and hence launches a coherent amplitude oscillation 59, 60, 61. Importantly, the photo-driven amplitude mode, in particular its magnitude and frequency, can inform us how the phase transition unfolds after the laser pulse incidence, such as whether the phase change proceeds in a spatially homogeneous fashion or heterogeneity like topological defects is involved 62, 63. As another example, in a prototypical CDW material LaTe3, the CDW order parameter can be suddenly quenched within a time set by the characteristic time of the amplitude mode oscillations 23. With varying photoexcitation densities, the suppression of the CDW order takes the longest time at the threshold photoexcitation density at which the order parameter transiently vanishes, demonstrating dynamical slowing-down in a nonequilibrium transition. Using more than one pump pulses, one can further harness the coherence in photoinduced collective excitations to precisely control the extent of the nonequilibrium transition. This coherent control was recently demonstrated in 1D Si(111)-In nanowires 64 and quasi-2D 1-TaS2 (ref. 65), both of which are archetypical CDW systems.
While most collective modes studied in the context of photoinduced phase transitions are driven by above-gap photoexcitation with ultraviolet to near-infrared light, they can also be induced by an intense mid-infrared or terahertz field. Instead of photo-doping and injecting mobile carriers into the solids, long-wavelength light can directly excite (i) dipole-active modes through the oscillating electric field, (ii) Raman-active modes through two-photon processes 66, or (iii) other strong-field effects 67 to initiate the collective mode that subsequently drives a transition. Novel metastable states and phenomenology can emerge, such as dynamic multiferroicity 68, 69, 70, 71, a change in electronic topology 72, 73, and magnetism induced by light-driven chiral phonons 74, 75, 76 or as a result of strong spin-lattice coupling 77. We again take CDW materials as an example. In metallic 3-Ta1+xSe2, when its CDW amplitude mode is excited by an intense terahertz pulse, an energy gap appears in the optical conductivity spectrum and executes a dynamical evolution in sync with the amplitude mode oscillation, suggesting a transition into a transient insulating state that has no equilibrium counterpart 78.
II.3 Dressed states
Aside from the role of collective modes during the course of a nonequilibrium phase transition, they can also help induce novel dynamical states of matter via coherent light-matter interactions. For instance, whereas Floquet-Bloch states have been conventionally realized by hybridization between electronic states and an external periodic light field 79, 80, 81, 82, 83, 84, 85, there are a number of theoretical proposals for the Floquet states achieved by replacing photons with phonons 86, 87 or excitons 88. One signature of Floquet-Bloch states is the appearance of electronic band replicas detected in tr-ARPES experiments, where replica bands follow specific intensity relations and polarization selection rules of the exciting photons. Importantly, these bands are spectrally spaced apart by , where is the angular frequency of the bosonic field, be it a photon, phonon, or exciton. In a recent photoemission experiments using a momentum microscope that was specifically designed to study very small samples, exciton-induced Floquet effects were reported in monolayer WS2 in the presence of a high density of resonantly-driven photoexcited excitons 89, which also bear the hallmark of an excitonic insulator phase.
II.4 Phase resolution of coherent collective excitations
When coherent dynamics are launched in solids upon photoexcitation, the motion of collective excitations manifests as wave-like oscillations. The waveform of these oscillations can be directly visualized via an ultrafast probe, an example of which is shown in Fig. 1b that depicts an oscillating Kerr rotation signal caused by a coherent magnon. Therefore, the phase of the collective excitation oscillations can be determined (labeled as in Fig. 1b, bottom panel), which is important information that equilibrium experiments cannot measure directly.
Phase-sensitivity can be used to isolate a single collective mode that affects multiple degrees of freedom, providing a way to measure the coupling strength between them 19, 22. More specifically, by selectively inducing a particular oscillatory response using a tailored pump, and by leveraging phase-locked coherent features from different observables for different degrees of freedom, one can obtain the coupling strength directly, independent of the theoretical model. For example, following the excitation of a coherent phonon, tr-ARPES can probe spectral shifts in the electronic band energy while time-resolved X-ray diffraction (tr-XRD) can quantify the shifts in each atomic position for the phonon of interest. By combining these two measurements, electron-phonon coupling of a specific electronic band to a specific phonon mode can be calculated. This procedure has been applied to an iron-based superconductor FeSe (ref. 19) and three-dimensional topological insulators Bi2Se3 and Bi2Te3 (ref. 22). The extracted electron-phonon coupling strength in Bi2Se3 and Bi2Te3 generally conforms to the expectation from density functional theory (DFT) computations while that in FeSe shows a significant deviation from DFT, which underestimates the strength. Instead, after electron-electron correlation effects were included in a self-consistent DFT-dynamical mean field theory, a quantitative agreement emerged, highlighting an enhancement of the coupling strength due to strong electron correlations. Besides electron-phonon coupling, this multi-probe experimental protocol has also been deployed to measure the dynamical magnetoelectric coupling in multiferroic materials. For instance, in NiI2, a giant coupling strength was obtained 90 where a single electromagnon mode induces periodic modulations in both the dipolar and magnetic orders in NiI2, revealing an interesting /2 phase shift and lock-in behavior between electric polarization and magnetization oscillations that are at terahertz frequencies.
Even with a single experimental probe, a changing oscillation phase with respect to some tuning parameter can be used to reveal material properties 91, 92, 93 and to identify microscopic processes underlying the collective excitation 94. For instance, far above the Curie temperature, the optical phonon mode in a layered ferromagnetic insulator CrSiTe3 can be excited due to the conventional displacive mechanism 15, but the coherent phonon phase experiences a shift when the material cools below a temperature scale where in-plane ferromagnetic correlations start to develop 92. Such a phase shift was only observed when the excitation photon energy is sufficiently large to induce a charge transfer from the Te2- to Cr3+ ions, a process that also increases the superexchange energy. The phase shift was interpreted as evidence for a coherent phonon mechanism due to superexchange enhancement via magneto-elastic coupling, which only applies in the presence of short-range ferromagnetic correlation in this material. Notably, the phase shift of an mode across the Néel temperature also appears in an antiferromagnetic insulator FePS3, providing evidence for a selective coupling between the magnetic order to the trigonal distortions in this layered system 93. In another example of a coherent acoustic phonon in a freestanding film of La2/3Ca1/3MnO3, three different observables in time-resolved diffraction (Bragg peak intensity, width, and position) were observed to synchronously oscillate with a particular phase relation — either in-phase or out-of-phase — after photoexcitation 94. By theoretically modeling the phase relation across the distinct observables from this coherent acoustic mode, Su and coworkers interpreted it as features of ultrafast wrinkle formation that results from delamination between the freestanding film and the substrate at its boundary, offering unique insights into the oft-neglected film–substrate interaction that governs the properties of freestanding structures.
III Time-domain probes of collective excitations
To explore collective excitations in the time domain and to study their couplings, a variety of advanced ultrafast tools have been developed, each with distinct advantages. Here, we briefly introduce a few selected techniques that serve as the workhorse in the experiments we reviewed in this article. A schematic for each category of techniques is depicted in Fig. 1b–e. For a typical table-top time-resolved probe introduced below, a femtosecond laser pulse is split into two beams, one of which is used to generate a tailored pump pulse to drive the nonequilibrium dynamics while the other is used to produce a probe pulse, which is either a photon or electron pulse. The frequency of the pump pulse can range from sub-terahertz up to ultraviolet and its polarization can be tuned anywhere between linear and circular, depending on the target excitation.
III.1 Time-resolved optical spectroscopy and polarimetry
The most commonly used ultrafast pump-probe technique is time-resolved optical spectroscopy and polarimetry, depicted in Fig. 1b. In a typical experiment, white light or monochromatic pulses are used as probes, and experimental observables can be transmission, reflectivity, photoluminescence, birefringence, and dichroism, all of which can be measured with varying energies and polarization states of the probe beam. These measurements are not confined to the fundamental frequency of the probe but also apply to its second or higher-order harmonics, which can be more sensitive to order parameter symmetry hidden in the linear order measurement. Based on these observables, different variants of the techniques have been developed, such as transient optical transmittance or reflectance spectroscopy 28, 29, 30, 95, 96, 31, 32, 25, time-resolved magneto-optical Kerr effect 97, time-resolved second harmonic generation (SHG) 98, 99, and time-resolved Raman scattering 100, 101, 27, 102, 103, 104. These measurements allow us to extract the complex optical constants of a material 32, which make it possible to infer collective excitations associated with carriers 105, 106, 107, 108, spins 51, 24, and polarizations 99 in quantum materials.
III.2 Ultrafast electron and X-ray scattering
In contrast to optical spectroscopy in the infrared to ultraviolet regime, high energy electrons and X-rays have much shorter wavelengths that are comparable to the length scale of lattice parameters of common crystals, making time-resolved electron and X-ray diffraction suitable for tracking phonon dynamics 36, 109, 34, 35 and charge or orbital order evolutions that are modulated by the lattice 110, 111. In a typical diffraction experiment, a tailored pump is used to launch the nonequilibrium dynamics of nuclear ions. It should be noted that the pump pulse duration should always be shorter than the period of the collective excitations involved. The ensuing periodic motion of ions modulates the Bragg condition and can lead to a coherent oscillation of the intensity or position of Bragg peaks — the former is commonly associated with an optical phonon while the latter with an acoustic mode 112.
Besides coherent phonons, incoherent phonons can also be populated following an excitation event, where incoherent means there is no net displacement of ionic positions, namely, the expectation value of phonon amplitude at wave vector is zero at all time , . Inelastic scattering of a specific phonon branch can lead to a particular distribution of diffuse scattering 113, 114, 115, 116, 117, 53 depending on how this phonon modulates the structure factor. The bottom panel of Fig. 1c shows a simulated nonequilibrium electron diffuse scattering pattern resulting from a nonthermal population of phonons induced by light 53.
It is worth noting that the absence of coherent phonons does not imply the absence of time-oscillatory features in a diffraction measurement. Without a net displacement of ions, the pump light can couple to the lattice to the second order in the phonon amplitude, where pairs of phonons at are generated to conserve the total momentum, leading to oscillations of the variance in the atomic displacement at twice the frequency of the phonon mode itself 118. This second-order effect has been demonstrated as intensity oscillations in either Bragg peaks 119 or diffuse scattering 120, and is also known as squeezed phonons 121, 122, 120, 123.
Another rapidly developing area that rides on recent progress in the X-ray free-electron laser technology is time-resolved resonant inelastic X-ray scattering (RIXS), which not only reveals phonon dynamics but can also access charge, orbital, and spin excitations as well as high-order, multi-particle correlation functions that describe coupled spin/charge/orbital excitations 124, 38, 125, 39. With ongoing effort on entering the sub-10-meV energy resolution in RIXS 126, this emerging technique provides a unique window into the nonequilibrium states in correlated materials through dynamical signatures across time, momentum, and energy axes.
III.3 Time-resolved photoemission
Time- and angle-resolved photoemission is a complementary method to ultrafast diffraction, enabling us to directly measure the evolution of electronic structure 40, 41, 42, 43, 44 (see Fig. 1d). In an equilibrium ARPES experiment, extreme ultraviolet to soft X-ray photons are illuminated onto a material to eject photoelectrons, whose energy and momentum inform us about the single-particle spectral function of the corresponding electrons inside a material 127, 128. By introducing a pump pulse, tr-ARPES yields further insights into nonequilibrium carrier dynamics and unoccupied states, reveals the electronic origin of ultrafast phase transitions 129, 130, and unambiguously demonstrates how band structure can be engineered by light 79, 80, 48, 49, 82, 81, 89.
The advent of momentum microscope in tr-ARPES experiments has also recently made it feasible to investigate micrometer-sized samples that were challenging to access before. For example, tr-ARPES experiments 131, 132, 89 have demonstrated the ability to probe dark excitons and moiré excitons in monolayer and heterobilayer transition metal dichalcogenides, respectively, whereas they were hard to probe by traditional photon-absorption or photon-emission spectroscopy. As illustrated in the bottom panel of Fig. 1d, only the occupied state at the valence band top can be visualized before the optical pump. With resonant photoexcitation at an exciton energy, photoemission signals from excitons can be observed upon the arrival of the pump pulses 131. The exciton photoemission signals feature an energy-momentum distribution centered below the conduction band bottom with a gap of binding energy (green dispersion line), conforming to theoretical expectations 133.
III.4 Time- and space-resolved probes
Time-resolved spectroscopic and diffraction methods introduced thus far can only provide spatially-averaged information inside the probed region, whereas in general ultrafast dynamics may vary in different regions due to different local environments. Therefore, techniques with both spatial and temporal resolutions play an important role in studying inhomogeneous effects and near-field distribution during ultrafast light-matter interaction. Techniques such as ultrafast electron microscopy 134, 135, photoinduced near-field electron microscopy (PINEM) 136, 137, coherent correlated X-ray imaging 138, 139, 140, and lightwave-driven terahertz scanning tunneling microscopy 141, 142 have all shown a great potential in studying heterogeneity and spatially-varying dynamics present in quantum materials.
As an example to illustrate the relation to collective excitations discussed in this article, ultrafast PINEM exhibits exceptional capabilities in studying the evolution of light-induced plasmonic field distribution 143 (see Fig. 1e), and we will return to specific examples in Sec. IV.5. PINEM is based on ultrafast electron microscopy, which typically uses elastically scattered electrons for imaging. In the PINEM mode, inelastically scattered electrons are used instead, where electrons can absorb and emit spatiotemporally overlapped photons and produce discrete peaks with a constant energy interval in the electron energy loss spectrum. Using the energy-filtered electrons for imaging, one can obtain information about dressed photons such as the near-field distribution of surface plasmon-polaritons 144 and propagation of phonon-polaritons 145. By focusing pump and probe pulses at distinct points and spatially scanning the probe electron pulse, the near-field distributions with space, time, and energy resolution can also be obtained 146.
The techniques introduced in this section are by no means exhaustive, and here we refer readers to other excellent topical reviews on different classes of ultrafast probes 28, 40, 29, 30, 41, 42, 36, 37, 38, 39, 31, 33, 43, 34, 25, 32, 35, 44.
IV Couplings between collective excitations at non-equilibrium
IV.1 Dressing photons with phonons: phonon-polariton
When solids are exposed to external electromagnetic waves, many dressed states or hybridized quasiparticles emerge. Polaritons are those effects that arise from a strong light-matter interaction between photons and dipole-active excitations. Polaritons can be of various types 55, 56 depending on which excitations they hybridize with. For instance, light can hybridize with magnons, phonons, excitons, and plasmons, forming magnon-polaritons, phonon-polaritons, exciton-polaritons, and plasmon-polaritons (Fig. 1). One hallmark feature of these interactions is the appearance of an avoided crossing between the dispersion of two quasiparticles 58, as illustrated in Fig. 2a. The size of the gap is a measure of the interaction strength, where a larger gap suggests a stronger interaction.
Here we focus on recent ultrafast studies of phonon-polaritons, which arise from the interaction between lattice vibration and electromagnetic waves. When light passes through a crystal, it can drive a transverse optical phonon mode if the frequency of the electromagnetic wave resonates with the phonon, which leads to dipole oscillations and re-radiates the electromagnetic wave. The re-radiated waves can further excite the dipolar vibrations, resulting in the propagation of hybridized light and phonon, namely, phonon-polaritons. As the hybridized quasiparticles inherit the attributes from both light and phonons, their frequencies lie in the frequency range between gigahertz and tens of terahertz, showing great potential for applications in THz spectroscopy, THz imaging, and high-bandwidth signal processing 149.
Many early ultrafast experiments of phonon-polaritons focused on the study of ferroelectrics 150, in which photoexcited coherent phonon-polaritons were used as a probe for low-frequency anharmonic phonons that play a crucial role in driving structural and ferroelectric phase transition near the critical point. Raman spectroscopy was first used to measure the frequency of phonon-polaritons 151 but lacked phase information of their wavepacket. With the development of femtosecond laser, ultrafast optical spectroscopy with spatially distinct pump and probe techniques has enabled us to directly measure the phonon-polaritons waveform in the time domain. In one experimental scheme, broadband polar phonon-polariton modes are produced by illuminating a femtosecond laser pulse onto the surface of a ferroelectric material, where the generation mechanism is based on impulsive stimulated Raman scattering (ISRS) through mixing different frequency components contained within the pulse bandwidth 152, 153. However, this method does not generate phonon-polariton mode with a well-defined frequency. To solve this problem, a four-wave mixing method has been developed 154, 147, 155, 156. In this configuration, two identical, spatiotemporally overlapped femtosecond pulses are used to launch the phonon-polaritons in a nonlinear crystal. The two pump pulses form a transient grating due to interference, and they selectively excite phonon-polaritons with a well-defined scattering wavevector that can be tuned by varying the incidence angle between the pump pulses. As the photoexcited phonon-polariton field modulates the index of refraction that in turn affects the scattering of a separate probe pulse, the periodic probe signal can reveal the presence of phonon-polaritons in a phase-resolved manner. Figure 2b shows the measured dispersion relation of phonon-polaritons in LiNbO3 using this four-wave mixing scheme 147, where five branches and four avoided crossings were observed, showing strong couplings between the photons and phonons in this material.
Experiments using the four-wave mixing methodology have some limitations: ISRS can only excite Raman-active modes 153, while to the lowest order only IR-active phonon modes can couple with electromagnetic waves to form phonon-polaritons. Therefore, the scheme is only feasible for noncentrosymmetric materials where IR phonons are also Raman-active, such as in LiNbO3 and LiTaO3. In addition, these experiments need careful adjustment of the geometry of the two pump beams in order to excite phonon-polaritons with a selected wavevector. An alternative method that overcomes these limitations is time-domain THz spectroscopy 157, where phonon-polaritons are excited by a single beam of THz radiation instead of two beams of visible or near-infrared light pulses. However, in a typical time-domain THz experiment that measures phonon-polariton dispersion, a reference crystal is required for accurate electro-optic sampling to retrieve the phase information across the spectrum of the THz pulse, hence complicating the data analysis and interpretation.
Recently, Luo and coworkers developed a time-of-flight detection method that overcame these challenges and was applied to probing phonon-polaritons in van der Waals magnetic materials 148, including NiI2 and MnPS3. In this scheme, a terahertz pulse centered around 3 THz was used to excite phonon-polaritons while the SHG signal of a 800-nm laser pulse was used as a probe to capture phonon-polaritons before or after they are reflected from the back surface of the sample. More specifically, when the THz pump pulse arrived at the sample surface, multiple oscillatory responses due to phonons and magnons were observed (Fig. 2c, window A). After the initial decay of the oscillations, there was a sudden amplitude revival at around 6 ps, suggesting a reflection of the propagating phonon-polariton wavepacket from the sample back surface (Fig. 2c, insets). After this revival, the signal frequency was also observed to increase towards a longer time delay (Fig. 2c, windows B and C). A clearer view of this phenomenon can be seen in the wavelet transform spectrum in Fig. 2d, which shows the time-dependent signal frequencies. Around time zero, as defined by the THz pulse incidence, the excitation spectrum and the second harmonic frequencies of the THz pump can be seen, including the photoexcited and phonon modes. After 6 ps, a chirped mode emerges, whose frequency increases from 2 to 4 THz over time and it corresponds to the phonon-polariton wavepacket. Given the known sample thickness, the time-dependent frequency spectrum in Fig. 2d can yield the group velocity at different frequencies, which in turn helps one reconstruct the energy-momentum dispersion of the phonon-polariton. Furthermore, the time-dependent attenuation of the phonon-polariton during its propagation allows one to compute the imaginary part of the dispersion, whose divergence near the phonon resonance conforms to the expectation that light quickly dissipates its energy through the phonon mode.
While the aforementioned experiments can detect the presence and dispersion of phonon-polariton wavepackets in the probed volume, they do not reveal the spatial structure of these hybridized quasiparticles. To image the spatiotemporal propagation of phonon-polaritons, ultrafast optical or electron microscopy was used. Nelson and colleagues first demonstrated direct imaging of optical phonon-polariton wavepackets in LiTaO3 using ultrafast optical microscopy that leveraged four-wave mixing 155, 156. In one of the experiments, an anisotropic spatial profile of the pump pulse was applied (see Fig. 3a), which excited phonon-polaritons that propagated mainly along the horizontal direction of the sample 155. The spot size of the probe pulse was significantly larger than the excitation region, so spatial information of the propagation dynamics was readily acquired. The left panel of Fig. 3b shows a typical snapshot of the wavepacket distribution at 1335 fs after the pump incidence. After averaging over the vertical direction in each image captured at different time delays, the spatiotemporal propagation of phonon-polaritons in LaTiO3 was obtained in the right panel of Fig. 3b. Under this experimental framework, the phonon-polariton response can be coherently controlled via spatiotemporal shaping of the optical pump pulses 149. For a more comprehensive survey on the coherent control of phonon-polaritons and THz polaritonics, we refer readers to ref. 149.
Apart from optical probes, electrons have also been successfully deployed to image phonon-polariton propagation dynamics in van der Waals materials 145, leveraging the spatial and spectral information of PINEM. As shown in Fig. 3c, an optical pump pulse first illuminates the edge of a hexagonal boron nitride (hBN) and launches multi-branch phonon-polariton wavepackets. The probing free electrons interact with the propagating phonon-polaritons, broadening the electron energy loss profile due to inelastic scattering (Fig. 3d). In this experiment by Kurman et al., it should be noted that the energy of each phonon-polariton is smaller than the spectral resolution of the instrument. Hence, instead of showing discrete peaks drawn in the schematic in Fig. 1e, the electron energy loss spectrum profile is broadened. Figure 3e shows the averaged energy-filtered electron counts along the edge direction (blue arrow in the inset of Fig. 3e) as a function of distance from the edge (red arrow in the inset of Fig. 3e), indicating the field profile of phonon-polariton along the propagation direction. By studying the distance-time relation of the traveling wavepackets (Fig. 3f), one can observe both acceleration (purple arrow) and deceleration (red arrow) during the propagation of the photoexcited phonon-polaritons. The observed acceleration and deceleration result from the excitation of different spectral components at different times during the pump-hBN interaction. Interestingly, Kurman et al. also showed that the accelerating and decelerating propagation can be modified by changing the exciting photon wavelength and sample thickness 145.
Note that the two featured examples of spatiotemporal imaging differ in the launching mechanism of the phonon-polaritons. While ultrafast optical microscopy used two near-infrared pulses to form a transient grating to excite phonon-polaritons based on the ISRS mechanism 155, the PINEM experiment directly used a mid-IR pulse to coherently excite them 145. Hence, the former methodology was deployed to noncentrosymmetric materials while the latter was used to directly excite dipole-active IR modes to launch the phonon-polaritons. In both cases, the additional spatial information has deepened our understanding of how phonon-polaritons propagate in materials, which will be especially helpful when one considers spatially heterogeneous samples in a real device setting 149, 158, 159.
IV.2 Phonon-phonon coupling
The lattice structure often dictates the electronic, optical, and mechanical properties of quantum materials. The collective excitations of atomic vibrations, or phonons, hence play an important role in determining these physical attributes as well. Under equilibrium conditions, phonons and their couplings can be characterized by techniques such as IR spectroscopy 160, Raman spectroscopy 161, and inelastic neutron, X-ray, or electron scatterings 162. Though these static methods are powerful, a lack of precise control over a particular phonon mode impedes us from revealing how individual phonon modes couple to other quasiparticles in solids. In this regard, ultrafast techniques allow us to tackle this challenge by extending the measurement into the time domain. For example, advances in generating intense mid-IR and THz pulses with either broad band or tunable narrow-band frequencies bring the paradigm of nonlinear phononics 163, 47. By selectively exciting particular phonon modes using tailored low-photon-energy pulses, we can study phonon-phonon couplings and engineer crystal structure in a nonequilibrium state.
The first experimental demonstration of nonlinear phononics was reported in 2011 by Först and coworkers 164. After the IR mode in a bulk single crystal of La0.7Sr0.3MnO3 was excited by 14.3-m light pulses, the reflectivity exhibited an oscillation at 1.2 THz, as shown by the dark curve in Fig. 5a. The oscillation was attributed to the Raman mode that is associated with oxygen octahedral rotation in this perovskite. Moreover, the amplitude of the mode showed a strong pump wavelength dependence and reached a maximum when it was resonant with the mode (see Fig. 5b). This intriguing phenomenon can be described by an anharmonic coupling between the IR-active mode and the Raman-active mode, as illustrated in Fig. 4. During this process, a mid-IR light pulse resonantly couples to the dipole of the IR-active phonon and sets off coherent oscillations of atoms around their equilibrium positions (Fig. 4a). The anharmonic coupling term () applies a directional force (Fig. 4b) on the Raman mode (Fig. 4c), thus driving its oscillations. For a more detailed description of this mechanism such as symmetry considerations during anharmonic interactions, we refer readers to refs. 165, 47. In contrast to the resonant excitation condition, 5.8-THz oscillations were observed when La0.7Sr0.3MnO3 was photoexcited by a 1.5-m pulse (Fig. 5a, light curve). This phonon mode was attributed to the mode, which was induced by the more conventional mechanism involving a displacive excitation of coherent phonons 15.
This work by Först et al. inspired ideas to engineer crystals using resonant mid-IR or terahertz excitations. This nonlinear phononic mechanism has since been utilized to manipulate properties in many perovskite systems, such as controlling an anisotropic strain wave in LaAlO3 166, as well as changing ferroelectric polarizations in systems such as LiNbO3 (refs. 167, 168), SrTiO3 (refs. 169, 170, 171), and BiFeO3 (ref. 172). Besides the effect on the lattice and hence the coupled polar orders, nonlinear phononics have also been shown to influence electronic and magnetic properties 47. A particularly intriguing case concerns the apparent resemblance to a transient superconducting state when resonant excitation of IR-active phonons of specific wavelengths are driven to a large amplitude 173, 174, 175, 176, 177, 178. However, other works have suggested that certain superconducting-like behavior may be a manifestation of the excitation of quasiparticles with a low scattering rate 179, 180, and that care needs to be taken in processing the transient optical data for drawing precise conclusions about the nonequilibrium state 181, 182, 183. This potential of light-induced superconductivity remains an intensely studied area with many interesting, open questions to be answered.
Compared with ultrafast optical spectroscopy, time-resolved X-ray diffraction can directly quantify the changes in a crystal structure, thus offering a powerful alternative to study phonon-phonon couplings 36, 35. Here, we highlight several experimental works on SrTiO3, a quantum paraelectric material where the competition between ferroelectric and antiferrodistortive instabilities prevents it from entering a ferroelectric phase at low temperature. Using THz pump pulses, the direct control of its ferroelectric soft phonon mode has been intensely studied and several nonlinear phononic effects have been observed 169, 170, 171. For example, phonon up-conversion in SrTiO3 was demonstrated in a THz pump, X-ray diffraction probe experiment 163. Kozina et al. used an intense THz pump pulse with a spectral range of 0.2–2.5 THz (red region in Fig. 5c), which overlapped with the IR-active TO1 soft mode in SrTiO3 at 100 K (the yellow peak). Interestingly, the oscillatory Bragg peak intensity observed in X-ray diffraction not only showed the frequency of the driving THz field but also exhibited several high frequencies that correspond to the TO2 and TO3 modes (turquoise and purple peaks in Fig. 5c). Upon a temperature increase where the hardened TO1 mode falls outside the THz excitation spectrum, no signatures of the TO2 and TO3 modes are seen, providing compelling evidence for phonon frequency up-conversion from the soft TO1 mode to these high-frequency modes.
Beyond the proof-of-principle demonstration by Kozina et al., nonlinear phononic effects have also been used to induce ferroelectricity in unstrained SrTiO3 under ambient pressure. In 2019, Li and coworkers reported that SrTiO3 can be driven to a transient ferroelectric phase via intense THz pumping as evidenced by the second harmonic signal of a 800-nm probe pulse 169, although later experiments performed on a cousin compound KTaO3 suggested that the THz-induced transient second-harmonic signal may come from defect-induced local polar structures without any long-range ferroelectric ordering 185. A contemporary work in 2019 reported a metastable ferroelectric phase of SrTiO3, which was induced by a mid-IR pump pulse instead of a THz pulse 170. In this experiment, Nova and coworkers resonantly drove the high-frequency IR-active mode in SrTiO3 at 4 K, and detected a metastable noncentrosymmetric phase via second-harmonic spectroscopy. The experiment was interpreted via anharmonic phonon-phonon interactions between the pulse-induced IR-active mode and acoustic strains, the latter of which suppressed the antiferrodistortive distortion and hence stabilized the ferroelectric phase. This mechanism was further clarified by a subsequent work using time-resolved X-ray diffusing scattering 171. The antiferrodistortive fluctuations can lead to a significant diffuse scattering intensity at the point in the Brillouin zone, as shown in the equilibrium X-ray detector image in Fig. 5d. At 135 K, upon photoexcitation with a 17-THz pulse that resonantly drives the IR-active mode, X-ray scattering intensity at shows a sudden increase followed by a decay (Fig. 5e). The sudden intensity increase is a result of the enhancement of antiferrodistortive fluctuations induced by their anharmonic coupling with the resonantly excited high-frequency IR-active mode. The same IR-active mode can also couple to low-frequency acoustic modes, which induce a strain that dominates at long time delays ( ps) and suppresses the antiferrodistortive fluctuations, resulting in the decrease of X-ray intensity at the point. Though the X-ray experiment was performed at 135 K that is above the antiferrodistortive transition, the authors expected the same mechanism would remain valid to explain the metastable ferroelectric property of SrTiO3 at 4 K, where a suppressed antiferro-distortion at long times leads to the emergence of a ferroelectric phase.
Aside from the standard pump-probe scheme employed in transient optical spectroscopy and X-ray scattering, 2D THz spectroscopy offers another route to determine the interaction pathway between different quasiparticles, showing particular relevance in revealing the phonon-phonon coupling process 29. In typical 2D THz spectroscopy setups, two intense THz pump pulses are used to induce nonlinear responses with a time delay between the two pulses, as shown in Fig. 5f. A third optical or near-infrared probe pulse with time delay is then applied to record the optical response of the sample. By performing a 2D Fourier transform on the time-domain data with varying and , spectra as a function of THz excitation frequency and detection frequency can be generated, yielding insights into the interaction process between different modes. As a recent example, 2D THz spectroscopy was employed to reveal the energy-flow pathway in a topological antiferromagnet MnBi2Te4 (ref. 184). The normalized 2D Fourier transform spectrum of the nonlinear signal is shown in Fig. 5g, while Fig. 5h shows three simulated spectra based on the different coupling pathways of exciting the Raman-active phonon: two-photon absorption (“photonic”), nonlinear phononics (“phononic”), and a combination of the two (“photophononic”) . Careful comparison with the experimental observation in Fig. 5g pinpoints a photophononic pathway as the most probable scenario, where the THz electric field first excites the IR-active mode and then interacts with this mode to further excite the Raman-active mode, as summarized in Fig. 5i.
IV.3 Phonon-magnon coupling
Novel properties in quantum materials often arise from the coupling between the spin and lattice degrees of freedom, for example, the Einstein-de Haas effect 190, the Barnett effect 191, 192, magnetostriction 193, and piezomagnetism 194, 195. In certain contexts, their strong coupling implies a unique spin configuration dictated by the crystalline lattice, and one prototypical example is the realization of quantum spin liquid predicted in a frustrated geometry such as in a triangular or kagome lattice 196.
While effects due to strong spin-lattice coupling have been extensively investigated in thermal equilibrium 197, ultrafast techniques allow us to probe the exact energy-flow pathway and unveils how angular momentum is transferred between the electron spin and the crystalline lattice. For instance, ferromagnetic systems are known to lose their net magnetization well within a picosecond when they are subjected to an intense ultrafast laser pulse 198, but how the conservation of total angular momentum is maintained at this timescale has been a long-standing puzzle. Recent works using time-resolved X-ray and electron diffraction offered some clues to this question 199, 200, where circularly-polarized high-frequency phonons are first excited to absorb the angular momentum, which are then transferred to low-frequency shear acoustic modes and a macroscopic rotation of the sample. Interestingly, acoustic shearing is similarly reported during ultrafast demagnetization of a van der Waals antiferromagnet whose net spin angular momentum is zero 201, 21, 202, and a detailed description of the angular momentum flow in the antiferromagnetic case remains an open question.
Strong spin-lattice couplings can manifest in the form of magnon-phonon coupling, where a hybridized magnon-phonon mode, magnon polarons, can readily be detected in equilibrium. Similar to phonon-polaritons discussed in Sec. IV.1, one characteristic feature of magnon polarons is the avoided crossing of energy levels of the constituent magnon and phonon modes. An example of avoided crossings is shown in Fig. 6a for a layered antiferromagnet FePSe3, measured by magneto-infrared transmission spectroscopy 186. In the detected spectral window of FePSe3, while the constituent phonons are individually chiral with opposite chiralities, their near degeneracy leads to no net chirality. However, the coupling to chiral magnons lifts such degeneracy, resulting in chiral magnon polarons that were demonstrated by magento-Raman spectroscopy. The hybridization of magnon and phonon can also be directly measured in the time domain using ultrafast techniques. For instance, in a cousin van der Waals antiferromagnet FePS3, two Raman-active optical phonons at 3.2 and 4.8 THz can be excited by a near-infrared pump pulse, evidenced by the coherent oscillations in the polarization rotation signal of the probe pulse (Fig. 6b) 187. As the applied magnetic field increases, the 3.2 THz phonon splits into two phonon-magnon modes due to phonon-magnon coupling (Fig. 6c), which is reminiscent of the hybridization-induced peak splittings detected in equilibrium 203, 204, 205.
In the study of phonon-magnon coupling, what sets nonequilibrium experiments apart from their equilibrium counterpart is the coherence of the dynamical motion of both atoms and spins, leading to a number of interesting phenomena. For example, circularly polarized phonons can be coherently excited, leading to an effective magnetic field that can in turn excite coherent spin precession. In an early demonstration of this effect, Nova and coworkers used a 20-THz mid-infrared pump to resonantly drive the in-plane and phonons in an antiferromagnetic insulator ErFeO3 (ref. 188), whose phonon eigenvectors are sketched in Fig. 7a. Due to the orthorhombic distortion in ErFeO3, these two phonons exhibit different eigenfrequencies and oscillator strengths. When the modes are excited simultaneously, the two modes begin with a nonzero relative phase and evolve at different rates, resulting in an elliptical motion of oxygen atoms, as illustrated in the bottom panel of Fig. 7a. The elliptically polarized phononic field mimics the application of a magnetic field and results in a spin precession, leading to the excitation of a coherent magnon at 0.75 THz frequency, which is readily detected in the Fourier spectrum of the pump-induced Faraday rotation signal (Fig. 7b). The elliptically polarized phonons are otherwise known as chiral phonons, which have attracted both theoretical 207, 208 and experimental 209, 76, 70, 69 interests because of their ability to create an effective magnetic field that paves the way for “chiral phonomagnetism” in magnetization-based devices.
Coherent magnons can also be excited by phonon-induced lattice distortions that modify the exchange interaction between neighboring atoms, a mechanism recently demonstrated in a closely related antiferromagnet DyFeO3 by Afanasiev and colleagues 189. In this material, the highest frequency IR-active transverse optical phonon mode () is associated with a periodic stretching of the Fe-O bond. Using a tailored terahertz pump pulse, this phonon can be resonantly driven to a large amplitude and couple to a finite lattice distortion along the coordinate of a Raman-active mode via the nonlinear phononics pathway discussed in Sec. IV.2 (see Fig. 4). This lattice distortion involves an antipolar motion of the Dy3+ ions, resulting in a rectification of the Fe-Dy exchange integral that can reach approximately 1–2 eV per unit cell according to density functional theory calculations, thus setting off a coherent magnon oscillation. Depending on whether DyFeO3 is in the antiferromagnetic state ( K) or the weakly ferromagnetic state ( K), distinct magnon modes are induced by the resonant driving of the same optical phonon (Fig. 7c,d). Importantly, the magnon mode frequencies under the resonant driving condition (orange peak in the insets) are clearly distinct from those under off-resonant pumping (black peak in the insets), the latter of which are excited via impulsive stimulated Raman scattering 210. Leveraging the modification of exchange interaction under a resonant pumping condition, Afanasiev and colleagues further showed that with a very large fluence of the resonant pump, an ultrafast magnetic transition from an antiferromagnetic state to a weakly ferromagnetic state can be induced coherently, a transition that proceeded much faster compared to an above-bandgap excitation scheme.
While the preceding examples illustrated how coherent magnons can be induced by resonant excitation of coherent phonons, the reverse pathway from magnons to phonons has also been demonstrated. In an experiment conducted on antiferromagnetic CoF2, Mashkovich et al. used an intense, nearly single-cycle terahertz pulse to pump the crystal below its Néel temperature, where the magnetic field of the pulse resonantly excited a magnon mode (Fig. 8a) 206. After approximately 10 ps, the amplitude of the magnon decreases, followed by a persistent phonon oscillation over the next tens of picoseconds. By contrast, the magnon and phonon oscillations both disappeared above the Néel temperature (Fig. 8b), strongly suggesting that the phonons are induced by the magnons and not vice versa. The researchers further performed 2D terahertz spectroscopy to demonstrate the causal relation that the nonlinear excitation of phonons was mediated by a coherent magnon, as shown by the Fourier spectrum of the nonlinear amplitude in Fig. 8c. Specifically, a peak was observed only when the excitation frequency coincided with the magnon mode and the detection frequency coincided with the phonon mode; on the other hand, no peak was seen above the noise level when one swapped the excitation and detection frequency. The magnon-mediated excitation pathway for the coherent phonon is summarized in Fig. 8e. In the first step, an intense terahertz field resonantly populates the coherent magnonic state at frequency , creating an intermediate state. In the second step, another terahertz photon at frequency interacts with this intermediate state and coherently excites the phonon.
IV.4 Phonon-exciton coupling
When a coherent phonon mode is driven out of equilibrium following an optical pump, the resulting atomic motions can in turn modulate the optical properties in a coherent manner. In materials whose optical response is strongly modified by the presence of excitons — or bound electron-hole pairs — the exciton binding energy 211, 212, 213, lifetime 214, 215, and transport behavior 216 can also be changed. Such phonon-exciton coupling has been widely studied in quantum dots and quantum wells 217, 218, 219, and more recently, in complex bulk and atomically-thin materials 211, 212, 213. By comparing the exciton lineshape in broadband absorption or photoluminescence spectra before and after the photoexcitation event, one can observe how excitons respond to phonons in the time domain, yielding insights for tuning exciton properties via ultrafast phonon engineering.
For example, in anatase TiO2, a bulk semiconductor, Baldini et al. used photoinduced coherent acoustic phonons to control its excitons at room temperature, observing an exciton energy shift up to 30 meV, which is one of the largest reported in solids under a time-dependent perturbation 211. In a prototypical two-dimensional lead iodide perovskite, (PEA)2PbI4 (PEA = phenylethylammonium), it was further discovered that different excitons have varying coupling strengths with different coherent phonons, suggesting distinct polaronic characters of each exciton in this system and possibly in a broader class of hybrid semiconductors 212. In the atomically thin limit, Trovatello et al. also showed a pronounced phonon-exciton coupling effect in a monolayer MoS2, whose excitons are strongly affected by a photoinduced coherent phonon 213. Using transient absorption spectroscopy, the authors studied the interaction between out-of-plane phonon modes and excitons, noting a marked enhancement in the phonon oscillation amplitude at resonant exciton frequencies.
The aforementioned studies of phonon-exciton coupling, however, lack spatial information, which is important for learning the transport behavior of excitons under the influence of phonons. Recently, Peng and coworkers demonstrated microscopic imaging of directional transport of interlayer excitons in bilayer WSe2 with the assistance of surface acoustic waves (ref. 216), and their experimental scheme is summarized in Fig. 9a. In this geometry, excitons were optically excited by a continuous-wave laser on the left edge of the sample and probed at different locations via photoluminescence imaging, giving rise to a spatially-resolved mapping of exciton propagation. As shown in Fig. 9b, the excitons were launched near the position at 20 m, where the strongest photoluminescence intensity appears. In the absence of a surface acoustic wave, the exciton is largely localized (Fig. 9b). However, by launching a surface acoustic wave using an interdigital transducer, excitons can travel a long distance (m, Fig. 9c), which is much longer than their typical diffusion length. Despite the lack of temporal resolution in this measurement, one can imagine a spatiotemporal experimental scheme by using a pulsed laser to launch the excitons. Crucially, this experiment introduces a new way for controlling exciton transport via electrically-excited or photoinduced acoustic waves, a versatile approach that can be adapted to many other 2D semiconductors and heterostructures.
These studies underscore the diversity and tunability of phonon-exciton interactions across material systems, providing insights into phononic control over excitonic properties, which has far-reaching implications for the next-generation design of optoelectronic devices.
IV.5 Other collective excitation couplings
Despite the special focus on phonons as one of the coupled modes in the preceding sections, the full combinatorics of collective excitation coupling has been applied to many other scenarios such as magnon-magnon 222, 221, 220, 223, exciton-magnon 224, plasmon-polariton 225, 144 among other interactions. In this section, we discuss a few examples that showcase the diversity of collective excitation couplings and their role in determining the ultrafast dynamics of quantum materials.
As an analogy of nonlinear phononics that opens a new pathway for coherent manipulation of the crystalline structure, magnon-magnon coupling provides an avenue for controlling the spin dynamics. Utilizing 2D THz spectroscopy, Blank and coworkers revealed the down-conversion coupling of two magnon modes in the canted antiferromagnet FeBO3 (ref. 222). In another canted antiferromagnet, ErFeO3, Zhang and coworkers also demonstrated a unidirectional up-conversion coupling between distinct magnon modes (ref. 221). The canted spins in these materials result from the competition between antiferromagnetic interaction and Dzyaloshinskii-Moriya interaction, yielding a small net magnetic moment . As a result, there are two primary cooperative motions of the sublattice spins, corresponding to two distinct magnon modes sketched in Fig. 10a: the quasi-ferromagnetic (qFM) mode where precesses around the -axis, and the quasi-antiferromagnetic (qAFM) mode where oscillates. Given the different contributions to from the two modes, they can be selectively excited by tuning the direction of the magnetic field in the incident THz pump pulse, and their dynamics can be detected in the radiated free-induction decay signals. For example, when the magnetic field of the emitted THz pulse is parallel to that of the excitation field (), qFM mode and qAFM mode are selectively excited and observed when and , respectively (upper panel in Fig. 10b). This observation is expected as qFM involves a precession in the - plane while qAFM involves an amplitude oscillation along . Surprisingly, when , qAFM mode was still observed with (lower panel in Fig. 10b), which was not expected if there were only linear interactions between the driving THz magnetic field and magnons. To elicit the nonlinear interactions, 2D THz spectroscopy measurements were employed, where the nonlinear spectrum reveals a unidirectional up-conversion process from qFM to qAFM. As shown in Fig. 10c, the unidirectionality is evidenced by a single peak at the qFM excitation frequency and qAFM detection frequency, and importantly, a peak is absent when the corresponding excitation and detection frequencies are interchanged. Besides magnon up-conversion that echoes the phonon up-conversion seen in SrTiO3 (Fig. 5c, ref. 163), the sum and difference frequencies between these two magnon modes have been demonstrated in a cousin rare-earth ferrite, YFeO3, which are featured by a plethora of off-diagonal peaks in the nonlinear 2D terahertz spectrum in Fig. 10d (ref. 220). Like nonlinear phononics, these nonlinear effects of magnons demonstrated in the time domain help usher in new ways to coherently manipulate the spin degree of freedom in magnetic materials.
The analogy between phonons and magnons goes beyond anharmonic couplings. Like phonons, a magnon wavepacket can propagate in materials, which forms the bedrock for spintronic and magnonic applications. When coupled with other collective excitations, magnon propagation can be modified. For instance, for magnons that are coupled with acoustic phonons, theoretical 226 and experimental 227, 228 works have shown that magnon polarons travel over long distances with a faster group velocity than magnons alone. In another study of a van der Waals antiferromagnetic semiconductor CrSBr, magnons were observed to coherently travel beyond several micrometers with a very fast group velocity, an effect that Bae and coworkers interpreted as the coupling between the magnon and acoustic phonon branches 224, 229. What is interesting about this experiment is that coherent magnons were probed by laser pulses tuned to the exciton resonances in CrSBr, demonstrating an efficient optical access to spintronic information due to strong exciton-magnon couplings in this material. A subsequent experiment suggested that optically launched magnon propagation in CrSBr and likely other van der Waals magnet is instead mediated by long-range magnetic-dipole interactions 230, calling for additional spatiotemporally-resolved studies to understand magnon dynamics in this important class of materials.
Non-local propagation effects mediated by the coupling between collective excitations are not restricted to magnons. Other examples include exciton propagation with the aid of surface acoustic waves in a GaAs quantum well 231, 232 or in a transition metal dichalcogenide 216, and surface-plasmon-polariton propagation at a metal-dielectric interface 233. Similar to a phonon-polariton, a plasmon-polariton is a hybridized excitation between charge oscillations and light. More specifically, when light illuminates a metal-dielectric interface, the oscillating electric field of light drives a collective oscillation of the surface charge, which in turn radiates electromagnetic waves. As a result, a hybridized wave, namely surface-plasmon-polariton that involves both the charge motion on the metal surface and electromagnetic field in the dielectric, propagates along the interface. To observe and control its propagation dynamics, PINEM is one of the leading techniques 146, 144. Figure 11a shows the spatiotemporally-resolved imaging of surface-plasmon-polaritons induced by near-infrared laser pulses on an isolated silver nanowire suspended on a graphene film 144, where surface-plasmon-polaritons are evidenced by the additional peaks that appear in the electron energy loss spectra when the light and electron pulses are temporally overlapped. These peaks result from the quantized gain and loss of energy when electrons interact with the surface-plasmon-polariton field. Using energy-filtered electrons, the temporal dynamics of the surface-plasmon-polariton field was imaged with high spatial resolution as it forms an evolving interferometric standing wave pattern (Fig. 11b–f). Moreover, with different light polarizations, different surface-plasmon-polariton field distributions were obtained, demonstrating the potential to control surface-plasmon-polariton using light.
V Summary and outlook
The study and manipulation of coupled collective excitations by ultrafast light pulses offer a new perspective to disentangle intertwined degrees of freedom that are difficult to isolate in equilibrium. In this review, we have highlighted several couplings that both yield an in-depth understanding of the equilibrium properties and help one to coherently control the nonequilibrium dynamics in the host material.
While examples mentioned so far only include two participating collective excitations, multipartite couplings have been theoretically pondered, such as a hybrid exciton-phonon-photon that gives rise to a “phonoriton” in an optical cavity (Fig. 12a) 236, 234, and experimentally realized, such as the hybrid plasmon-exciton-photon — known as “plexcitons” — in WS2 (Fig. 12b), where cavity effects were achieved through nano-patterning (refs. 235, 237). These studies fall into a larger body of works that aims to manipulate quantum materials via cavity control, in which the light-matter coupling strength can be greatly enhanced by replacing the classical laser fields with quantum-mechanical photon modes. One of the most celebrated examples is microcavity exciton–polaritons in GaAs quantum wells, which arise from the strong coupling between photons and excitons and exhibit remarkable effects such as Bose-Einstein condensation 238, 239, 240. More recently, a multitude of theoretical proposals have been put forward for tailoring the unique properties of quantum materials via cavity design 241, 242, with experimental progress in the study of topological state 243, superconductivity 244, and phase transition 245 in different material platforms.
Besides the potential in the experimental realization of such multi-way couplings in the context of cavity quantum materials, we comment on a few other directions in the study of coupled excitations from the viewpoint of ultrafast science.
First, advances in engineering the waveform of the excitation light pulse — such as its polarization, carrier envelope phase, and orbital angular momentum — can greatly expand the tool kits for resonant excitation of collective modes and hence a deterministic control over material properties. A recent example is the synthesis of an intense, circularly-polarized THz pulse used to induce circular motions of atoms, otherwise known as chiral phonons, which can in turn induce a large effective magnetic field 76, 71, 70, 69. In particular, Davies et al. demonstrated that the induced field can significantly alter the magnetic order even at considerable distances from the phonon creation point 70, underscoring the need for a deeper insight into chiral phonon transport. Equipped with the capability to deterministically excite particular phonon and magnon modes using the mechanisms introduced in this review, one can envision future experiments where chiral phonons and magnons are created at distinct sample locations, and their mutual interaction over space and time can be used to gain a mechanistic understanding of the propagation mechanism for realizing ultrafast non-local control of magnetic ordering.
Second, from a technical point of view, great care needs to be taken when comparing the dynamics of different collective excitations probed by distinct experimental setups, making a quantitative study of their couplings challenging. An example of this issue arises when one extracts the electron-phonon coupling strength by comparing the electronic band change and atomic displacement, which are probed by time-resolved photoemission and X-ray diffraction respectively 19, 22. For the same incident pump fluence, the effective excitation fluence can differ substantially because photoemission is a surface-sensitive probe (sub- to few nanometers), whereas X-rays can probe much deeper into the bulk (tens of nanometers to micrometers). Therefore, careful fluence calibrations are needed when comparing these probes unless monolayer samples are used to eliminate the variance. An alternative route is to look for complementary probes with a similar probing depth, such as high-energy electron diffraction and extreme ultraviolet absorption spectroscopy, both operated under a transmission geometry. While the former is sensitive to the lattice degree of freedom, the latter excels in capturing carrier, magnetic, and excitonic excitations with an exceptional time resolution 25, 246. One can envision a chamber design that integrates both probes in a single setup, making it possible to concurrently access lattice and electronic information following the same excitation pulse.
Third, while the review has largely focused on coherent collective dynamics with a well-defined phase, not all light-induced collective excitations are coherent or fully reproducible over multiple pump pulses, particularly in systems with spatial inhomogeneities where defects and disorders are prevalent 247, 248, 249, 250. In this aspect, techniques with both spatial and temporal resolutions that operate in the single-shot mode are important for revealing the local, incoherent, and stochastic dynamics that vary from pulse to pulse, as recently demonstrated in the study of magnetic fluctuations in magnetic multilayer thin films 138, 139 and the heterogeneous insulator-to-metal transition in VO2 (ref. 140), both of which utilized a coherent X-ray source in large-scale user facilities. It is worth noting that stochastic dynamics can also be accessed without relying on single-shot detection 251, which makes it possible to study fluctuating systems with the stroboscopic scheme employed in the vast majority of ultrafast measurements. Along with rapid progress in facility-based ultrashort coherent X-ray sources, we also envision two key developments in table-top experiments for probing local and incoherent collective modes. In the extreme spatial scale, innovations in scanning tunneling microscopy has enabled new ways to image collective excitations at the atomic scale 141, 142, paving the way for controlling these local dynamics through creative engineering of nanoscopic defects. In the extreme temporal scale, ultrafast electron diffraction and microscopy that have witnessed a rapid improvement of their temporal resolution 252, 109, 253, 254 would provide a useful platform to study the initial emergence of incoherent phonons following photoexcitation 255, 256, which play an important role in determining the ensuing defect structure 53 and energy flow in the host material system 257.
As we conclude this review that has touched a diverse set of experimental techniques and observables, it is an apt place to be reminded that the overarching goal of studying coupled collective excitations is to uncover a set of unifying principles that govern the interaction among the different degrees of freedom in solids. While studying collective excitations is certainly not the only way to unravel the intricate microscopic interactions, they offer us a way to describe a many-body system without the need to keep track of individual dynamics of an astronomical number of particles, making it tractable to study and engineer material properties both in and out of equilibrium.
Acknowledgements.
We thank helpful input from Edoardo Baldini, Henry Bell, Thomas Blank, Yu-Che Chien, Shaofeng Duan, Nuh Gedik, Patrick Liu, Tianchuang Luo, Mariano Trigo, and Xunqing Yin during the preparation of this manuscript.References
- Lee et al. [2006] P. A. Lee, N. Nagaosa, and X.-G. Wen, Doping a Mott insulator: Physics of high-temperature superconductivity, Rev. Mod. Phys. 78, 17 (2006).
- Stewart [2011] G. R. Stewart, Superconductivity in iron compounds, Rev. Mod. Phys. 83, 1589 (2011).
- Keimer et al. [2015] B. Keimer, S. A. Kivelson, M. R. Norman, S. Uchida, and J. Zaanen, From quantum matter to high-temperature superconductivity in copper oxides, Nature 518, 179 (2015).
- Fernandes et al. [2022] R. M. Fernandes, A. I. Coldea, H. Ding, I. R. Fisher, P. J. Hirschfeld, and G. Kotliar, Iron pnictides and chalcogenides: a new paradigm for superconductivity, Nature 601, 35 (2022).
- Cheong and Mostovoy [2007] S.-W. Cheong and M. Mostovoy, Multiferroics: a magnetic twist for ferroelectricity, Nat. Mater. 6, 13 (2007).
- Salamon and Jaime [2001] M. B. Salamon and M. Jaime, The physics of manganites: Structure and transport, Rev. Mod. Phys. 73, 583 (2001).
- Wang and Zhang [2017] J. Wang and S.-C. Zhang, Topological states of condensed matter, Nat. Mater. 16, 1062 (2017).
- Tokura et al. [2017] Y. Tokura, M. Kawasaki, and N. Nagaosa, Emergent functions of quantum materials, Nat. Phys. 13, 1056 (2017).
- Nayak et al. [2008] C. Nayak, S. H. Simon, A. Stern, M. Freedman, and S. Das Sarma, Non-Abelian anyons and topological quantum computation, Rev. Mod. Phys. 80, 1083 (2008).
- Žutić et al. [2004] I. Žutić, J. Fabian, and S. Das Sarma, Spintronics: Fundamentals and applications, Rev. Mod. Phys. 76, 323 (2004).
- Cooper [1956] L. N. Cooper, Bound electron pairs in a degenerate Fermi gas, Phys. Rev. 104, 1189 (1956).
- Bardeen et al. [1957] J. Bardeen, L. N. Cooper, and J. R. Schrieffer, Theory of superconductivity, Phys. Rev. 108, 1175 (1957).
- Petek and Ogawa [1997] H. Petek and S. Ogawa, Femtosecond time-resolved two-photon photoemission studies of electron dynamics in metals, Prog. Surf. Sci. 56, 239–310 (1997).
- Bovensiepen and Kirchmann [2012] U. Bovensiepen and P. S. Kirchmann, Elementary relaxation processes investigated by femtosecond photoelectron spectroscopy of two-dimensional materials, Laser Photonics Rev. 6, 589–606 (2012).
- Zeiger et al. [1992] H. J. Zeiger, J. Vidal, T. K. Cheng, E. P. Ippen, G. Dresselhaus, and M. S. Dresselhaus, Theory for displacive excitation of coherent phonons, Phys. Rev. B 45, 768 (1992).
- Stevens et al. [2002] T. E. Stevens, J. Kuhl, and R. Merlin, Coherent phonon generation and the two stimulated Raman tensors, Phys. Rev. B 65, 144304 (2002).
- Bugayev et al. [2011] A. Bugayev, A. Esmail, M. Abdel-Fattah, and H. E. Elsayed-Ali, Coherent phonons in bismuth film observed by ultrafast electron diffraction, AIP Adv. 1, 012117 (2011).
- Chatelain et al. [2014] R. P. Chatelain, V. R. Morrison, B. L. M. Klarenaar, and B. J. Siwick, Coherent and incoherent electron-phonon coupling in graphite observed with radio-frequency compressed ultrafast electron diffraction, Phys. Rev. Lett. 113, 235502 (2014).
- Gerber et al. [2017] S. Gerber, S. L. Yang, D. Zhu, H. Soifer, J. A. Sobota, S. Rebec, J. J. Lee, T. Jia, B. Moritz, C. Jia, A. Gauthier, Y. Li, D. Leuenberger, Y. Zhang, L. Chaix, W. Li, H. Jang, J. S. Lee, M. Yi, G. L. Dakovski, S. Song, J. M. Glownia, S. Nelson, K. W. Kim, Y. D. Chuang, Z. Hussain, R. G. Moore, T. P. Devereaux, W. S. Lee, P. S. Kirchmann, and Z. X. Shen, Femtosecond electron-phonon lock-in by photoemission and x-ray free-electron laser, Science 357, 71 (2017).
- Zong et al. [2018] A. Zong, X. Shen, A. Kogar, L. Ye, C. Marks, D. Chowdhury, T. Rohwer, B. Freelon, S. Weathersby, R. Li, J. Yang, J. Checkelsky, X. Wang, and N. Gedik, Ultrafast manipulation of mirror domain walls in a charge density wave, Sci. Adv. 4, eaau5501 (2018).
- Zong et al. [2023a] A. Zong, Q. Zhang, F. Zhou, Y. Su, K. Hwangbo, X. Shen, Q. Jiang, H. Liu, T. E. Gage, D. A. Walko, M. E. Kozina, D. Luo, A. H. Reid, J. Yang, S. Park, S. H. Lapidus, J.-H. Chu, I. Arslan, X. Wang, D. Xiao, X. Xu, N. Gedik, and H. Wen, Spin-mediated shear oscillators in a van der Waals antiferromagnet, Nature 620, 988 (2023a).
- Huang et al. [2023] Y. Huang, J. D. Querales-Flores, S. W. Teitelbaum, J. Cao, T. Henighan, H. Liu, M. Jiang, G. De la Peña, V. Krapivin, J. Haber, T. Sato, M. Chollet, D. Zhu, T. Katayama, R. Power, M. Allen, C. R. Rotundu, T. P. Bailey, C. Uher, M. Trigo, P. S. Kirchmann, É. D. Murray, Z.-X. Shen, I. Savić, S. Fahy, J. A. Sobota, and D. A. Reis, Ultrafast measurements of mode-specific deformation potentials of Bi2Te3 and Bi2Se3, Phys. Rev. X 13, 041050 (2023).
- Zong et al. [2019a] A. Zong, P. E. Dolgirev, A. Kogar, E. Ergeçen, M. B. Yilmaz, Y.-Q. Bie, T. Rohwer, I. C. Tung, J. Straquadine, X. Wang, Y. Yang, X. Shen, R. Li, J. Yang, S. Park, M. C. Hoffmann, B. K. Ofori-Okai, M. E. Kozina, H. Wen, X. Wang, I. R. Fisher, P. Jarillo-Herrero, and N. Gedik, Dynamical slowing-down in an ultrafast photoinduced phase transition, Phys. Rev. Lett. 123, 097601 (2019a).
- Bartram et al. [2022] F. M. Bartram, Y.-C. Leng, Y. Wang, L. Liu, X. Chen, H. Peng, H. Li, P. Yu, Y. Wu, M.-L. Lin, J. Zhang, P.-H. Tan, and L. Yang, Ultrafast coherent interlayer phonon dynamics in atomically thin layers of MnBi2Te4, npj Quantum Mater. 7, 84 (2022).
- Zong et al. [2023b] A. Zong, B. R. Nebgen, S.-C. Lin, J. A. Spies, and M. Zuerch, Emerging ultrafast techniques for studying quantum materials, Nat. Rev. Mater. 8, 224 (2023b).
- de la Torre et al. [2021] A. de la Torre, D. M. Kennes, M. Claassen, S. Gerber, J. W. McIver, and M. A. Sentef, Colloquium: Nonthermal pathways to ultrafast control in quantum materials, Rev. Mod. Phys. 93, 041002 (2021).
- Fausti and van Loosdrecht [2011] D. Fausti and P. H. M. van Loosdrecht, Optical techniques for solid-state materials characterization (CRC Press Boca Raton, 2011).
- Orenstein [2012] J. Orenstein, Ultrafast spectroscopy of quantum materials, Phys. Today 65, 44 (2012).
- Woerner et al. [2013] M. Woerner, W. Kuehn, P. Bowlan, K. Reimann, and T. Elsaesser, Ultrafast two-dimensional terahertz spectroscopy of elementary excitations in solids, New J. Phys. 15, 025039 (2013).
- Giannetti et al. [2016] C. Giannetti, M. Capone, D. Fausti, M. Fabrizio, F. Parmigiani, and D. Mihailovic, Ultrafast optical spectroscopy of strongly correlated materials and high-temperature superconductors: a non-equilibrium approach, Adv. Phys. 65, 58 (2016).
- Lloyd-Hughes et al. [2021] J. Lloyd-Hughes, P. M. Oppeneer, T. Pereira dos Santos, A. Schleife, S. Meng, M. A. Sentef, M. Ruggenthaler, A. Rubio, I. Radu, M. Murnane, X. Shi, H. Kapteyn, B. Stadtmüller, K. M. Dani, F. H. da Jornada, E. Prinz, M. Aeschlimann, R. L. Milot, M. Burdanova, J. Boland, T. Cocker, and F. Hegmann, The 2021 ultrafast spectroscopic probes of condensed matter roadmap, J. Phys. Condens. Matter 33, 353001 (2021).
- Dong et al. [2023] T. Dong, S.-J. Zhang, and N.-L. Wang, Recent development of ultrafast optical characterizations for quantum materials, Adv. Mater. 35, 2110068 (2023).
- Zong et al. [2021] A. Zong, A. Kogar, and N. Gedik, Unconventional light-induced states visualized by ultrafast electron diffraction and microscopy, MRS Bull. 46, 720 (2021).
- Filippetto et al. [2022] D. Filippetto, P. Musumeci, R. K. Li, B. J. Siwick, M. R. Otto, M. Centurion, and J. P. F. Nunes, Ultrafast electron diffraction: Visualizing dynamic states of matter, Rev. Mod. Phys. 94, 045004 (2022).
- Lee et al. [2024] Y. Lee, K. Y. Oang, D. Kim, and H. Ihee, A comparative review of time-resolved x-ray and electron scattering to probe structural dynamics, Struct. Dyn. 11, 031301 (2024).
- Buzzi et al. [2018] M. Buzzi, M. Först, R. Mankowsky, and A. Cavalleri, Probing dynamics in quantum materials with femtosecond X-rays, Nat. Rev. Mater. 3, 299 (2018).
- Cao et al. [2019] Y. Cao, D. G. Mazzone, D. Meyers, J. P. Hill, X. Liu, S. Wall, and M. P. M. Dean, Ultrafast dynamics of spin and orbital correlations in quantum materials: an energy- and momentum-resolved perspective, Philos. Trans. R. Soc. A 377, 20170480 (2019).
- Mitrano and Wang [2020] M. Mitrano and Y. Wang, Probing light-driven quantum materials with ultrafast resonant inelastic X-ray scattering, Commun. Phys. 3, 184 (2020).
- Mitrano et al. [2024] M. Mitrano, S. Johnston, Y.-J. Kim, and M. P. M. Dean, Exploring quantum materials with resonant inelastic x-ray scattering, Phys. Rev. X 14, 040501 (2024).
- Bovensiepen and Kirchmann [2012] U. Bovensiepen and P. S. Kirchmann, Elementary relaxation processes investigated by femtosecond photoelectron spectroscopy of two-dimensional materials, Laser Photonics Rev. 6, 589 (2012).
- Smallwood et al. [2016] C. L. Smallwood, R. A. Kaindl, and A. Lanzara, Ultrafast angle-resolved photoemission spectroscopy of quantum materials, EPL 115, 27001 (2016).
- Zhou et al. [2018] X. Zhou, S. He, G. Liu, L. Zhao, L. Yu, and W. Zhang, New developments in laser-based photoemission spectroscopy and its scientific applications: a key issues review, Rep. Prog. Phys. 81, 062101 (2018).
- Huang et al. [2022] C. Huang, S. Duan, and W. Zhang, High-resolution time-and angle-resolved photoemission studies on quantum materials, Quantum Frontiers 1, 15 (2022).
- Boschini et al. [2024] F. Boschini, M. Zonno, and A. Damascelli, Time-resolved ARPES studies of quantum materials, Rev. Mod. Phys. 96, 015003 (2024).
- Mankowsky et al. [2016] R. Mankowsky, M. Först, and A. Cavalleri, Non-equilibrium control of complex solids by nonlinear phononics, Rep. Prog. Phys. 79, 064503 (2016).
- Nicoletti and Cavalleri [2016] D. Nicoletti and A. Cavalleri, Nonlinear light–matter interaction at terahertz frequencies, Adv. Opt. Photonics 8, 401 (2016).
- Disa et al. [2021] A. S. Disa, T. F. Nova, and A. Cavalleri, Engineering crystal structures with light, Nat. Phys. 17, 1087 (2021).
- Oka and Kitamura [2019] T. Oka and S. Kitamura, Floquet engineering of quantum materials, Annu. Rev. Condens. Matter Phys 10, 387 (2019).
- Rudner and Lindner [2020] M. S. Rudner and N. H. Lindner, Band structure engineering and non-equilibrium dynamics in Floquet topological insulators, Nat. Rev. Phys. 2, 229 (2020).
- Bao et al. [2022] C. Bao, P. Tang, D. Sun, and S. Zhou, Light-induced emergent phenomena in 2D materials and topological materials, Nat. Rev. Phys. 4, 33 (2022).
- Kirilyuk et al. [2010] A. Kirilyuk, A. V. Kimel, and T. Rasing, Ultrafast optical manipulation of magnetic order, Rev. Mod. Phys. 82, 2731 (2010).
- Zhang and Averitt [2014] J. Zhang and R. D. Averitt, Dynamics and control in complex transition metal oxides, Annu. Rev. Mater. Res. 44, 19 (2014).
- Cheng et al. [2024] Y. Cheng, A. Zong, L. Wu, Q. Meng, W. Xia, F. Qi, P. Zhu, X. Zou, T. Jiang, Y. Guo, J. van Wezel, A. Kogar, M. W. Zuerch, J. Zhang, Y. Zhu, and D. Xiang, Ultrafast formation of topological defects in a two-dimensional charge density wave, Nat. Phys. 20, 54 (2024).
- Baldini [2018] E. Baldini, Nonequilibrium dynamics of collective excitations in quantum materials (Springer, 2018).
- Basov et al. [2016] D. N. Basov, M. M. Fogler, and F. J. García de Abajo, Polaritons in van der Waals materials, Science 354, aag1992 (2016).
- Basov et al. [2020] D. N. Basov, A. Asenjo-Garcia, P. J. Schuck, X. Zhu, and A. Rubio, Polariton panorama, Nanophotonics 10, 449 (2020).
- Baldini et al. [2020] E. Baldini, C. A. Belvin, M. Rodriguez-Vega, I. O. Ozel, D. Legut, A. Kozłowski, A. M. Oleś, K. Parlinski, P. Piekarz, J. Lorenzana, G. A. Fiete, and N. Gedik, Discovery of the soft electronic modes of the trimeron order in magnetite, Nat. Phys. 16, 541–545 (2020).
- Basov et al. [2017] D. N. Basov, R. D. Averitt, and D. Hsieh, Towards properties on demand in quantum materials, Nat. Mater. 16, 1077 (2017).
- Demsar et al. [1999] J. Demsar, K. Biljaković , and D. Mihailovic, Single particle and collective excitations in the one-dimensional charge density wave solid K0.3MoO3 probed in real time by femtosecond spectroscopy, Phys. Rev. Lett. 83, 800 (1999).
- Demsar et al. [2002] J. Demsar, L. Forró, H. Berger, and D. Mihailovic, Femtosecond snapshots of gap-forming charge-density-wave correlations in quasi-two-dimensional dichalcogenides 1T-TaS2 and 2H-TaSe2, Phys. Rev. B 66, 041101 (2002).
- Schmitt et al. [2008] F. Schmitt, P. S. Kirchmann, U. Bovensiepen, R. G. Moore, L. Rettig, M. Krenz, J. H. Chu, N. Ru, L. Perfetti, D. H. Lu, M. Wolf, I. R. Fisher, and Z. X. Shen, Transient electronic structure and melting of a charge density wave in TbTe3, Science 321, 1649 (2008).
- Yusupov et al. [2010] R. Yusupov, T. Mertelj, V. V. Kabanov, S. Brazovskii, P. Kusar, J.-H. Chu, I. R. Fisher, and D. Mihailovic, Coherent dynamics of macroscopic electronic order through a symmetry breaking transition, Nat. Phys. 6, 681–684 (2010).
- Mihailovic [2019] D. Mihailovic, The importance of topological defects in photoexcited phase transitions including memory applications, Appl. Sci. 9, 890 (2019).
- Horstmann et al. [2020] J. G. Horstmann, H. Böckmann, B. Wit, F. Kurtz, G. Storeck, and C. Ropers, Coherent control of a surface structural phase transition, Nature 583, 232–236 (2020).
- Maklar et al. [2023] J. Maklar, J. Sarkar, S. Dong, Y. A. Gerasimenko, T. Pincelli, S. Beaulieu, P. S. Kirchmann, J. A. Sobota, S. Yang, D. Leuenberger, R. G. Moore, Z.-X. Shen, M. Wolf, D. Mihailovic, R. Ernstorfer, and L. Rettig, Coherent light control of a metastable hidden state, Sci. Adv. 9, eadi4661 (2023).
- Maehrlein et al. [2017] S. Maehrlein, A. Paarmann, M. Wolf, and T. Kampfrath, Terahertz sum-frequency excitation of a Raman-active phonon, Phys. Rev. Lett. 119, 127402 (2017).
- Shi et al. [2023] J. Shi, Y.-Q. Bie, A. Zong, S. Fang, W. Chen, J. Han, Z. Cao, Y. Zhang, T. Taniguchi, K. Watanabe, X. Fu, V. Bulović, E. Kaxiras, E. Baldini, P. Jarillo-Herrero, and K. A. Nelson, Intrinsic phase induced in atomically thin 2-MoTe2 by a single terahertz pulse, Nat. Commun. 14, 5905 (2023).
- Juraschek et al. [2017] D. M. Juraschek, M. Fechner, A. V. Balatsky, and N. A. Spaldin, Dynamical multiferroicity, Phys. Rev. Mater. 1, 014401 (2017).
- Basini et al. [2024] M. Basini, M. Pancaldi, B. Wehinger, M. Udina, V. Unikandanunni, T. Tadano, M. C. Hoffmann, A. V. Balatsky, and S. Bonetti, Terahertz electric-field-driven dynamical multiferroicity in SrTiO3, Nature 628, 534 (2024).
- Davies et al. [2024] C. S. Davies, F. G. N. Fennema, A. Tsukamoto, I. Razdolski, A. V. Kimel, and A. Kirilyuk, Phononic switching of magnetization by the ultrafast Barnett effect, Nature 628, 540 (2024).
- Romao and Juraschek [2024] C. P. Romao and D. M. Juraschek, Light makes atoms behave like electromagnetic coils, Nature 628, 505 (2024).
- Sie et al. [2019] E. J. Sie, C. M. Nyby, C. D. Pemmaraju, S. J. Park, X. Shen, J. Yang, M. C. Hoffmann, B. K. Ofori-Okai, R. Li, A. H. Reid, S. Weathersby, E. Mannebach, N. Finney, D. Rhodes, D. Chenet, A. Antony, L. Balicas, J. Hone, T. P. Devereaux, T. F. Heinz, X. Wang, and A. M. Lindenberg, An ultrafast symmetry switch in a Weyl semimetal, Nature 565, 61–66 (2019).
- Tang et al. [2023] R. Tang, F. Boi, and Y.-H. Cheng, Light-induced topological phase transition via nonlinear phononics in superconductor CsV3Sb5, npj Quantum Mater. 8, 78 (2023).
- Juraschek et al. [2022] D. M. Juraschek, T. Neuman, and P. Narang, Giant effective magnetic fields from optically driven chiral phonons in 4 paramagnets, Phys. Rev. Res. 4, 013129 (2022).
- Kahana et al. [2024] T. Kahana, D. A. Bustamante Lopez, and D. M. Juraschek, Light-induced magnetization from magnonic rectification, Sci. Adv. 10, eado0722 (2024).
- Luo et al. [2023a] J. Luo, T. Lin, J. Zhang, X. Chen, E. R. Blackert, R. Xu, B. I. Yakobson, and H. Zhu, Large effective magnetic fields from chiral phonons in rare-earth halides, Science 382, 698 (2023a).
- Ilyas et al. [2024] B. Ilyas, T. Luo, A. von Hoegen, E. Viñas Boström, Z. Zhang, J. Park, J. Kim, J.-G. Park, K. A. Nelson, A. Rubio, and N. Gedik, Terahertz field-induced metastable magnetization near criticality in FePS3, Nature 636, 609 (2024).
- Yoshikawa et al. [2021] N. Yoshikawa, H. Suganuma, H. Matsuoka, Y. Tanaka, P. Hemme, M. Cazayous, Y. Gallais, M. Nakano, Y. Iwasa, and R. Shimano, Ultrafast switching to an insulating-like metastable state by amplitudon excitation of a charge density wave, Nat. Phys. 17, 909 (2021).
- Wang et al. [2013] Y. H. Wang, H. Steinberg, P. Jarillo-Herrero, and N. Gedik, Observation of Floquet-Bloch states on the surface of a topological insulator, Science 342, 453 (2013).
- Mahmood et al. [2016] F. Mahmood, C.-K. Chan, Z. Alpichshev, D. Gardner, Y. Lee, P. A. Lee, and N. Gedik, Selective scattering between Floquet-Bloch and Volkov states in a topological insulator, Nat. Phys. 12, 306 (2016).
- Zhou et al. [2023a] S. Zhou, C. Bao, B. Fan, H. Zhou, Q. Gao, H. Zhong, T. Lin, H. Liu, P. Yu, P. Tang, S. Meng, W. Duan, and S. Zhou, Pseudospin-selective Floquet band engineering in black phosphorus, Nature 614, 75 (2023a).
- Zhou et al. [2023b] S. Zhou, C. Bao, B. Fan, F. Wang, H. Zhong, H. Zhang, P. Tang, W. Duan, and S. Zhou, Floquet engineering of black phosphorus upon below-gap pumping, Phys. Rev. Lett. 131, 116401 (2023b).
- Ito et al. [2023] S. Ito, M. Schüler, M. Meierhofer, S. Schlauderer, J. Freudenstein, J. Reimann, D. Afanasiev, K. A. Kokh, O. E. Tereshchenko, J. Güdde, M. A. Sentef, U. Höfer, and R. Huber, Build-up and dephasing of Floquet–Bloch bands on subcycle timescales, Nature 616, 696–701 (2023).
- Choi et al. [2024] D. Choi, M. Mogi, U. De Giovannini, D. Azoury, B. Lv, Y. Su, H. Hübener, A. Rubio, and N. Gedik, Direct observation of Floquet-Bloch states in monolayer graphene (2024), arXiv:2404.14392 .
- Merboldt et al. [2024] M. Merboldt, M. Schüler, D. Schmitt, J. P. Bange, W. Bennecke, K. Gadge, K. Pierz, H. W. Schumacher, D. Momeni, D. Steil, S. R. Manmana, M. Sentef, M. Reutzel, and S. Mathias, Observation of Floquet states in graphene (2024), arXiv:2404.12791 .
- Hübener et al. [2018] H. Hübener, U. De Giovannini, and A. Rubio, Phonon driven Floquet matter, Nano Lett. 18, 1535–1542 (2018).
- Chaudhary et al. [2020] S. Chaudhary, A. Haim, Y. Peng, and G. Refael, Phonon-induced Floquet topological phases protected by space-time symmetries, Phys. Rev. Res. 2, 043431 (2020).
- Chan et al. [2023] Y.-H. Chan, D. Y. Qiu, F. H. da Jornada, and S. G. Louie, Giant self-driven exciton-Floquet signatures in time-resolved photoemission spectroscopy of MoS2 from time-dependent GW approach, Proc. Natl. Acad. Sci. U.S.A. 120, e2301957120 (2023).
- Pareek et al. [2024] V. Pareek, D. R. Bacon, X. Zhu, Y.-H. Chan, F. Bussolotti, N. S. Chan, J. Pérez Urquizo, K. Watanabe, T. Taniguchi, M. K. L. Man, J. Madéo, D. Y. Qiu, K. Eng Johnson Goh, F. H. da Jornada, and K. M. Dani, Driving non-trivial quantum phases in conventional semiconductors with intense excitonic fields (2024), arXiv:2403.08725 .
- Gao et al. [2024] F. Y. Gao, X. Peng, X. Cheng, E. Viñas Boström, D. S. Kim, R. K. Jain, D. Vishnu, K. Raju, R. Sankar, S.-F. Lee, M. A. Sentef, T. Kurumaji, X. Li, P. Tang, A. Rubio, and E. Baldini, Giant chiral magnetoelectric oscillations in a van der Waals multiferroic, Nature 632, 273 (2024).
- Golias and Sánchez-Barriga [2016] E. Golias and J. Sánchez-Barriga, Observation of antiphase coherent phonons in the warped Dirac cone of Bi2Te3, Phys. Rev. B 94, 161113 (2016).
- Ron et al. [2020] A. Ron, S. Chaudhary, G. Zhang, H. Ning, E. Zoghlin, S. D. Wilson, R. D. Averitt, G. Refael, and D. Hsieh, Ultrafast enhancement of ferromagnetic spin exchange induced by ligand-to-metal charge transfer, Phys. Rev. Lett. 125, 197203 (2020).
- Ergeçen et al. [2023] E. Ergeçen, B. Ilyas, J. Kim, J. Park, M. B. Yilmaz, T. Luo, D. Xiao, S. Okamoto, J.-G. Park, and N. Gedik, Coherent detection of hidden spin-lattice coupling in a van der Waals antiferromagnet, Proc. Natl. Acad. Sci. U.S.A. 120, e2208968120 (2023).
- Su et al. [2023] Y. Su, A. Zong, A. Kogar, D. Lu, S. S. Hong, B. Freelon, T. Rohwer, B. Y. Wang, H. Y. Hwang, and N. Gedik, Delamination-assisted ultrafast wrinkle formation in a freestanding film, Nano Lett. 23, 10772 (2023).
- Geneaux et al. [2019] R. Geneaux, H. J. B. Marroux, A. Guggenmos, D. M. Neumark, and S. R. Leone, Transient absorption spectroscopy using high harmonic generation: a review of ultrafast X-ray dynamics in molecules and solids, Philos. Trans. R. Soc. A 377, 20170463 (2019).
- Li et al. [2020] J. Li, J. Lu, A. Chew, S. Han, J. Li, Y. Wu, H. Wang, S. Ghimire, and Z. Chang, Attosecond science based on high harmonic generation from gases and solids, Nat. Commun. 11, 2748 (2020).
- Zhang et al. [2009] G. P. Zhang, W. Hübner, G. Lefkidis, Y. Bai, and T. F. George, Paradigm of the time-resolved magneto-optical Kerr effect for femtosecond magnetism, Nat. Phys. 5, 499 (2009).
- Nowakowski et al. [2015] P. J. Nowakowski, D. A. Woods, C. D. Bain, and J. R. R. Verlet, Time-resolved phase-sensitive second harmonic generation spectroscopy, J. Chem. Phys. 142, 084201 (2015).
- Zhang et al. [2021a] Y. Zhang, J. Dai, X. Zhong, D. Zhang, G. Zhong, and J. Li, Probing ultrafast dynamics of ferroelectrics by time-resolved pump-probe spectroscopy, Adv. Sci. 8, 2102488 (2021a).
- Yoshizawa and Kurosawa [1999] M. Yoshizawa and M. Kurosawa, Femtosecond time-resolved Raman spectroscopy using stimulated Raman scattering, Phys. Rev. A 61, 013808 (1999).
- Saichu et al. [2009] R. P. Saichu, I. Mahns, A. Goos, S. Binder, P. May, S. G. Singer, B. Schulz, A. Rusydi, J. Unterhinninghofen, D. Manske, P. Guptasarma, M. S. Williamsen, and M. Rübhausen, Two-component dynamics of the order parameter of high temperature Bi2Sr2CaCu2O8+δ superconductors revealed by time-resolved Raman scattering, Phys. Rev. Lett. 102, 177004 (2009).
- Versteeg et al. [2018] R. B. Versteeg, J. Zhu, P. Padmanabhan, C. Boguschewski, R. German, M. Goedecke, P. Becker, and P. H. M. van Loosdrecht, A tunable time-resolved spontaneous Raman spectroscopy setup for probing ultrafast collective excitation and quasiparticle dynamics in quantum materials, Struct. Dyn. 5, 044301 (2018).
- Yang et al. [2020] J.-A. Yang, N. Pellatz, T. Wolf, R. Nandkishore, and D. Reznik, Ultrafast magnetic dynamics in insulating YBa2Cu3O6.1 revealed by time resolved two-magnon Raman scattering, Nat. Commun. 11, 2548 (2020).
- Buchenau et al. [2023] S. Buchenau, B. Grimm-Lebsanft, F. Biebl, T. Glier, L. Westphal, J. Reichstetter, D. Manske, M. Fechner, A. Cavalleri, S. Herres-Pawlis, and M. Rübhausen, Optically induced avoided crossing in graphene, Phys. Rev. B 108, 075419 (2023).
- Breusing et al. [2009] M. Breusing, C. Ropers, and T. Elsaesser, Ultrafast carrier dynamics in graphite, Phys. Rev. Lett. 102, 086809 (2009).
- Ulbricht et al. [2011] R. Ulbricht, E. Hendry, J. Shan, T. F. Heinz, and M. Bonn, Carrier dynamics in semiconductors studied with time-resolved terahertz spectroscopy, Rev. Mod. Phys. 83, 543 (2011).
- Lagarde et al. [2014] D. Lagarde, L. Bouet, X. Marie, C. R. Zhu, B. L. Liu, T. Amand, P. H. Tan, and B. Urbaszek, Carrier and polarization dynamics in monolayer MoS2, Phys. Rev. Lett. 112, 047401 (2014).
- Robert et al. [2016a] C. Robert, D. Lagarde, F. Cadiz, G. Wang, B. Lassagne, T. Amand, A. Balocchi, P. Renucci, S. Tongay, B. Urbaszek, and X. Marie, Exciton radiative lifetime in transition metal dichalcogenide monolayers, Phys. Rev. B 93, 205423 (2016a).
- Qi et al. [2020] F. Qi, Z. Ma, L. Zhao, Y. Cheng, W. Jiang, C. Lu, T. Jiang, D. Qian, Z. Wang, W. Zhang, P. Zhu, X. Zou, W. Wan, D. Xiang, and J. Zhang, Breaking 50 femtosecond resolution barrier in MeV ultrafast electron diffraction with a double bend achromat compressor, Phys. Rev. Lett. 124, 134803 (2020).
- Beaud et al. [2009] P. Beaud, S. L. Johnson, E. Vorobeva, U. Staub, R. A. D. Souza, C. J. Milne, Q. X. Jia, and G. Ingold, Ultrafast structural phase transition driven by photoinduced melting of charge and orbital order, Phys. Rev. Lett. 103, 155702 (2009).
- Li et al. [2016] J. Li, W.-G. Yin, L. Wu, P. Zhu, T. Konstantinova, J. Tao, J. Yang, S.-W. Cheong, F. Carbone, J. A. Misewich, J. P. Hill, X. Wang, R. J. Cava, and Y. Zhu, Dichotomy in ultrafast atomic dynamics as direct evidence of polaron formation in manganites, npj Quantum Mater. 1, 16026 (2016).
- Kirkland [2020] E. J. Kirkland, Advanced computing in electron microscopy, Vol. 12 (Springer, 2020).
- Wall et al. [2018] S. Wall, S. Yang, L. Vidas, M. Chollet, J. M. Glownia, M. Kozina, T. Katayama, T. Henighan, M. Jiang, T. A. Miller, D. A. Reis, L. A. Boatner, O. Delaire, and M. Trigo, Ultrafast disordering of vanadium dimers in photoexcited VO2, Science 362, 572 (2018).
- Dürr et al. [2021] H. A. Dürr, R. Ernstorfer, and B. J. Siwick, Revealing momentum-dependent electron-phonon and phonon-phonon coupling in complex materials with ultrafast electron diffuse scattering, MRS Bull. 46, 731 (2021).
- Zong et al. [2021] A. Zong, P. E. Dolgirev, A. Kogar, Y. Su, X. Shen, J. A. W. Straquadine, X. Wang, D. Luo, M. E. Kozina, A. H. Reid, R. Li, J. Yang, S. P. Weathersby, S. Park, E. J. Sie, P. Jarillo-Herrero, I. R. Fisher, X. Wang, E. Demler, and N. Gedik, Role of equilibrium fluctuations in light-induced order, Physical Review Letters 127, 227401 (2021).
- Pan and Caruso [2023] Y. Pan and F. Caruso, Vibrational dichroism of chiral valley phonons, Nano Lett. 23, 7463 (2023).
- Britt and Siwick [2023] T. L. Britt and B. J. Siwick, Ultrafast phonon diffuse scattering as a tool for observing chiral phonons in monolayer hexagonal lattices, Phys. Rev. B 107, 214306 (2023).
- Trigo [2018] M. Trigo, Ultrafast Fourier transform inelastic X-ray scattering, MRS Bull. 43, 520–526 (2018).
- Johnson et al. [2009] S. L. Johnson, P. Beaud, E. Vorobeva, C. J. Milne, É. D. Murray, S. Fahy, and G. Ingold, Directly observing squeezed phonon states with femtosecond x-ray diffraction, Phys. Rev. Lett. 102, 175503 (2009).
- Trigo et al. [2013] M. Trigo, M. Fuchs, J. Chen, M. P. Jiang, M. Cammarata, S. Fahy, D. M. Fritz, K. Gaffney, S. Ghimire, A. Higginbotham, S. L. Johnson, M. E. Kozina, J. Larsson, H. Lemke, A. M. Lindenberg, G. Ndabashimiye, F. Quirin, K. Sokolowski-Tinten, C. Uher, G. Wang, J. S. Wark, D. Zhu, and D. A. Reis, Fourier-transform inelastic X-ray scattering from time- and momentum-dependent phonon–phonon correlations, Nat. Phys. 9, 790–794 (2013).
- Hu and Nori [1996] X. Hu and F. Nori, Squeezed phonon states: Modulating quantum fluctuations of atomic displacements, Phys. Rev. Lett. 76, 2294–2297 (1996).
- Garrett et al. [1997] G. A. Garrett, A. G. Rojo, A. K. Sood, J. F. Whitaker, and R. Merlin, Vacuum squeezing of solids: Macroscopic quantum states driven by light pulses, Science 275, 1638–1640 (1997).
- Henighan et al. [2016] T. Henighan, M. Trigo, M. Chollet, J. N. Clark, S. Fahy, J. M. Glownia, M. P. Jiang, M. Kozina, H. Liu, S. Song, D. Zhu, and D. A. Reis, Control of two-phonon correlations and the mechanism of high-wavevector phonon generation by ultrafast light pulses, Phys. Rev. B 94, 020302 (2016).
- Dean et al. [2016] M. P. M. Dean, Y. Cao, X. Liu, S. Wall, D. Zhu, R. Mankowsky, V. Thampy, X. M. Chen, J. G. Vale, D. Casa, J. Kim, A. H. Said, P. Juhas, R. Alonso-Mori, J. M. Glownia, A. Robert, J. Robinson, M. Sikorski, S. Song, M. Kozina, H. Lemke, L. Patthey, S. Owada, T. Katayama, M. Yabashi, Y. Tanaka, T. Togashi, J. Liu, C. Rayan Serrao, B. J. Kim, L. Huber, C. L. Chang, D. F. McMorrow, M. Först, and J. P. Hill, Ultrafast energy- and momentum-resolved dynamics of magnetic correlations in the photo-doped Mott insulator Sr2IrO4, Nat. Mater. 15, 601 (2016).
- Mazzone et al. [2021] D. G. Mazzone, D. Meyers, Y. Cao, J. G. Vale, C. D. Dashwood, Y. Shi, A. J. A. James, N. J. Robinson, J. Lin, V. Thampy, Y. Tanaka, A. S. Johnson, H. Miao, R. Wang, T. A. Assefa, J. Kim, D. Casa, R. Mankowsky, D. Zhu, R. Alonso-Mori, S. Song, H. Yavas, T. Katayama, M. Yabashi, Y. Kubota, S. Owada, J. Liu, J. Yang, R. M. Konik, I. K. Robinson, J. P. Hill, D. F. McMorrow, M. Först, S. Wall, X. Liu, and M. P. M. Dean, Laser-induced transient magnons in Sr3Ir2O7 throughout the Brillouin zone, Proc. Natl. Acad. Sci. U.S.A. 118, e2103696118 (2021).
- Kim et al. [2018] J. Kim, D. Casa, A. Said, R. Krakora, B. J. Kim, E. Kasman, X. Huang, and T. Gog, Quartz-based flat-crystal resonant inelastic x-ray scattering spectrometer with sub-10 meV energy resolution, Sci. Rep. 8, 1958 (2018).
- Damascelli et al. [2003] A. Damascelli, Z. Hussain, and Z.-X. Shen, Angle-resolved photoemission studies of the cuprate superconductors, Rev. Mod. Phys. 75, 473 (2003).
- Sobota et al. [2021] J. A. Sobota, Y. He, and Z.-X. Shen, Angle-resolved photoemission studies of quantum materials, Rev. Mod. Phys. 93, 025006 (2021).
- Zong et al. [2019b] A. Zong, A. Kogar, Y.-Q. Bie, T. Rohwer, C. Lee, E. Baldini, E. Ergeçen, M. B. Yilmaz, B. Freelon, E. J. Sie, H. Zhou, J. Straquadine, P. Walmsley, P. E. Dolgirev, A. V. Rozhkov, I. R. Fisher, P. Jarillo-Herrero, B. V. Fine, and N. Gedik, Evidence for topological defects in a photoinduced phase transition, Nat. Phys. 15, 27 (2019b).
- Duan et al. [2023] S. Duan, W. Xia, C. Huang, S. Wang, L. Gu, H. Liu, D. Xiang, D. Qian, Y. Guo, and W. Zhang, Ultrafast switching from the charge density wave phase to a metastable metallic state in 1-TiSe2, Phys. Rev. Lett. 130, 226501 (2023).
- Madéo et al. [2020] J. Madéo, M. K. L. Man, C. Sahoo, M. Campbell, V. Pareek, E. L. Wong, A. Al-Mahboob, N. S. Chan, A. Karmakar, B. M. K. Mariserla, X. Li, T. F. Heinz, T. Cao, and K. M. Dani, Directly visualizing the momentum-forbidden dark excitons and their dynamics in atomically thin semiconductors, Science 370, 1199 (2020).
- Karni et al. [2022] O. Karni, E. Barré, V. Pareek, J. D. Georgaras, M. K. L. Man, C. Sahoo, D. R. Bacon, X. Zhu, H. B. Ribeiro, A. L. O’Beirne, J. Hu, A. Al-Mahboob, M. M. M. Abdelrasoul, N. S. Chan, A. Karmakar, A. J. Winchester, B. Kim, K. Watanabe, T. Taniguchi, K. Barmak, J. Madéo, F. H. da Jornada, T. F. Heinz, and K. M. Dani, Structure of the moiré exciton captured by imaging its electron and hole, Nature 603, 247–252 (2022).
- Rustagi and Kemper [2018] A. Rustagi and A. F. Kemper, Photoemission signature of excitons, Phys. Rev. B 97, 235310 (2018).
- Zewail and Thomas [2009] A. H. Zewail and J. M. Thomas, 4D electron microscopy: imaging in space and time (World Scientific, 2009).
- Zewail [2010] A. H. Zewail, Four-dimensional electron microscopy, Science 328, 187 (2010).
- Barwick et al. [2009] B. Barwick, D. J. Flannigan, and A. H. Zewail, Photon-induced near-field electron microscopy, Nature 462, 902 (2009).
- Liu et al. [2016] H. Liu, J. S. Baskin, and A. H. Zewail, Infrared PINEM developed by diffraction in 4D UEM, Proc. Natl. Acad. Sci. U.S.A. 113, 2041 (2016).
- Büttner et al. [2021] F. Büttner, B. Pfau, M. Böttcher, M. Schneider, G. Mercurio, C. M. Günther, P. Hessing, C. Klose, A. Wittmann, K. Gerlinger, L.-M. Kern, C. Strüber, C. von Korff Schmising, J. Fuchs, D. Engel, A. Churikova, S. Huang, D. Suzuki, I. Lemesh, M. Huang, L. Caretta, D. Weder, J. H. Gaida, M. Möller, T. R. Harvey, S. Zayko, K. Bagschik, R. Carley, L. Mercadier, J. Schlappa, A. Yaroslavtsev, L. Le Guyarder, N. Gerasimova, A. Scherz, C. Deiter, R. Gort, D. Hickin, J. Zhu, M. Turcato, D. Lomidze, F. Erdinger, A. Castoldi, S. Maffessanti, M. Porro, A. Samartsev, J. Sinova, C. Ropers, J. H. Mentink, B. Dupé, G. S. D. Beach, and S. Eisebitt, Observation of fluctuation-mediated picosecond nucleation of a topological phase, Nat. Mater. 20, 30 (2021).
- Klose et al. [2023] C. Klose, F. Büttner, W. Hu, C. Mazzoli, K. Litzius, R. Battistelli, I. Lemesh, J. M. Bartell, M. Huang, C. M. Günther, M. Schneider, A. Barbour, S. B. Wilkins, G. S. D. Beach, S. Eisebitt, and B. Pfau, Coherent correlation imaging for resolving fluctuating states of matter, Nature 614, 256 (2023).
- Johnson et al. [2023] A. S. Johnson, D. Perez-Salinas, K. M. Siddiqui, S. Kim, S. Choi, K. Volckaert, P. E. Majchrzak, S. Ulstrup, N. Agarwal, K. Hallman, R. F. Haglund, C. M. Günther, B. Pfau, S. Eisebitt, D. Backes, F. Maccherozzi, A. Fitzpatrick, S. S. Dhesi, P. Gargiani, M. Valvidares, N. Artrith, F. de Groot, H. Choi, D. Jang, A. Katoch, S. Kwon, S. H. Park, H. Kim, and S. E. Wall, Ultrafast X-ray imaging of the light-induced phase transition in VO2, Nat. Phys. 19, 215 (2023).
- Jelic et al. [2024] V. Jelic, S. Adams, M. Hassan, K. Cleland-Host, S. E. Ammerman, and T. L. Cocker, Atomic-scale terahertz time-domain spectroscopy, Nat. Photonics 18, 898–904 (2024).
- Sheng et al. [2024] S. Sheng, M. Abdo, S. Rolf-Pissarczyk, K. Lichtenberg, S. Baumann, J. A. J. Burgess, L. Malavolti, and S. Loth, Terahertz spectroscopy of collective charge density wave dynamics at the atomic scale, Nat. Phys. 10.1038/s41567-024-02552-7 (2024).
- Barwick and Zewail [2015] B. Barwick and A. H. Zewail, Photonics and plasmonics in 4D ultrafast electron microscopy, ACS Photonics 2, 1391 (2015).
- Piazza et al. [2015] L. Piazza, T. T. A. Lummen, E. Quiñonez, Y. Murooka, B. W. Reed, B. Barwick, and F. Carbone, Simultaneous observation of the quantization and the interference pattern of a plasmonic near-field, Nat. Commun. 6, 6407 (2015).
- Kurman et al. [2021] Y. Kurman, R. Dahan, H. H. Sheinfux, K. Wang, M. Yannai, Y. Adiv, O. Reinhardt, L. H. G. Tizei, S. Y. Woo, J. Li, J. H. Edgar, M. Kociak, F. H. L. Koppens, and I. Kaminer, Spatiotemporal imaging of 2D polariton wave packet dynamics using free electrons, Science 372, 1181 (2021).
- Yurtsever et al. [2012] A. Yurtsever, R. M. van der Veen, and A. H. Zewail, Subparticle ultrafast spectrum imaging in 4D electron microscopy, Science 335, 59 (2012).
- Bakker et al. [1994] H. J. Bakker, S. Hunsche, and H. Kurz, Investigation of anharmonic lattice vibrations with coherent phonon polaritons, Phys. Rev. B 50, 914 (1994).
- Luo et al. [2024] T. Luo, B. Ilyas, A. v. Hoegen, Y. Lee, J. Park, J.-G. Park, and N. Gedik, Time-of-flight detection of terahertz phonon-polariton, Nat. Commun. 15, 2276 (2024).
- Feurer et al. [2007] T. Feurer, N. S. Stoyanov, D. W. Ward, J. C. Vaughan, E. R. Statz, and K. A. Nelson, Terahertz polaritonics, Annu. Rev. Mater. Res. 37, 317 (2007).
- Bakker et al. [1998] H. J. Bakker, S. Hunsche, and H. Kurz, Coherent phonon polaritons as probes of anharmonic phonons in ferroelectrics, Rev. Mod. Phys. 70, 523 (1998).
- Henry and Hopfield [1965] C. H. Henry and J. J. Hopfield, Raman scattering by polaritons, Phys. Rev. Lett. 15, 964 (1965).
- Cheung and Auston [1985] K. P. Cheung and D. H. Auston, Excitation of coherent phonon polaritons with femtosecond optical pulses, Phys. Rev. Lett. 55, 2152 (1985).
- Yan et al. [1985] Y. X. Yan, J. Gamble, E. B., and K. A. Nelson, Impulsive stimulated scattering: General importance in femtosecond laser pulse interactions with matter, and spectroscopic applications, J. Chem. Phys. 83, 5391 (1985).
- Planken et al. [1992] P. C. M. Planken, L. D. Noordam, J. T. M. Kennis, and A. Lagendijk, Femtosecond time-resolved study of the generation and propagation of phonon polaritons in LiNbO3, Phys. Rev. B 45, 7106 (1992).
- Koehl et al. [1999a] R. M. Koehl, S. Adachi, and K. A. Nelson, Direct visualization of collective wavepacket dynamics, J. Phys. Chem. A 103, 10260 (1999a).
- Koehl et al. [1999b] R. M. Koehl, S. Adachi, and K. A. Nelson, Real-space polariton wave packet imaging, J. Chem. Phys. 110, 1317–1320 (1999b).
- Kojima et al. [2003] S. Kojima, N. Tsumura, M. W. Takeda, and S. Nishizawa, Far-infrared phonon-polariton dispersion probed by terahertz time-domain spectroscopy, Phys. Rev. B 67, 035102 (2003).
- Alfaro-Mozaz et al. [2019] F. J. Alfaro-Mozaz, S. G. Rodrigo, P. Alonso-González, S. Vélez, I. Dolado, F. Casanova, L. E. Hueso, L. Martín-Moreno, R. Hillenbrand, and A. Y. Nikitin, Deeply subwavelength phonon-polaritonic crystal made of a van der Waals material, Nat. Commun. 10, 42 (2019).
- Foteinopoulou et al. [2019] S. Foteinopoulou, G. C. R. Devarapu, G. S. Subramania, S. Krishna, and D. Wasserman, Phonon-polaritonics: enabling powerful capabilities for infrared photonics, Nanophotonics 8, 232 (2019).
- Stuart [2004] B. H. Stuart, Infrared spectroscopy: fundamentals and applications (John Wiley & Sons, Ltd, 2004).
- Gardiner [1989] D. J. Gardiner, in Practical Raman Spectroscopy (Springer, 1989) p. 1.
- Burkel [2000] E. Burkel, Phonon spectroscopy by inelastic x-ray scattering, Rep. Prog. Phys. 63, 171 (2000).
- Kozina et al. [2019] M. Kozina, M. Fechner, P. Marsik, T. van Driel, J. M. Glownia, C. Bernhard, M. Radovic, D. Zhu, S. Bonetti, U. Staub, and M. C. Hoffmann, Terahertz-driven phonon upconversion in SrTiO3, Nat. Phys. 15, 387 (2019).
- Först et al. [2011] M. Först, C. Manzoni, S. Kaiser, Y. Tomioka, Y. Tokura, R. Merlin, and A. Cavalleri, Nonlinear phononics as an ultrafast route to lattice control, Nat. Phys. 7, 854 (2011).
- Radaelli [2018] P. G. Radaelli, Breaking symmetry with light: Ultrafast ferroelectricity and magnetism from three-phonon coupling, Phys. Rev. B 97, 085145 (2018).
- Hortensius et al. [2020] J. R. Hortensius, D. Afanasiev, A. Sasani, E. Bousquet, and A. D. Caviglia, Ultrafast strain engineering and coherent structural dynamics from resonantly driven optical phonons in LaAlO3, npj Quantum Mater. 5, 95 (2020).
- Mankowsky et al. [2017] R. Mankowsky, A. von Hoegen, M. Först, and A. Cavalleri, Ultrafast reversal of the ferroelectric polarization, Phys. Rev. Lett. 118, 197601 (2017).
- Henstridge et al. [2022] M. Henstridge, M. Först, E. Rowe, M. Fechner, and A. Cavalleri, Nonlocal nonlinear phononics, Nat. Phys. 18, 457 (2022).
- Li et al. [2019] X. Li, T. Qiu, J. Zhang, E. Baldini, J. Lu, A. M. Rappe, and K. A. Nelson, Terahertz field-induced ferroelectricity in quantum paraelectric SrTiO3, Science 364, 1079 (2019).
- Nova et al. [2019] T. F. Nova, A. S. Disa, M. Fechner, and A. Cavalleri, Metastable ferroelectricity in optically strained SrTiO3, Science 364, 1075 (2019).
- Fechner et al. [2024] M. Fechner, M. Först, G. Orenstein, V. Krapivin, A. S. Disa, M. Buzzi, A. von Hoegen, G. de la Pena, Q. L. Nguyen, R. Mankowsky, M. Sander, H. Lemke, Y. Deng, M. Trigo, and A. Cavalleri, Quenched lattice fluctuations in optically driven SrTiO3, Nat. Mater. 23, 363 (2024).
- Bustamante Lopez et al. [2023] D. A. Bustamante Lopez, D. M. Juraschek, M. Fechner, X. Xu, S.-W. Cheong, and W. Hu, Ultrafast simultaneous manipulation of multiple ferroic orders through nonlinear phonon excitation (2023), arXiv:2305.08250 .
- Fausti et al. [2011] D. Fausti, R. I. Tobey, N. Dean, S. Kaiser, A. Dienst, M. C. Hoffmann, S. Pyon, T. Takayama, H. Takagi, and A. Cavalleri, Light-induced superconductivity in a stripe-ordered cuprate, Science 331, 189 (2011).
- Mankowsky et al. [2014] R. Mankowsky, A. Subedi, M. Först, S. O. Mariager, M. Chollet, H. T. Lemke, J. S. Robinson, J. M. Glownia, M. P. Minitti, A. Frano, M. Fechner, N. A. Spaldin, T. Loew, B. Keimer, A. Georges, and A. Cavalleri, Nonlinear lattice dynamics as a basis for enhanced superconductivity in YBa2Cu3O6.5, Nature 516, 71–73 (2014).
- Mankowsky et al. [2017] R. Mankowsky, M. Fechner, M. Först, A. von Hoegen, J. Porras, T. Loew, G. L. Dakovski, M. Seaberg, S. Möller, G. Coslovich, B. Keimer, S. S. Dhesi, and A. Cavalleri, Optically induced lattice deformations, electronic structure changes, and enhanced superconductivity in YBa2Cu3O6.48, Struct. Dyn. 4, 044007 (2017).
- Kaiser [2017] S. Kaiser, Light-induced superconductivity in high- cuprates, Phys. Scr. 92, 103001 (2017).
- Liu et al. [2020] B. Liu, M. Först, M. Fechner, D. Nicoletti, J. Porras, T. Loew, B. Keimer, and A. Cavalleri, Pump frequency resonances for light-induced incipient superconductivity in YBa2Cu3O6.5, Phys. Rev. X 10, 011053 (2020).
- Fava et al. [2024] S. Fava, G. De Vecchi, G. Jotzu, M. Buzzi, T. Gebert, Y. Liu, B. Keimer, and A. Cavalleri, Magnetic field expulsion in optically driven YBa2Cu3O6.48, Nature 632, 75 (2024).
- Zhang et al. [2020] S. J. Zhang, Z. X. Wang, H. Xiang, X. Yao, Q. M. Liu, L. Y. Shi, T. Lin, T. Dong, D. Wu, and N. L. Wang, Photoinduced nonequilibrium response in underdoped YBa2Cu3O6+x probed by time-resolved terahertz spectroscopy, Phys. Rev. X 10, 011056 (2020).
- Zhang et al. [2024a] S. J. Zhang, X. Y. Zhou, S. X. Xu, Q. Wu, L. Yue, Q. M. Liu, T. C. Hu, R. S. Li, J. Y. Yuan, C. C. Homes, G. D. Gu, T. Dong, and N. L. Wang, Light-induced melting of competing stripe orders without introducing superconductivity in La2-xBaxCuO4, Phys. Rev. X 14, 011036 (2024a).
- Dodge et al. [2023a] J. S. Dodge, L. Lopez, and D. G. Sahota, Optical saturation produces spurious evidence for photoinduced superconductivity in K3C60, Phys. Rev. Lett. 130, 146002 (2023a).
- Dodge et al. [2023b] J. S. Dodge, L. Lopez, and D. G. Sahota, in 48th International Conference on Infrared, Millimeter, and Terahertz Waves (IRMMW-THz) (IEEE, 2023) p. 1.
- Buzzi et al. [2023] M. Buzzi, D. Nicoletti, E. Rowe, E. Wang, and A. Cavalleri, Comment on arXiv:2210.01114: Optical Saturation Produces Spurious Evidence for Photoinduced Superconductivity in K3C60 (2023), arXiv:2303.10169 .
- Blank et al. [2023] T. G. H. Blank, K. A. Grishunin, K. A. Zvezdin, N. T. Hai, J. C. Wu, S. H. Su, J. C. A. Huang, A. K. Zvezdin, and A. V. Kimel, Two-dimensional terahertz spectroscopy of nonlinear phononics in the topological Insulator MnBi2Te4, Phys. Rev. Lett. 131, 026902 (2023).
- Cheng et al. [2023] B. Cheng, P. L. Kramer, Z.-X. Shen, and M. C. Hoffmann, Terahertz-driven local dipolar correlation in a quantum paraelectric, Phys. Rev. Lett. 130, 126902 (2023).
- Cui et al. [2023] J. Cui, E. V. Boström, M. Ozerov, F. Wu, Q. Jiang, J.-H. Chu, C. Li, F. Liu, X. Xu, A. Rubio, and Q. Zhang, Chirality selective magnon-phonon hybridization and magnon-induced chiral phonons in a layered zigzag antiferromagnet, Nat. Commun. 14, 3396 (2023).
- Mertens et al. [2023] F. Mertens, D. Mönkebüscher, U. Parlak, C. Boix-Constant, S. Mañas-Valero, M. Matzer, R. Adhikari, A. Bonanni, E. Coronado, A. M. Kalashnikova, D. Bossini, and M. Cinchetti, Ultrafast coherent THz lattice dynamics coupled to spins in the van der Waals antiferromagnet FePS3, Adv. Mater. 35, 2208355 (2023).
- Nova et al. [2017] T. F. Nova, A. Cartella, A. Cantaluppi, M. Först, D. Bossini, R. V. Mikhaylovskiy, A. V. Kimel, R. Merlin, and A. Cavalleri, An effective magnetic field from optically driven phonons, Nat. Phys. 13, 132 (2017).
- Afanasiev et al. [2021] D. Afanasiev, J. R. Hortensius, B. A. Ivanov, A. Sasani, E. Bousquet, Y. M. Blanter, R. V. Mikhaylovskiy, A. V. Kimel, and A. D. Caviglia, Ultrafast control of magnetic interactions via light-driven phonons, Nat. Mater. 20, 607 (2021).
- Frenkel’ [1979] V. Y. Frenkel’, On the history of the Einstein-de Haas effect, Soviet Physics Uspekhi 22, 580 (1979).
- Barnett [1915] S. J. Barnett, Magnetization by rotation, Phys. Rev. 6, 239 (1915).
- Barnett [1935] S. J. Barnett, Gyromagnetic and electron-inertia effects, Rev. Mod. Phys 7, 129 (1935).
- Lee [1955] E. W. Lee, Magnetostriction and magnetomechanical effects, Rep. Prog. Phys. 18, 184 (1955).
- Dzialoshinskii [1958] I. E. Dzialoshinskii, The problem of piezomagnetism, Soviet Journal of Experimental and Theoretical Physics 6, 621 (1958).
- Borovik-romanov [1994] A. S. Borovik-romanov, Piezomagnetism, linear magnetostriction and magnetooptic effect, Ferroelectrics 162, 153 (1994).
- Broholm et al. [2020] C. Broholm, R. J. Cava, S. A. Kivelson, D. G. Nocera, M. R. Norman, and T. Senthil, Quantum spin liquids, Science 367, eaay0668 (2020).
- Hu et al. [2023] L. Hu, K.-z. Du, Y. Chen, Y. Zhai, X. Wang, and Q. Xiong, Spin-phonon coupling in two-dimensional magnetic materials, Natl. Sci. Open 2, 20230002 (2023).
- Beaurepaire et al. [1996] E. Beaurepaire, J.-C. Merle, A. Daunois, and J.-Y. Bigot, Ultrafast spin dynamics in ferromagnetic nickel, Phys. Rev. Lett. 76, 4250–4253 (1996).
- Dornes et al. [2019] C. Dornes, Y. Acremann, M. Savoini, M. Kubli, M. J. Neugebauer, E. Abreu, L. Huber, G. Lantz, C. A. F. Vaz, H. Lemke, E. M. Bothschafter, M. Porer, V. Esposito, L. Rettig, M. Buzzi, A. Alberca, Y. W. Windsor, P. Beaud, U. Staub, D. Zhu, S. Song, J. M. Glownia, and S. L. Johnson, The ultrafast Einstein-de Haas effect, Nature 565, 209 (2019).
- Tauchert et al. [2022] S. R. Tauchert, M. Volkov, D. Ehberger, D. Kazenwadel, M. Evers, H. Lange, A. Donges, A. Book, W. Kreuzpaintner, U. Nowak, and P. Baum, Polarized phonons carry angular momentum in ultrafast demagnetization, Nature 602, 73 (2022).
- Zhou et al. [2022] F. Zhou, K. Hwangbo, Q. Zhang, C. Wang, L. Shen, J. Zhang, Q. Jiang, A. Zong, Y. Su, M. Zajac, Y. Ahn, D. A. Walko, R. D. Schaller, J.-h. Chu, N. Gedik, X. Xu, D. Xiao, and H. Wen, Dynamical criticality of spin-shear coupling in van der Waals antiferromagnets, Nat. Commun. 13, 6598 (2022).
- Zhou et al. [2023] F. Zhou, H. Liu, M. Zajac, K. Hwangbo, Q. Jiang, J.-h. Chu, X. Xu, I. Arslan, T. E. Gage, and H. Wen, Ultrafast nanoimaging of spin-mediated shear waves in an acoustic cavity, Nano Lett. 23, 10213–10220 (2023).
- Liu et al. [2021] S. Liu, A. Granados del Águila, D. Bhowmick, C. K. Gan, T. Thu Ha Do, M. A. Prosnikov, D. Sedmidubský, Z. Sofer, P. C. M. Christianen, P. Sengupta, and Q. Xiong, Direct observation of magnon-phonon strong coupling in two-dimensional antiferromagnet at high magnetic fields, Phys. Rev. Lett. 127, 097401 (2021).
- Vaclavkova et al. [2021] D. Vaclavkova, M. Palit, J. Wyzula, S. Ghosh, A. Delhomme, S. Maity, P. Kapuscinski, A. Ghosh, M. Veis, M. Grzeszczyk, C. Faugeras, M. Orlita, S. Datta, and M. Potemski, Magnon polarons in the van der Waals antiferromagnet FePS3, Phys. Rev. B 104, 134437 (2021).
- Zhang et al. [2021b] Q. Zhang, M. Ozerov, E. V. Boström, J. Cui, N. Suri, Q. Jiang, C. Wang, F. Wu, K. Hwangbo, J.-H. Chu, D. Xiao, A. Rubio, and X. Xu, Coherent strong-coupling of terahertz magnons and phonons in a Van der Waals antiferromagnetic insulator (2021b), arXiv:2108.11619 .
- Mashkovich et al. [2021] E. A. Mashkovich, K. A. Grishunin, R. M. Dubrovin, A. K. Zvezdin, R. V. Pisarev, and A. V. Kimel, Terahertz light-driven coupling of antiferromagnetic spins to lattice, Science 374, 1608 (2021).
- Zhang and Niu [2015] L. Zhang and Q. Niu, Chiral phonons at high-symmetry points in monolayer hexagonal lattices, Phys. Rev. Lett. 115, 115502 (2015).
- Suri et al. [2021] N. Suri, C. Wang, Y. Zhang, and D. Xiao, Chiral phonons in Moiré superlattices, Nano Lett. 21, 10026 (2021).
- Zhu et al. [2018] H. Zhu, J. Yi, M.-Y. Li, J. Xiao, L. Zhang, C.-W. Yang, R. A. Kaindl, L.-J. Li, Y. Wang, and X. Zhang, Observation of chiral phonons, Science 359, 579 (2018).
- Kalashnikova et al. [2007] A. M. Kalashnikova, A. V. Kimel, R. V. Pisarev, V. N. Gridnev, A. Kirilyuk, and T. Rasing, Impulsive generation of coherent magnons by linearly polarized light in the easy-plane antiferromagnet , Phys. Rev. Lett. 99, 167205 (2007).
- Baldini et al. [2019] E. Baldini, A. Dominguez, T. Palmieri, O. Cannelli, A. Rubio, P. Ruello, and M. Chergui, Exciton control in a room temperature bulk semiconductor with coherent strain pulses, Sci. Adv. 5, eaax2937 (2019).
- Thouin et al. [2019] F. Thouin, D. A. Valverde-Chávez, C. Quarti, D. Cortecchia, I. Bargigia, D. Beljonne, A. Petrozza, C. Silva, and A. R. Srimath Kandada, Phonon coherences reveal the polaronic character of excitons in two-dimensional lead halide perovskites, Nat. Mater. 18, 349 (2019).
- Trovatello et al. [2020a] C. Trovatello, H. P. C. Miranda, A. Molina-Sánchez, R. Borrego Varillas, C. Manzoni, L. Moretti, L. Ganzer, M. Maiuri, J. Wang, D. Dumcenco, A. Kis, L. Wirtz, A. Marini, G. Soavi, A. C. Ferrari, G. Cerullo, D. Sangalli, and S. Dal Conte, Strongly coupled coherent phonons in single-layer MoS2, ACS nano 14, 5700 (2020a).
- Robert et al. [2016b] C. Robert, D. Lagarde, F. Cadiz, G. Wang, B. Lassagne, T. Amand, A. Balocchi, P. Renucci, S. Tongay, B. Urbaszek, and X. Marie, Exciton radiative lifetime in transition metal dichalcogenide monolayers, Phys. Rev. B 93, 205423 (2016b).
- Trovatello et al. [2020b] C. Trovatello, F. Katsch, N. J. Borys, M. Selig, K. Yao, R. Borrego-Varillas, F. Scotognella, I. Kriegel, A. Yan, A. Zettl, P. J. Schuck, A. Knorr, G. Cerullo, and S. D. Conte, The ultrafast onset of exciton formation in 2D semiconductors, Nat. Commun. , 5277 (2020b).
- Peng et al. [2022] R. Peng, A. Ripin, Y. Ye, J. Zhu, C. Wu, S. Lee, H. Li, T. Taniguchi, K. Watanabe, T. Cao, X. Xu, and M. Li, Long-range transport of 2D excitons with acoustic waves, Nat. Commun. 13, 1334 (2022).
- Morello et al. [2007] G. Morello, M. De Giorgi, S. Kudera, L. Manna, R. Cingolani, and M. Anni, Temperature and size dependence of nonradiative relaxation and exciton- phonon coupling in colloidal CdTe quantum dots, J. Phys. Chem. C 111, 5846 (2007).
- Sagar et al. [2008] D. M. Sagar, R. R. Cooney, S. L. Sewall, E. A. Dias, M. M. Barsan, I. S. Butler, and P. Kambhampati, Size dependent, state-resolved studies of exciton-phonon couplings in strongly confined semiconductor quantum dots, Phys. Rev. B 77, 235321 (2008).
- Nazir and McCutcheon [2016] A. Nazir and D. P. S. McCutcheon, Modelling exciton-phonon interactions in optically driven quantum dots, J. Phys. Condens. Matter 28, 103002 (2016).
- Zhang et al. [2024b] Z. Zhang, F. Y. Gao, J. B. Curtis, Z.-J. Liu, Y.-C. Chien, A. von Hoegen, M. T. Wong, T. Kurihara, T. Suemoto, P. Narang, E. Baldini, and K. A. Nelson, Terahertz field-induced nonlinear coupling of two magnon modes in an antiferromagnet, Nat. Phys. 20, 801 (2024b).
- Zhang et al. [2024c] Z. Zhang, F. Y. Gao, Y.-C. Chien, Z.-J. Liu, J. B. Curtis, E. R. Sung, X. Ma, W. Ren, S. Cao, P. Narang, A. von Hoegen, E. Baldini, and K. A. Nelson, Terahertz-field-driven magnon upconversion in an antiferromagnet, Nat. Phys. 20, 788 (2024c).
- Blank et al. [2023] T. G. H. Blank, K. A. Grishunin, B. A. Ivanov, E. A. Mashkovich, D. Afanasiev, and A. V. Kimel, Empowering control of antiferromagnets by THz-induced spin coherence, Physical Review Letters 131, 096701 (2023).
- Leenders et al. [2024] R. A. Leenders, D. Afanasiev, A. V. Kimel, and R. V. Mikhaylovskiy, Canted spin order as a platform for ultrafast conversion of magnons, Nature 630, 335–339 (2024).
- Bae et al. [2022] Y. J. Bae, J. Wang, A. Scheie, J. Xu, D. G. Chica, G. M. Diederich, J. Cenker, M. E. Ziebel, Y. Bai, H. Ren, C. R. Dean, M. Delor, X. Xu, X. Roy, A. D. Kent, and X. Zhu, Exciton-coupled coherent magnons in a 2D semiconductor, Nature 609, 282 (2022).
- Zayats et al. [2005] A. V. Zayats, I. I. Smolyaninov, and A. A. Maradudin, Nano-optics of surface plasmon polaritons, Phys. Rep. 408, 131 (2005).
- Shen and Bauer [2018] K. Shen and G. E. W. Bauer, Theory of spin and lattice wave dynamics excited by focused laser pulses, J. Phys. D Appl. Phys. 51, 224008 (2018).
- Ogawa et al. [2015] N. Ogawa, W. Koshibae, A. J. Beekman, N. Nagaosa, M. Kubota, M. Kawasaki, and Y. Tokura, Photodrive of magnetic bubbles via magnetoelastic waves, Proc. Natl. Acad. Sci. U.S.A. 112, 8977 (2015).
- Wang et al. [2023] H. Wang, Y. Yang, J. Chen, J. Wang, H. Jia, P. Chen, Y. Zhang, C. Wan, S. Liu, D. Yu, X. Han, J.-P. Ansermet, J. Zhang, and H. Yu, Long-distance coherent propagation of magnon polarons in a ferroelectric-ferromagnetic heterostructure, Phys. Rev. B 108, 144425 (2023).
- Bae et al. [2024] Y. J. Bae, T. Handa, Y. Dai, J. Wang, H. Liu, A. Scheie, D. G. Chica, M. E. Ziebel, A. D. Kent, X. Xu, K. Shen, X. Roy, and X. Zhu, Transient magnetoelastic coupling in CrSBr, Phys. Rev. B 109, 104401 (2024).
- Sun et al. [2024] Y. Sun, F. Meng, C. Lee, A. Soll, H. Zhang, R. Ramesh, J. Yao, Z. Sofer, and J. Orenstein, Dipolar spin wave packet transport in a van der Waals antiferromagnet, Nat. Phys. 20, 794 (2024).
- Violante et al. [2014] A. Violante, K. Cohen, S. Lazić, R. Hey, R. Rapaport, and P. V. Santos, Dynamics of indirect exciton transport by moving acoustic fields, New J. Phys. 16, 033035 (2014).
- Grasselli et al. [2018] F. Grasselli, A. Bertoni, and G. Goldoni, Classical and quantum dynamics of indirect excitons driven by surface acoustic waves, Phys. Rev. B 98, 165407 (2018).
- Zhang et al. [2012] J. Zhang, L. Zhang, and W. Xu, Surface plasmon polaritons: physics and applications, J. Phys. D 45, 113001 (2012).
- Latini et al. [2021] S. Latini, U. De Giovannini, E. J. Sie, N. Gedik, H. Hübener, and A. Rubio, Phonoritons as hybridized exciton-photon-phonon excitations in a monolayer h-BN optical cavity, Phys. Rev. Lett. 126, 227401 (2021).
- Zhang et al. [2020] H. Zhang, B. Abhiraman, Q. Zhang, J. Miao, K. Jo, S. Roccasecca, M. W. Knight, A. R. Davoyan, and D. Jariwala, Hybrid exciton-plasmon-polaritons in van der Waals semiconductor gratings, Nat. Commun. 11, 3552 (2020).
- Keldysh [2024] L. V. Keldysh, in Selected Papers of Leonid V Keldysh (World Scientific, 2024) p. 155.
- Tang et al. [2022] Y. Tang, Y. Zhang, Q. Liu, K. Wei, X. Cheng, L. Shi, and T. Jiang, Interacting plexcitons for designed ultrafast optical nonlinearity in a monolayer semiconductor, Light Sci. Appl. 11, 94 (2022).
- Deng et al. [2002] H. Deng, G. Weihs, C. Santori, J. Bloch, and Y. Yamamoto, Condensation of semiconductor microcavity exciton polaritons, Science 298, 199 (2002).
- Deng et al. [2010] H. Deng, H. Haug, and Y. Yamamoto, Exciton-polariton Bose-Einstein condensation, Rev. Mod. Phys. 82, 1489–1537 (2010).
- Luo et al. [2023b] S. Luo, H. Zhou, L. Zhang, and Z. Chen, Nanophotonics of microcavity exciton-polaritons, Appl. Phys. Rev. 10, 011316 (2023b).
- Ashida et al. [2020] Y. Ashida, A. Imamoǧlu, J. Faist, D. Jaksch, A. Cavalleri, and E. Demler, Quantum electrodynamic control of matter: Cavity-enhanced ferroelectric phase transition, Phys. Rev. X 10, 041027 (2020).
- Schlawin et al. [2022] F. Schlawin, D. M. Kennes, and M. A. Sentef, Cavity quantum materials, Appl. Phys. Rev. 9, 011312 (2022).
- Peng et al. [2024] K. Peng, W. Li, M. Sun, J. D. H. Rivero, C. Ti, X. Han, L. Ge, L. Yang, X. Zhang, and W. Bao, Topological valley Hall polariton condensation, Nat. Nanotechnol. 19, 1283–1289 (2024).
- Garcia-Vidal et al. [2021] F. J. Garcia-Vidal, C. Ciuti, and T. W. Ebbesen, Manipulating matter by strong coupling to vacuum fields, Science 373, eabd0336 (2021).
- Jarc et al. [2023] G. Jarc, S. Y. Mathengattil, A. Montanaro, F. Giusti, E. M. Rigoni, R. Sergo, F. Fassioli, S. Winnerl, S. Dal Zilio, D. Mihailovic, P. Prelovšek, M. Eckstein, and D. Fausti, Cavity-mediated thermal control of metal-to-insulator transition in 1-TaS2, Nature 622, 487–492 (2023).
- Zong et al. [2024] A. Zong, S.-C. Lin, S. A. Sato, E. Berger, B. R. Nebgen, M. Hui, B. Q. Lv, Y. Cheng, W. Xia, Y. Guo, D. Xiang, and M. W. Zuerch, Core-level signature of long-range density-wave order and short-range excitonic correlations probed by attosecond broadband spectroscopy (2024), arXiv:2407.00772 .
- O’Callahan et al. [2015] B. T. O’Callahan, A. C. Jones, J. Hyung Park, D. H. Cobden, J. M. Atkin, and M. B. Raschke, Inhomogeneity of the ultrafast insulator-to-metal transition dynamics of VO2, Nat. Commun. 6, 6849 (2015).
- Cremons et al. [2016] D. R. Cremons, D. A. Plemmons, and D. J. Flannigan, Femtosecond electron imaging of defect-modulated phonon dynamics, Nat. Commun. 7, 11230 (2016).
- Perez-Salinas et al. [2022] D. Perez-Salinas, A. S. Johnson, D. Prabhakaran, and S. Wall, Multi-mode excitation drives disorder during the ultrafast melting of a C4-symmetry-broken phase, Nat. Commun. 13, 238 (2022).
- Carbin et al. [2023] T. Carbin, X. Zhang, A. B. Culver, H. Zhao, A. Zong, R. Acharya, C. J. Abbamonte, R. Roy, G. Cao, and A. Kogar, Evidence for bootstrap percolation dynamics in a photoinduced phase transition, Phys. Rev. Lett. 130, 186902 (2023).
- McClellan et al. [2024] J. McClellan, A. Zong, K. H. Pham, H. Liu, Z. W. B. Iton, B. Guzelturk, D. A. Walko, H. Wen, S. K. Cushing, and M. W. Zuerch, Hidden correlations in stochastic photoinduced dynamics of a solid-state electrolyte (2024), arXiv:2406.06832 .
- Kim et al. [2020] H. W. Kim, N. A. Vinokurov, I. H. Baek, K. Y. Oang, M. H. Kim, Y. C. Kim, K.-H. Jang, K. Lee, S. H. Park, S. Park, J. Shin, J. Kim, F. Rotermund, S. Cho, T. Feurer, and Y. U. Jeong, Towards jitter-free ultrafast electron diffraction technology, Nat. Photonics 14, 245–249 (2020).
- Nabben et al. [2023] D. Nabben, J. Kuttruff, L. Stolz, A. Ryabov, and P. Baum, Attosecond electron microscopy of sub-cycle optical dynamics, Nature 619, 63–67 (2023).
- Hui et al. [2024] D. Hui, H. Alqattan, M. Sennary, N. V. Golubev, and M. T. Hassan, Attosecond electron microscopy and diffraction, Sci. Adv. 10, eadp5805 (2024).
- René de Cotret et al. [2019] L. P. René de Cotret, J.-H. Pöhls, M. J. Stern, M. R. Otto, M. Sutton, and B. J. Siwick, Time- and momentum-resolved phonon population dynamics with ultrafast electron diffuse scattering, Phys. Rev. B 100, 214115 (2019).
- Britt et al. [2022] T. L. Britt, Q. Li, L. P. René de Cotret, N. Olsen, M. Otto, S. A. Hassan, M. Zacharias, F. Caruso, X. Zhu, and B. J. Siwick, Direct view of phonon dynamics in atomically thin MoS2, Nano Lett. 22, 4718 (2022).
- Sidiropoulos et al. [2021] T. P. H. Sidiropoulos, N. Di Palo, D. E. Rivas, S. Severino, M. Reduzzi, B. Nandy, B. Bauerhenne, S. Krylow, T. Vasileiadis, T. Danz, P. Elliott, S. Sharma, K. Dewhurst, C. Ropers, Y. Joly, M. E. Garcia, M. Wolf, R. Ernstorfer, and J. Biegert, Probing the energy conversion pathways between light, carriers, and lattice in real time with attosecond core-level spectroscopy, Phys. Rev. X 11, 041060 (2021).

### v2

@@
 Abstract
-Quantum materials hold immense promises for future applications due to their intriguing electronic, magnetic, thermal, and mechanical properties that are often traced to a complex interplay among different microscopic degrees of freedom.
+Quantum materials hold immense promises for future applications due to their intriguing electronic, magnetic, thermal, and mechanical properties that often arise from a complex interplay between microscopic degrees of freedom.
 Important insights of such interactions come from studying the collective excitations of electrons, spins, orbitals, and lattice, whose cooperative motions play a crucial role in determining the novel behavior of these systems and offer us a key tuning knob to modify material properties on-demand through external perturbations.
-In this regard, ultrafast light-matter interaction has shown great potential in controlling the couplings of collective excitations, and rapid progress in a plethora of time-resolved techniques down to the attosecond regime has significantly advanced our understanding of the coupling mechanism and guided us in manipulating the dynamical property of quantum materials.
-This review aims to highlight recent experiments on visualizing collective excitations in the time domain, focusing on the coupling mechanism between different collective modes such as phonon-phonon, phonon-magnon, phonon-exciton, magnon-magnon, magnon-exciton, and various polaritons.
+In this regard, ultrafast light-matter interaction has shown great potential in controlling the couplings of collective excitations, and rapid progress in a plethora of time-resolved techniques down to the attosecond regime has significantly advanced our understanding of the coupling mechanisms and guided us in manipulating the dynamical properties of quantum materials.
+This review aims to highlight recent experiments on visualizing collective excitations in the time domain, focusing on the coupling mechanisms between different collective modes such as phonon-phonon, phonon-magnon, phonon-exciton, magnon-magnon, magnon-exciton, and various polaritons.
 We introduce how these collective modes are excited by an ultrashort laser pulse and probed by different ultrafast techniques, and we explain how the coupling between collective excitations governs the ensuing nonequilibrium dynamics.
-We also provide some perspectives on future studies that can lead to further discoveries regarding the emergent properties of quantum materials both in and out of equilibrium.
+We also provide some perspectives on future studies that can lead to discoveries of the emergent properties of quantum materials both in and out of equilibrium.
 Contents
@@
 Quantum materials have garnered considerable interest due to their remarkable properties associated with, for instance, unconventional superconductivity 1, 2, 3, 4, multiferroicity 5, colossal magnetoresistance 6, and topological quantum states 7.
-These properties hold great promise for future applications 8 such as topological quantum computing 9 and spintronic devices 10, yet a precise understanding of the underlying mechanism remains elusive in many cases.
+These properties hold great promise for future applications 8 such as topological quantum computing 9 and spintronic devices 10, yet a precise understanding of the underlying mechanisms remains elusive in many cases.
 At the microscopic level, these properties emerge as a result of complex interactions among the charge, spin, orbital, and lattice degrees of freedom.
@@
 Unraveling the interplay among these degrees of freedom is crucial for advancing our knowledge of the rich phase diagrams of quantum systems.
-However, the key challenge has been the coexistence of multi-way couplings of a myriad of quasiparticles and excitations, which makes it difficult to isolate a specific type of interaction we are interested in.
+However, one key challenge has been the coexistence of multi-way couplings of a myriad of quasiparticles and excitations, which makes it difficult to isolate a specific type of interaction we are interested in.
 To tackle this challenge, recent advances in time-resolved techniques based on ultrafast light-matter interaction have identified time as a new dimension to investigate the couplings among the microscopic degrees of freedom.
@@
 One of the most notable examples is photoinduced coherent phonons 15, 16, where a macroscopic number of atoms move in perfect synchronization.
-Such lattice dynamics show up in a variety of time-resolved probes, such as the periodic oscillations of diffraction peak intensity in time-resolved scattering 17, 18, 19, 20, 21, 22 or the oscillation of optical reflectivity and transmissivity in transient optical spectroscopy 23, 24, 25.
-Oftentimes, more than one collective excitations emerge after a photoexcitation event, and their coupling is governed by the underlying interaction among the constituent degrees of freedom.
-Therefore, investigating these collective modes and their couplings provides an important perspective to study the microscopic interactions in quantum materials, which in turn offers an opportunity to use light to modify dynamical material properties.
+Such lattice dynamics show up in a variety of time-resolved probes, such as the periodic oscillations of diffraction peak intensity in time-resolved scattering 17, 18, 19, 20, 21, 22 and the oscillations of optical reflectivity and transmissivity in transient optical spectroscopy 23, 24, 25.
+Oftentimes, more than one type of collective excitation emerges after a photoexcitation event, and their coupling is governed by the underlying interaction among the constituent degrees of freedom.
+Investigating these collective modes and their couplings provides an important perspective to study the microscopic interactions in quantum materials, which in turn offers an opportunity to use light to modify dynamical material properties.
 This review focuses on recent experimental progress in understanding the couplings between collective excitations thanks to the development of ultrafast techniques with ever better spatiotemporal resolution and the capability to generate intense laser pulses with highly tailored properties.
@@
 In Sec.
-IV, we delve deeper into the couplings of collective excitations by introducing specific cases such as photon-phonon, phonon-phonon, phonon-magnon, phonon-exciton, among other coupling mechanisms.
+IV, we delve deeper into the couplings of collective excitations by introducing specific cases such as photon-phonon, phonon-phonon, phonon-magnon, phonon-exciton, and other coupling mechanisms.
 An outlook for future studies in this field is included in Sec.
 V.
-As this review does not aim to cover the more general area of ultrafast science in condensed matter systems yet an understanding of the broader field is helpful for contextualizing the present article, we hereby refer readers to a more general survey of theoretical advances and experimental techniques in ref.
+While this review does not aim to cover the more general area of ultrafast science in condensed matter systems, such understanding of the broader field is helpful for contextualizing the present article.
+Therefore, we refer readers to a more general survey of theoretical advances and experimental techniques in ref.
 26 as well as a few other topical reviews.
@@
 Regarding precision engineering of material properties through ultrafast light pulses, we refer readers to excellent reviews on nonequilibrium engineering such as nonlinear phononics 45, 46, 47 and Floquet engineering 48, 49.
-From the perspective of different material classes, readers can check out reviews on two-dimensional materials 50, magnetic materials 51, high-temperature superconductors 30, 46, complex oxides 52, and topological insulators 49.
+From the perspective of different material classes, see reviews on two-dimensional materials 50, magnetic materials 51, high-temperature superconductors 30, 46, complex oxides 52, and topological insulators 49.
 These lists are by no means exhaustive but provide sufficient conceptual background for appreciating the experiments highlighted in this review.
@@
 For example, structural or magnetic transitions are often accompanied by soft phonons or magnons, respectively, which signify an instability in the system.
-In addition, these collective modes are highly responsive to external stimuli, making it possible to achieve on-demand properties of quantum materials as one applies electromagnetic fields, pressure, strain, or ultrafast photoexcitation to modify the collective dynamics 58.
+In addition, these collective modes are highly responsive to external stimuli, making it possible to achieve on-demand properties of quantum materials as one applies electric or magnetic fields, pressure, strain, or ultrafast photoexcitation to modify the collective dynamics 58.
 While collective excitations are readily detected in an equilibrium setting and have hallmarks in a variety of measurements ranging from transport to scattering experiments, the underlying microscopic motions are incoherent and hence only time-averaged dynamics are obtained.
@@
 Under the paradigm of pump-probe measurements, coherent dynamics can be excited by an ultrashort laser pulse, offering a general framework for inducing and controlling coherent dynamics in solids.
-For example, coherent phonons or magnons are often observed following a photoexcitation event, where ionic displacement or spin precession occurs in complete synchrony with the same phase over a macroscopic sample volume.
+For example, coherent phonons or magnons are often observed following a photoexcitation event, where ionic displacement or spin precession occurs in complete synchrony with the same temporal phase over a macroscopic sample volume.
 Such coherent motions are in stark contrast to the random movements that account for incoherent phonons and magnons in thermal equilibrium.
@@
 One extensively studied example is transitions that involves a symmetry change under the Ginzburg-Landau framework, such as those found in material systems hosting a charge density wave (CDW).
-One type of collective excitations of the CDW ground state is the amplitude mode, which has been widely observed following above-gap photoexcitation that transiently suppresses the order parameter amplitude and hence launches a coherent amplitude oscillation 59, 60, 61.
-Importantly, the photo-driven amplitude mode, in particular its magnitude and frequency, can inform us how the phase transition unfolds after the laser pulse incidence, such as whether the phase change proceeds in a spatially homogeneous fashion or heterogeneity like topological defects is involved 62, 63.
+One type of collective excitations of the CDW ground state is the amplitude mode, which has been widely observed following an above-gap photoexcitation that transiently suppresses the order parameter amplitude and hence launches a coherent amplitude oscillation 59, 60, 61.
+Importantly, the photo-driven amplitude mode, in particular its magnitude and frequency, can inform us how the phase transition unfolds after the laser pulse incidence, such as whether the phase change proceeds in a spatially homogeneous fashion or involves heterogeneity like the formation of topological defects62, 63.
 As another example, in a prototypical CDW material LaTe3, the CDW order parameter can be suddenly quenched within a time set by the characteristic time of the amplitude mode oscillations 23.
@@
 Novel metastable states and phenomenology can emerge, such as dynamic multiferroicity 68, 69, 70, 71, a change in electronic topology 72, 73, and magnetism induced by light-driven chiral phonons 74, 75, 76 or as a result of strong spin-lattice coupling 77.
-We again take CDW materials as an example.
+We again use a CDW material as an example.
 In metallic 3-Ta1+xSe2, when its CDW amplitude mode is excited by an intense terahertz pulse, an energy gap appears in the optical conductivity spectrum and executes a dynamical evolution in sync with the amplitude mode oscillation, suggesting a transition into a transient insulating state that has no equilibrium counterpart 78.
 II.3 Dressed states
-Aside from the role of collective modes during the course of a nonequilibrium phase transition, they can also help induce novel dynamical states of matter via coherent light-matter interactions.
-For instance, whereas Floquet-Bloch states have been conventionally realized by hybridization between electronic states and an external periodic light field 79, 80, 81, 82, 83, 84, 85, there are a number of theoretical proposals for the Floquet states achieved by replacing photons with phonons 86, 87 or excitons 88.
-One signature of Floquet-Bloch states is the appearance of electronic band replicas detected in tr-ARPES experiments, where replica bands follow specific intensity relations and polarization selection rules of the exciting photons.
+Collective modes, aside from their role during the course of a nonequilibrium phase transition, can also help induce novel dynamical states of matter via coherent light-matter interactions.
+For instance, whereas Floquet-Bloch states have been conventionally realized by hybridization between electronic states and an external periodic light field 79, 80, 81, 82, 83, 84, 85, there are a number of theoretical proposals for achieving the Floquet states by replacing photons with phonons 86, 87 or excitons 88.
+One signature of Floquet-Bloch states is the appearance of electronic band replicas detected in tr-ARPES experiments, where these replica bands follow specific intensity relations and polarization selection rules given by the exciting photons.
 Importantly, these bands are spectrally spaced apart by , where is the angular frequency of the bosonic field, be it a photon, phonon, or exciton.
-In a recent photoemission experiments using a momentum microscope that was specifically designed to study very small samples, exciton-induced Floquet effects were reported in monolayer WS2 in the presence of a high density of resonantly-driven photoexcited excitons 89, which also bear the hallmark of an excitonic insulator phase.
+In a recent photoemission experiment using a momentum microscope specifically designed to study extremely small samples, exciton-induced Floquet effects were reported in monolayer WS2 in the presence of a high density of resonantly-driven photoexcited excitons 89, which also bear the hallmark of an excitonic insulator phase.
 II.4 Phase resolution of coherent collective excitations
@@
 Therefore, the phase of the collective excitation oscillations can be determined (labeled as in Fig.
-1b, bottom panel), which is important information that equilibrium experiments cannot measure directly.
+1b, bottom panel), which equilibrium experiments cannot measure directly.
 Phase-sensitivity can be used to isolate a single collective mode that affects multiple degrees of freedom, providing a way to measure the coupling strength between them 19, 22.
-More specifically, by selectively inducing a particular oscillatory response using a tailored pump, and by leveraging phase-locked coherent features from different observables for different degrees of freedom, one can obtain the coupling strength directly, independent of the theoretical model.
+More specifically, by selectively inducing a particular oscillatory response using a tailored pump, and by leveraging phase-locked coherent features from different observables for the corresponding degrees of freedom, one can obtain the coupling strength directly, independent of the theoretical model.
 For example, following the excitation of a coherent phonon, tr-ARPES can probe spectral shifts in the electronic band energy while time-resolved X-ray diffraction (tr-XRD) can quantify the shifts in each atomic position for the phonon of interest.
@@
 22).
-The extracted electron-phonon coupling strength in Bi2Se3 and Bi2Te3 generally conforms to the expectation from density functional theory (DFT) computations while that in FeSe shows a significant deviation from DFT, which underestimates the strength.
-Instead, after electron-electron correlation effects were included in a self-consistent DFT-dynamical mean field theory, a quantitative agreement emerged, highlighting an enhancement of the coupling strength due to strong electron correlations.
+The extracted electron-phonon coupling strength in Bi2Se3 and Bi2Te3 is generally consistent with the expectation from density functional theory (DFT) computations while that in FeSe shows a significant deviation from DFT, which underestimates the strength.
+Instead, after electron-electron correlation effects were included in a self-consistent DFT-dynamical mean field theory, a quantitative agreement emerged, indicating an enhancement of the coupling strength due to strong electron correlations.
 Besides electron-phonon coupling, this multi-probe experimental protocol has also been deployed to measure the dynamical magnetoelectric coupling in multiferroic materials.
-For instance, in NiI2, a giant coupling strength was obtained 90 where a single electromagnon mode induces periodic modulations in both the dipolar and magnetic orders in NiI2, revealing an interesting /2 phase shift and lock-in behavior between electric polarization and magnetization oscillations that are at terahertz frequencies.
-Even with a single experimental probe, a changing oscillation phase with respect to some tuning parameter can be used to reveal material properties 91, 92, 93 and to identify microscopic processes underlying the collective excitation 94.
+For instance, in NiI2, a giant coupling strength was obtained 90 where a single electromagnon mode induces periodic modulations in the dipolar and magnetic orders, revealing an interesting /2 phase shift and lock-in behavior between electric polarization and magnetization oscillations that are at terahertz frequencies.
+Even with a single experimental probe, a change in the oscillation phase with respect to some tuning parameter can be used to reveal material properties 91, 92, 93 and identify microscopic processes underlying the collective excitation 94.
 For instance, far above the Curie temperature, the optical phonon mode in a layered ferromagnetic insulator CrSiTe3 can be excited due to the conventional displacive mechanism 15, but the coherent phonon phase experiences a shift when the material cools below a temperature scale where in-plane ferromagnetic correlations start to develop 92.
-Such a phase shift was only observed when the excitation photon energy is sufficiently large to induce a charge transfer from the Te2- to Cr3+ ions, a process that also increases the superexchange energy.
-The phase shift was interpreted as evidence for a coherent phonon mechanism due to superexchange enhancement via magneto-elastic coupling, which only applies in the presence of short-range ferromagnetic correlation in this material.
-Notably, the phase shift of an mode across the Néel temperature also appears in an antiferromagnetic insulator FePS3, providing evidence for a selective coupling between the magnetic order to the trigonal distortions in this layered system 93.
-In another example of a coherent acoustic phonon in a freestanding film of La2/3Ca1/3MnO3, three different observables in time-resolved diffraction (Bragg peak intensity, width, and position) were observed to synchronously oscillate with a particular phase relation — either in-phase or out-of-phase — after photoexcitation 94.
-By theoretically modeling the phase relation across the distinct observables from this coherent acoustic mode, Su and coworkers interpreted it as features of ultrafast wrinkle formation that results from delamination between the freestanding film and the substrate at its boundary, offering unique insights into the oft-neglected film–substrate interaction that governs the properties of freestanding structures.
+Such a phase shift is only observed when the excitation photon energy is sufficiently large to induce charge transfer from the Te2- to Cr3+ ions, a process that also increases the superexchange energy.
+The phase shift is interpreted as evidence for a coherent phonon mechanism due to superexchange enhancement via magneto-elastic coupling, which only applies in the presence of short-range ferromagnetic correlation in this material.
+Notably, the phase shift of an mode across the Néel temperature also appears in an antiferromagnetic insulator FePS3, providing evidence for selective coupling between the magnetic order and the trigonal distortions in this layered system 93.
+Another example that highlights the importance of the oscillation phase comes from a coherent acoustic phonon in a freestanding film of La2/3Ca1/3MnO3.
+Three different observables in time-resolved diffraction (Bragg peak intensity, width, and position) were observed to synchronously oscillate with a particular phase relation — either in-phase or out-of-phase — after photoexcitation 94.
+By theoretically modeling the phase relation across the distinct observables from this coherent acoustic mode, Su and coworkers interpreted the phase relation as features of ultrafast wrinkle formation that results from delamination between the freestanding film and the substrate at its boundary, offering unique insights into the oft-neglected film–substrate interaction that governs the properties of freestanding structures.
 III Time-domain probes of collective excitations
-To explore collective excitations in the time domain and to study their couplings, a variety of advanced ultrafast tools have been developed, each with distinct advantages.
-Here, we briefly introduce a few selected techniques that serve as the workhorse in the experiments we reviewed in this article.
+To explore collective excitations and their couplings in the time domain, a variety of advanced ultrafast tools have been developed, each with distinct advantages.
+Here, we briefly introduce a few selected techniques that serve as the workhorses in the experiments we reviewed in this article.
 A schematic for each category of techniques is depicted in Fig.
 1b–e.
-For a typical table-top time-resolved probe introduced below, a femtosecond laser pulse is split into two beams, one of which is used to generate a tailored pump pulse to drive the nonequilibrium dynamics while the other is used to produce a probe pulse, which is either a photon or electron pulse.
+For a typical table-top time-resolved probe introduced below, a femtosecond laser pulse is split into two beams, one of which is used to generate a tailored pump pulse to drive the nonequilibrium dynamics while the other is used to produce a probe pulse, which is either a photon or an electron pulse.
 The frequency of the pump pulse can range from sub-terahertz up to ultraviolet and its polarization can be tuned anywhere between linear and circular, depending on the target excitation.
@@
 Without a net displacement of ions, the pump light can couple to the lattice to the second order in the phonon amplitude, where pairs of phonons at are generated to conserve the total momentum, leading to oscillations of the variance in the atomic displacement at twice the frequency of the phonon mode itself 118.
-This second-order effect has been demonstrated as intensity oscillations in either Bragg peaks 119 or diffuse scattering 120, and is also known as squeezed phonons 121, 122, 120, 123.
+This second-order effect induces intensity oscillations in either Bragg peaks 119 or diffuse scattering 120, and is also known as squeezed phonons 121, 122, 120, 123.
 Another rapidly developing area that rides on recent progress in the X-ray free-electron laser technology is time-resolved resonant inelastic X-ray scattering (RIXS), which not only reveals phonon dynamics but can also access charge, orbital, and spin excitations as well as high-order, multi-particle correlation functions that describe coupled spin/charge/orbital excitations 124, 38, 125, 39.
-With ongoing effort on entering the sub-10-meV energy resolution in RIXS 126, this emerging technique provides a unique window into the nonequilibrium states in correlated materials through dynamical signatures across time, momentum, and energy axes.
+With ongoing effort towards entering the sub-10-meV energy resolution in RIXS 126, this emerging technique provides a unique window into the nonequilibrium states in correlated materials through dynamical signatures across time, momentum, and energy axes.
 III.3 Time-resolved photoemission
@@
 By introducing a pump pulse, tr-ARPES yields further insights into nonequilibrium carrier dynamics and unoccupied states, reveals the electronic origin of ultrafast phase transitions 129, 130, and unambiguously demonstrates how band structure can be engineered by light 79, 80, 48, 49, 82, 81, 89.
-The advent of momentum microscope in tr-ARPES experiments has also recently made it feasible to investigate micrometer-sized samples that were challenging to access before.
+The advent of momentum microscopes in tr-ARPES experiments has also recently made it feasible to investigate micrometer-sized samples that were challenging to access before.
 For example, tr-ARPES experiments 131, 132, 89 have demonstrated the ability to probe dark excitons and moiré excitons in monolayer and heterobilayer transition metal dichalcogenides, respectively, whereas they were hard to probe by traditional photon-absorption or photon-emission spectroscopy.
 As illustrated in the bottom panel of Fig.
-1d, only the occupied state at the valence band top can be visualized before the optical pump.
+1d, only the occupied states at the valence band top can be visualized before the optical pump.
 With resonant photoexcitation at an exciton energy, photoemission signals from excitons can be observed upon the arrival of the pump pulses 131.
-The exciton photoemission signals feature an energy-momentum distribution centered below the conduction band bottom with a gap of binding energy (green dispersion line), conforming to theoretical expectations 133.
+The exciton photoemission signals feature an energy-momentum distribution centered below the conduction band bottom with a gap of binding energy (green dispersion line), in agreement with theoretical expectations 133, 134, 135, 136, 137.
 III.4 Time- and space-resolved probes
 Time-resolved spectroscopic and diffraction methods introduced thus far can only provide spatially-averaged information inside the probed region, whereas in general ultrafast dynamics may vary in different regions due to different local environments.
-Therefore, techniques with both spatial and temporal resolutions play an important role in studying inhomogeneous effects and near-field distribution during ultrafast light-matter interaction.
-Techniques such as ultrafast electron microscopy 134, 135, photoinduced near-field electron microscopy (PINEM) 136, 137, coherent correlated X-ray imaging 138, 139, 140, and lightwave-driven terahertz scanning tunneling microscopy 141, 142 have all shown a great potential in studying heterogeneity and spatially-varying dynamics present in quantum materials.
-As an example to illustrate the relation to collective excitations discussed in this article, ultrafast PINEM exhibits exceptional capabilities in studying the evolution of light-induced plasmonic field distribution 143 (see Fig.
+Therefore, techniques with both spatial and temporal resolutions play an important role in studying inhomogeneous effects and near-field distribution during ultrafast light-matter interactions.
+Techniques such as ultrafast electron microscopy 138, 139, photoinduced near-field electron microscopy (PINEM) 140, 141, coherent correlated X-ray imaging 142, 143, 144, and lightwave-driven terahertz scanning tunneling microscopy 145, 146 have all shown great potential in studying heterogeneity and spatially-varying dynamics present in quantum materials.
+As an example to illustrate the relation to collective excitations discussed in this article, ultrafast PINEM exhibits exceptional capabilities in studying the evolution of light-induced plasmonic field distribution 147 (see Fig.
 1e), and we will return to specific examples in Sec.
@@
 In the PINEM mode, inelastically scattered electrons are used instead, where electrons can absorb and emit spatiotemporally overlapped photons and produce discrete peaks with a constant energy interval in the electron energy loss spectrum.
-Using the energy-filtered electrons for imaging, one can obtain information about dressed photons such as the near-field distribution of surface plasmon-polaritons 144 and propagation of phonon-polaritons 145.
-By focusing pump and probe pulses at distinct points and spatially scanning the probe electron pulse, the near-field distributions with space, time, and energy resolution can also be obtained 146.
+Using the energy-filtered electrons for imaging, one can obtain information about dressed photons such as the near-field distribution of surface plasmon-polaritons 148 and propagation of phonon-polaritons 149.
+By focusing pump and probe pulses at distinct points and spatially scanning the probe electron pulse, the near-field distributions with space, time, and energy resolution can also be obtained 150.
 The techniques introduced in this section are by no means exhaustive, and here we refer readers to other excellent topical reviews on different classes of ultrafast probes 28, 40, 29, 30, 41, 42, 36, 37, 38, 39, 31, 33, 43, 34, 25, 32, 35, 44.
-IV Couplings between collective excitations at non-equilibrium
+IV Couplings between collective excitations at nonequilibrium
 IV.1 Dressing photons with phonons: phonon-polariton
@@
 1).
-One hallmark feature of these interactions is the appearance of an avoided crossing between the dispersion of two quasiparticles 58, as illustrated in Fig.
+One hallmark feature of these interactions is the appearance of an avoided crossing between the dispersions of two quasiparticles 58, as illustrated in Fig.
 2a.
 The size of the gap is a measure of the interaction strength, where a larger gap suggests a stronger interaction.
-Here we focus on recent ultrafast studies of phonon-polaritons, which arise from the interaction between lattice vibration and electromagnetic waves.
-When light passes through a crystal, it can drive a transverse optical phonon mode if the frequency of the electromagnetic wave resonates with the phonon, which leads to dipole oscillations and re-radiates the electromagnetic wave.
-The re-radiated waves can further excite the dipolar vibrations, resulting in the propagation of hybridized light and phonon, namely, phonon-polaritons.
-As the hybridized quasiparticles inherit the attributes from both light and phonons, their frequencies lie in the frequency range between gigahertz and tens of terahertz, showing great potential for applications in THz spectroscopy, THz imaging, and high-bandwidth signal processing 149.
-Many early ultrafast experiments of phonon-polaritons focused on the study of ferroelectrics 150, in which photoexcited coherent phonon-polaritons were used as a probe for low-frequency anharmonic phonons that play a crucial role in driving structural and ferroelectric phase transition near the critical point.
-Raman spectroscopy was first used to measure the frequency of phonon-polaritons 151 but lacked phase information of their wavepacket.
-With the development of femtosecond laser, ultrafast optical spectroscopy with spatially distinct pump and probe techniques has enabled us to directly measure the phonon-polaritons waveform in the time domain.
-In one experimental scheme, broadband polar phonon-polariton modes are produced by illuminating a femtosecond laser pulse onto the surface of a ferroelectric material, where the generation mechanism is based on impulsive stimulated Raman scattering (ISRS) through mixing different frequency components contained within the pulse bandwidth 152, 153.
-However, this method does not generate phonon-polariton mode with a well-defined frequency.
-To solve this problem, a four-wave mixing method has been developed 154, 147, 155, 156.
+Here we focus on recent ultrafast studies of phonon-polaritons, which arise from the interaction between lattice vibrations and electromagnetic waves.
+When light passes through a crystal, it can drive a transverse optical phonon mode if the frequency of the electromagnetic wave resonates with the phonon, which leads to dipole oscillations and consequent re-radiation of the electromagnetic wave.
+The re-radiated waves can further excite the dipolar vibrations, resulting in the propagation of hybridized light and phonons, namely, phonon-polaritons.
+As the hybridized quasiparticles inherit the attributes from both light and phonons, their frequencies lie in the frequency range between gigahertz and tens of terahertz, showing great potential for applications in THz spectroscopy, THz imaging, and high-bandwidth signal processing 153.
+Many early ultrafast experiments of phonon-polaritons focused on the study of ferroelectrics 154, in which photoexcited coherent phonon-polaritons were used as a probe for low-frequency anharmonic phonons that play a crucial role in driving structural and ferroelectric phase transition near the critical point.
+Raman spectroscopy was first used to measure the frequency of phonon-polaritons 155 but lacked phase information of their wavepacket.
+With the development of femtosecond lasers, ultrafast optical spectroscopy with spatially distinct pump and probe techniques has enabled us to directly measure the phonon-polaritons waveform in the time domain.
+In one experimental scheme, broadband polar phonon-polariton modes are produced by illuminating the surface of a ferroelectric material with a femtosecond laser pulse.
+These modes are generated from impulsive stimulated Raman scattering (ISRS) through mixing different frequency components contained within the pulse bandwidth 156, 157.
+However, this method does not generate phonon-polariton modes with a well-defined frequency.
+To solve this problem, a four-wave mixing method has been developed 158, 151, 159, 160.
 In this configuration, two identical, spatiotemporally overlapped femtosecond pulses are used to launch the phonon-polaritons in a nonlinear crystal.
@@
 As the photoexcited phonon-polariton field modulates the index of refraction that in turn affects the scattering of a separate probe pulse, the periodic probe signal can reveal the presence of phonon-polaritons in a phase-resolved manner.
-Figure 2b shows the measured dispersion relation of phonon-polaritons in LiNbO3 using this four-wave mixing scheme 147, where five branches and four avoided crossings were observed, showing strong couplings between the photons and phonons in this material.
-Experiments using the four-wave mixing methodology have some limitations: ISRS can only excite Raman-active modes 153, while to the lowest order only IR-active phonon modes can couple with electromagnetic waves to form phonon-polaritons.
-Therefore, the scheme is only feasible for noncentrosymmetric materials where IR phonons are also Raman-active, such as in LiNbO3 and LiTaO3.
+Figure 2b shows the measured dispersion relation of phonon-polaritons in LiNbO3 using this four-wave mixing scheme 151, where five branches and four avoided crossings were observed, showing strong couplings between the photons and phonons in this material.
+Experiments using four-wave mixing have some limitations: ISRS can only excite Raman-active modes 157, while to the lowest order only IR-active phonon modes can couple with electromagnetic waves to form phonon-polaritons.
+Therefore, this scheme is only feasible for noncentrosymmetric materials in which IR phonons are also Raman-active, such as in LiNbO3 and LiTaO3.
 In addition, these experiments need careful adjustment of the geometry of the two pump beams in order to excite phonon-polaritons with a selected wavevector.
-An alternative method that overcomes these limitations is time-domain THz spectroscopy 157, where phonon-polaritons are excited by a single beam of THz radiation instead of two beams of visible or near-infrared light pulses.
+An alternative method that overcomes these limitations is time-domain THz spectroscopy 161, where phonon-polaritons are excited by a single beam of THz radiation instead of two beams of visible or near-infrared light pulses.
 However, in a typical time-domain THz experiment that measures phonon-polariton dispersion, a reference crystal is required for accurate electro-optic sampling to retrieve the phase information across the spectrum of the THz pulse, hence complicating the data analysis and interpretation.
-Recently, Luo and coworkers developed a time-of-flight detection method that overcame these challenges and was applied to probing phonon-polaritons in van der Waals magnetic materials 148, including NiI2 and MnPS3.
-In this scheme, a terahertz pulse centered around 3 THz was used to excite phonon-polaritons while the SHG signal of a 800-nm laser pulse was used as a probe to capture phonon-polaritons before or after they are reflected from the back surface of the sample.
+Recently, Luo and coworkers developed a time-of-flight detection method that overcame these challenges and was applied to probing phonon-polaritons in van der Waals magnetic materials 152, including NiI2 and MnPS3.
+In this scheme, a pulse centered around 3 THz was used to excite phonon-polaritons while the SHG signal of a 800-nm laser pulse was used as a probe to capture phonon-polaritons before or after they were reflected from the back surface of the sample.
 More specifically, when the THz pump pulse arrived at the sample surface, multiple oscillatory responses due to phonons and magnons were observed (Fig.
@@
 2c, insets).
-After this revival, the signal frequency was also observed to increase towards a longer time delay (Fig.
+After this revival, the signal frequency was also observed to increase towards longer time delay (Fig.
 2c, windows B and C).
@@
 Around time zero, as defined by the THz pulse incidence, the excitation spectrum and the second harmonic frequencies of the THz pump can be seen, including the photoexcited and phonon modes.
-After 6 ps, a chirped mode emerges, whose frequency increases from 2 to 4 THz over time and it corresponds to the phonon-polariton wavepacket.
+After 6 ps, a chirped mode emerges, whose frequency increases from 2 to 4 THz over, which corresponds to the phonon-polariton wavepacket.
 Given the known sample thickness, the time-dependent frequency spectrum in Fig.
-2d can yield the group velocity at different frequencies, which in turn helps one reconstruct the energy-momentum dispersion of the phonon-polariton.
+2d can yield the group velocity at different frequencies, which in turn helps reconstruct the energy-momentum dispersion of the phonon-polariton.
 Furthermore, the time-dependent attenuation of the phonon-polariton during its propagation allows one to compute the imaginary part of the dispersion, whose divergence near the phonon resonance conforms to the expectation that light quickly dissipates its energy through the phonon mode.
 While the aforementioned experiments can detect the presence and dispersion of phonon-polariton wavepackets in the probed volume, they do not reveal the spatial structure of these hybridized quasiparticles.
-To image the spatiotemporal propagation of phonon-polaritons, ultrafast optical or electron microscopy was used.
-Nelson and colleagues first demonstrated direct imaging of optical phonon-polariton wavepackets in LiTaO3 using ultrafast optical microscopy that leveraged four-wave mixing 155, 156.
-In one of the experiments, an anisotropic spatial profile of the pump pulse was applied (see Fig.
-3a), which excited phonon-polaritons that propagated mainly along the horizontal direction of the sample 155.
+To image the spatiotemporal propagation of phonon-polaritons, ultrafast optical or electron microscopy can be used.
+Nelson and colleagues first demonstrated direct imaging of optical phonon-polariton wavepackets in LiTaO3 using ultrafast optical microscopy that leveraged four-wave mixing 159, 160.
+In one of the experiments, a pump pulse with an anisotropic spatial profile was applied (see Fig.
+3a), which excited phonon-polaritons that propagated mainly along the horizontal direction of the sample 159.
 The spot size of the probe pulse was significantly larger than the excitation region, so spatial information of the propagation dynamics was readily acquired.
 The left panel of Fig.
-3b shows a typical snapshot of the wavepacket distribution at 1335 fs after the pump incidence.
+3b shows a representative snapshot of the wavepacket distribution at 1335 fs after the pump incidence.
 After averaging over the vertical direction in each image captured at different time delays, the spatiotemporal propagation of phonon-polaritons in LaTiO3 was obtained in the right panel of Fig.
 3b.
-Under this experimental framework, the phonon-polariton response can be coherently controlled via spatiotemporal shaping of the optical pump pulses 149.
+Under this experimental framework, the phonon-polariton response can be coherently controlled via spatiotemporal shaping of the optical pump pulses 153.
 For a more comprehensive survey on the coherent control of phonon-polaritons and THz polaritonics, we refer readers to ref.
-149.
-Apart from optical probes, electrons have also been successfully deployed to image phonon-polariton propagation dynamics in van der Waals materials 145, leveraging the spatial and spectral information of PINEM.
+153.
+Apart from optical probes, electrons have been successfully deployed to image phonon-polariton propagation dynamics in van der Waals materials 149, leveraging the spatial and spectral information of PINEM.
 As shown in Fig.
-3c, an optical pump pulse first illuminates the edge of a hexagonal boron nitride (hBN) and launches multi-branch phonon-polariton wavepackets.
+3c, an optical pump pulse first illuminates the edge of a hexagonal boron nitride (-BN) and launches multi-branch phonon-polariton wavepackets.
 The probing free electrons interact with the propagating phonon-polaritons, broadening the electron energy loss profile due to inelastic scattering (Fig.
@@
 3e) as a function of distance from the edge (red arrow in the inset of Fig.
-3e), indicating the field profile of phonon-polariton along the propagation direction.
+3e), indicating the field profile of phonon-polaritons along the propagation direction.
 By studying the distance-time relation of the traveling wavepackets (Fig.
 3f), one can observe both acceleration (purple arrow) and deceleration (red arrow) during the propagation of the photoexcited phonon-polaritons.
-The observed acceleration and deceleration result from the excitation of different spectral components at different times during the pump-hBN interaction.
+The observed acceleration and deceleration result from the excitation of different spectral components at different times during the pump--BN interaction.
 Interestingly, Kurman et al.
-also showed that the accelerating and decelerating propagation can be modified by changing the exciting photon wavelength and sample thickness 145.
+also showed that the accelerating and decelerating propagation can be modified by changing the exciting photon wavelength and sample thickness 149.
 Note that the two featured examples of spatiotemporal imaging differ in the launching mechanism of the phonon-polaritons.
-While ultrafast optical microscopy used two near-infrared pulses to form a transient grating to excite phonon-polaritons based on the ISRS mechanism 155, the PINEM experiment directly used a mid-IR pulse to coherently excite them 145.
+While ultrafast optical microscopy used two near-infrared pulses to form a transient grating to excite phonon-polaritons based on the ISRS mechanism 159, the PINEM experiment directly used a mid-IR pulse to coherently excite them 149.
 Hence, the former methodology was deployed to noncentrosymmetric materials while the latter was used to directly excite dipole-active IR modes to launch the phonon-polaritons.
-In both cases, the additional spatial information has deepened our understanding of how phonon-polaritons propagate in materials, which will be especially helpful when one considers spatially heterogeneous samples in a real device setting 149, 158, 159.
+In both cases, the additional spatial information has deepened our understanding of how phonon-polaritons propagate in materials, which will be especially helpful when one considers spatially heterogeneous samples in a real device setting 153, 162, 163.
 IV.2 Phonon-phonon coupling
@@
 The collective excitations of atomic vibrations, or phonons, hence play an important role in determining these physical attributes as well.
-Under equilibrium conditions, phonons and their couplings can be characterized by techniques such as IR spectroscopy 160, Raman spectroscopy 161, and inelastic neutron, X-ray, or electron scatterings 162.
-Though these static methods are powerful, a lack of precise control over a particular phonon mode impedes us from revealing how individual phonon modes couple to other quasiparticles in solids.
+Under equilibrium conditions, phonons and their couplings can be characterized by techniques such as IR spectroscopy 164, Raman spectroscopy 165, and inelastic neutron, X-ray, or electron scatterings 166.
+Though these static methods are powerful, a lack of precise control over particular phonon modes impedes us from revealing how individual phonon modes couple to other quasiparticles in solids.
 In this regard, ultrafast techniques allow us to tackle this challenge by extending the measurement into the time domain.
-For example, advances in generating intense mid-IR and THz pulses with either broad band or tunable narrow-band frequencies bring the paradigm of nonlinear phononics 163, 47.
+For example, advances in generating intense mid-IR and THz pulses with either broadband or tunable narrow-band frequencies have been instrumental in the emergence of the field of nonlinear phononics 167, 47.
 By selectively exciting particular phonon modes using tailored low-photon-energy pulses, we can study phonon-phonon couplings and engineer crystal structure in a nonequilibrium state.
-The first experimental demonstration of nonlinear phononics was reported in 2011 by Först and coworkers 164.
-After the IR mode in a bulk single crystal of La0.7Sr0.3MnO3 was excited by 14.3-m light pulses, the reflectivity exhibited an oscillation at 1.2 THz, as shown by the dark curve in Fig.
+The first experimental demonstration of nonlinear phononics was reported in 2011 by Först and coworkers 167.
+After the IR mode in a bulk single crystal of La0.7Sr0.3MnO3 was excited by 14.3-m light pulses, the reflectivity exhibited oscillations at 1.2 THz, as shown by the dark curve in Fig.
 5a.
-The oscillation was attributed to the Raman mode that is associated with oxygen octahedral rotation in this perovskite.
+The oscillations were attributed to the Raman mode that is associated with oxygen octahedral rotation in this perovskite.
 Moreover, the amplitude of the mode showed a strong pump wavelength dependence and reached a maximum when it was resonant with the mode (see Fig.
@@
 For a more detailed description of this mechanism such as symmetry considerations during anharmonic interactions, we refer readers to refs.
-165, 47.
+168, 47.
 In contrast to the resonant excitation condition, 5.8-THz oscillations were observed when La0.7Sr0.3MnO3 was photoexcited by a 1.5-m pulse (Fig.
@@
 inspired ideas to engineer crystals using resonant mid-IR or terahertz excitations.
-This nonlinear phononic mechanism has since been utilized to manipulate properties in many perovskite systems, such as controlling an anisotropic strain wave in LaAlO3 166, as well as changing ferroelectric polarizations in systems such as LiNbO3 (refs.
-167, 168), SrTiO3 (refs.
-169, 170, 171), and BiFeO3 (ref.
-172).
-Besides the effect on the lattice and hence the coupled polar orders, nonlinear phononics have also been shown to influence electronic and magnetic properties 47.
-A particularly intriguing case concerns the apparent resemblance to a transient superconducting state when resonant excitation of IR-active phonons of specific wavelengths are driven to a large amplitude 173, 174, 175, 176, 177, 178.
-However, other works have suggested that certain superconducting-like behavior may be a manifestation of the excitation of quasiparticles with a low scattering rate 179, 180, and that care needs to be taken in processing the transient optical data for drawing precise conclusions about the nonequilibrium state 181, 182, 183.
+This nonlinear phononic mechanism has since been utilized to manipulate properties in many perovskite systems, such as controlling an anisotropic strain wave in LaAlO3 (ref.
+169), as well as changing ferroelectric polarizations in systems such as LiNbO3 (refs.
+170, 171), SrTiO3 (refs.
+172, 173, 174), and BiFeO3 (ref.
+175).
+Besides the effect on the lattice and hence coupled polar orders, nonlinear phononics has also been shown to influence electronic and magnetic properties 47.
+A particularly intriguing case concerns the apparent resemblance to a transient superconducting state when IR-active phonons of specific wavelengths are driven to a large amplitude by resonant excitation 176, 177, 178, 179, 180, 181.
+However, other works have suggested that certain superconducting-like behavior may be a manifestation of the excitation of quasiparticles with a low scattering rate 182, 183.
+Therefore, care needs to be taken in processing the transient optical data for drawing precise conclusions about the nonequilibrium state 184, 185, 186.
 This potential of light-induced superconductivity remains an intensely studied area with many interesting, open questions to be answered.
@@
 Here, we highlight several experimental works on SrTiO3, a quantum paraelectric material where the competition between ferroelectric and antiferrodistortive instabilities prevents it from entering a ferroelectric phase at low temperature.
-Using THz pump pulses, the direct control of its ferroelectric soft phonon mode has been intensely studied and several nonlinear phononic effects have been observed 169, 170, 171.
-For example, phonon up-conversion in SrTiO3 was demonstrated in a THz pump, X-ray diffraction probe experiment 163.
+Using THz pump pulses, the direct control of its ferroelectric soft phonon mode has been intensely studied and several nonlinear phononic effects have been observed 172, 173, 174.
+For example, phonon up-conversion in SrTiO3 was demonstrated in a THz pump, X-ray diffraction probe experiment 187.
 Kozina et al.
@@
 5c).
-Upon a temperature increase where the hardened TO1 mode falls outside the THz excitation spectrum, no signatures of the TO2 and TO3 modes are seen, providing compelling evidence for phonon frequency up-conversion from the soft TO1 mode to these high-frequency modes.
+When temperature increased such that the hardened TO1 mode fell outside the THz excitation spectrum, no signatures of the TO2 and TO3 modes were seen, providing compelling evidence for phonon frequency up-conversion from the soft TO1 mode to these high-frequency modes.
 Beyond the proof-of-principle demonstration by Kozina et al., nonlinear phononic effects have also been used to induce ferroelectricity in unstrained SrTiO3 under ambient pressure.
-In 2019, Li and coworkers reported that SrTiO3 can be driven to a transient ferroelectric phase via intense THz pumping as evidenced by the second harmonic signal of a 800-nm probe pulse 169, although later experiments performed on a cousin compound KTaO3 suggested that the THz-induced transient second-harmonic signal may come from defect-induced local polar structures without any long-range ferroelectric ordering 185.
-A contemporary work in 2019 reported a metastable ferroelectric phase of SrTiO3, which was induced by a mid-IR pump pulse instead of a THz pulse 170.
+In 2019, Li and coworkers reported that SrTiO3 can be driven to a transient ferroelectric phase via intense THz pumping as evidenced by the second harmonic signal of a 800-nm probe pulse 172, although later experiments performed on a cousin compound KTaO3 suggested that the THz-induced transient second-harmonic signal may come from defect-induced local polar structures without any long-range ferroelectric ordering 189.
+A contemporary work in 2019 reported a metastable ferroelectric phase of SrTiO3, which was induced by a mid-IR pump pulse instead of a THz pulse 173.
 In this experiment, Nova and coworkers resonantly drove the high-frequency IR-active mode in SrTiO3 at 4 K, and detected a metastable noncentrosymmetric phase via second-harmonic spectroscopy.
-The experiment was interpreted via anharmonic phonon-phonon interactions between the pulse-induced IR-active mode and acoustic strains, the latter of which suppressed the antiferrodistortive distortion and hence stabilized the ferroelectric phase.
-This mechanism was further clarified by a subsequent work using time-resolved X-ray diffusing scattering 171.
+The results were interpreted via anharmonic phonon-phonon interactions between the pulse-induced IR-active mode and acoustic strain, the latter of which suppressed the antiferrodistortive distortion and hence stabilized the ferroelectric phase.
+This mechanism was further clarified by a subsequent work using time-resolved X-ray diffusing scattering 174.
 The antiferrodistortive fluctuations can lead to a significant diffuse scattering intensity at the point in the Brillouin zone, as shown in the equilibrium X-ray detector image in Fig.
 5d.
-At 135 K, upon photoexcitation with a 17-THz pulse that resonantly drives the IR-active mode, X-ray scattering intensity at shows a sudden increase followed by a decay (Fig.
+At 135 K, upon photoexcitation with a 17-THz pulse that resonantly drives the IR-active mode, the X-ray scattering intensity at shows a sudden increase followed by a decay (Fig.
 5e).
@@
 The same IR-active mode can also couple to low-frequency acoustic modes, which induce a strain that dominates at long time delays ( ps) and suppresses the antiferrodistortive fluctuations, resulting in the decrease of X-ray intensity at the point.
-Though the X-ray experiment was performed at 135 K that is above the antiferrodistortive transition, the authors expected the same mechanism would remain valid to explain the metastable ferroelectric property of SrTiO3 at 4 K, where a suppressed antiferro-distortion at long times leads to the emergence of a ferroelectric phase.
-Aside from the standard pump-probe scheme employed in transient optical spectroscopy and X-ray scattering, 2D THz spectroscopy offers another route to determine the interaction pathway between different quasiparticles, showing particular relevance in revealing the phonon-phonon coupling process 29.
+Though the X-ray experiment was performed at 135 K, which is above the antiferrodistortive transition, the authors expect the same mechanism would remain valid to explain the metastable ferroelectric property of SrTiO3 at 4 K, where a suppressed antiferro-distortion at long times leads to the emergence of a ferroelectric phase.
+Aside from the standard pump-probe scheme employed in transient optical spectroscopy and X-ray scattering, 2D THz spectroscopy offers another route to determine the interaction pathway between different quasiparticles, showing particular relevance in revealing phonon-phonon coupling 29.
 In typical 2D THz spectroscopy setups, two intense THz pump pulses are used to induce nonlinear responses with a time delay between the two pulses, as shown in Fig.
@@
 As a recent example, 2D THz spectroscopy was employed to reveal the energy-flow pathway in a topological antiferromagnet MnBi2Te4 (ref.
-184).
+188).
 The normalized 2D Fourier transform spectrum of the nonlinear signal is shown in Fig.
@@
 IV.3 Phonon-magnon coupling
-Novel properties in quantum materials often arise from the coupling between the spin and lattice degrees of freedom, for example, the Einstein-de Haas effect 190, the Barnett effect 191, 192, magnetostriction 193, and piezomagnetism 194, 195.
-In certain contexts, their strong coupling implies a unique spin configuration dictated by the crystalline lattice, and one prototypical example is the realization of quantum spin liquid predicted in a frustrated geometry such as in a triangular or kagome lattice 196.
-While effects due to strong spin-lattice coupling have been extensively investigated in thermal equilibrium 197, ultrafast techniques allow us to probe the exact energy-flow pathway and unveils how angular momentum is transferred between the electron spin and the crystalline lattice.
-For instance, ferromagnetic systems are known to lose their net magnetization well within a picosecond when they are subjected to an intense ultrafast laser pulse 198, but how the conservation of total angular momentum is maintained at this timescale has been a long-standing puzzle.
-Recent works using time-resolved X-ray and electron diffraction offered some clues to this question 199, 200, where circularly-polarized high-frequency phonons are first excited to absorb the angular momentum, which are then transferred to low-frequency shear acoustic modes and a macroscopic rotation of the sample.
-Interestingly, acoustic shearing is similarly reported during ultrafast demagnetization of a van der Waals antiferromagnet whose net spin angular momentum is zero 201, 21, 202, and a detailed description of the angular momentum flow in the antiferromagnetic case remains an open question.
-Strong spin-lattice couplings can manifest in the form of magnon-phonon coupling, where a hybridized magnon-phonon mode, magnon polarons, can readily be detected in equilibrium.
+Novel properties in quantum materials often arise from the coupling between the spin and lattice degrees of freedom, for example, the Einstein-de Haas effect 194, the Barnett effect 195, 196, magnetostriction 197, and piezomagnetism 198, 199.
+In certain contexts, their strong coupling implies a unique spin configuration dictated by the crystalline lattice, and one prototypical example is the realization of quantum spin liquid predicted in a frustrated geometry such as in a triangular or kagome lattice 200.
+While effects due to strong spin-lattice coupling have been extensively investigated in thermal equilibrium 201, ultrafast techniques allow us to probe the exact energy-flow pathway and reveal how angular momentum is transferred between the electron spin and the crystalline lattice.
+For instance, ferromagnetic systems are known to lose their net magnetization well within a picosecond when they are subjected to an intense ultrafast laser pulse 202, but how the conservation of total angular momentum is maintained at this timescale has been a long-standing puzzle.
+Recent works using time-resolved X-ray and electron diffraction offered some clues to this question 203, 204, where circularly-polarized high-frequency phonons were first excited to absorb the angular momentum, which was then transferred to low-frequency shear acoustic modes and a macroscopic rotation of the sample.
+Interestingly, acoustic shearing has been similarly reported during ultrafast demagnetization of a van der Waals antiferromagnet whose net spin angular momentum is zero 205, 21, 206, and a detailed description of the angular momentum flow in the antiferromagnetic case remains an open question.
+Strong spin-lattice couplings can sometimes manifest in the form of magnon-phonon coupling, where hybridized magnon-phonon modes, magnon polarons, can readily be detected in equilibrium.
 Similar to phonon-polaritons discussed in Sec.
@@
 An example of avoided crossings is shown in Fig.
-6a for a layered antiferromagnet FePSe3, measured by magneto-infrared transmission spectroscopy 186.
+6a for a layered antiferromagnet FePSe3, measured by magneto-infrared transmission spectroscopy 190.
 In the detected spectral window of FePSe3, while the constituent phonons are individually chiral with opposite chiralities, their near degeneracy leads to no net chirality.
-However, the coupling to chiral magnons lifts such degeneracy, resulting in chiral magnon polarons that were demonstrated by magento-Raman spectroscopy.
+However, the coupling to chiral magnons lifts such degeneracy, resulting in chiral magnon polarons that were demonstrated by magneto-Raman spectroscopy.
 The hybridization of magnon and phonon can also be directly measured in the time domain using ultrafast techniques.
 For instance, in a cousin van der Waals antiferromagnet FePS3, two Raman-active optical phonons at 3.2 and 4.8 THz can be excited by a near-infrared pump pulse, evidenced by the coherent oscillations in the polarization rotation signal of the probe pulse (Fig.
-6b) 187.
+6b) 191.
 As the applied magnetic field increases, the 3.2 THz phonon splits into two phonon-magnon modes due to phonon-magnon coupling (Fig.
-6c), which is reminiscent of the hybridization-induced peak splittings detected in equilibrium 203, 204, 205.
+6c), which is reminiscent of the hybridization-induced peak splittings detected in equilibrium 207, 208, 209.
 In the study of phonon-magnon coupling, what sets nonequilibrium experiments apart from their equilibrium counterpart is the coherence of the dynamical motion of both atoms and spins, leading to a number of interesting phenomena.
@@
 In an early demonstration of this effect, Nova and coworkers used a 20-THz mid-infrared pump to resonantly drive the in-plane and phonons in an antiferromagnetic insulator ErFeO3 (ref.
-188), whose phonon eigenvectors are sketched in Fig.
+192), whose phonon eigenvectors are sketched in Fig.
 7a.
@@
 7b).
-The elliptically polarized phonons are otherwise known as chiral phonons, which have attracted both theoretical 207, 208 and experimental 209, 76, 70, 69 interests because of their ability to create an effective magnetic field that paves the way for “chiral phonomagnetism” in magnetization-based devices.
-Coherent magnons can also be excited by phonon-induced lattice distortions that modify the exchange interaction between neighboring atoms, a mechanism recently demonstrated in a closely related antiferromagnet DyFeO3 by Afanasiev and colleagues 189.
-In this material, the highest frequency IR-active transverse optical phonon mode () is associated with a periodic stretching of the Fe-O bond.
+The elliptically polarized phonons are otherwise known as chiral phonons, which have attracted both theoretical 211, 212 and experimental 213, 76, 70, 69 interests because of their ability to create an effective magnetic field that paves the way for “chiral phonomagnetism” in magnetization-based devices.
+Coherent magnons can also be excited by phonon-induced lattice distortions that modify the exchange interaction between neighboring atoms, a mechanism recently demonstrated in a closely related antiferromagnet DyFeO3 by Afanasiev and colleagues 193.
+In this material, the highest-frequency IR-active transverse optical phonon mode () is associated with a periodic stretching of the Fe-O bond.
 Using a tailored terahertz pump pulse, this phonon can be resonantly driven to a large amplitude and couple to a finite lattice distortion along the coordinate of a Raman-active mode via the nonlinear phononics pathway discussed in Sec.
@@
 4).
-This lattice distortion involves an antipolar motion of the Dy3+ ions, resulting in a rectification of the Fe-Dy exchange integral that can reach approximately 1–2 eV per unit cell according to density functional theory calculations, thus setting off a coherent magnon oscillation.
+This lattice distortion involves an antipolar motion of the Dy3+ ions, resulting in a rectification of the Fe-Dy exchange integral that can reach approximately 1–2 eV per unit cell according to DFT calculations, thus setting off coherent magnon oscillations.
 Depending on whether DyFeO3 is in the antiferromagnetic state ( K) or the weakly ferromagnetic state ( K), distinct magnon modes are induced by the resonant driving of the same optical phonon (Fig.
 7c,d).
-Importantly, the magnon mode frequencies under the resonant driving condition (orange peak in the insets) are clearly distinct from those under off-resonant pumping (black peak in the insets), the latter of which are excited via impulsive stimulated Raman scattering 210.
+Importantly, the magnon mode frequencies under the resonant driving condition (orange peak in the insets) are clearly distinct from those under off-resonant pumping (black peak in the insets), the latter of which are excited via impulsive stimulated Raman scattering 214.
 Leveraging the modification of exchange interaction under a resonant pumping condition, Afanasiev and colleagues further showed that with a very large fluence of the resonant pump, an ultrafast magnetic transition from an antiferromagnetic state to a weakly ferromagnetic state can be induced coherently, a transition that proceeded much faster compared to an above-bandgap excitation scheme.
@@
 used an intense, nearly single-cycle terahertz pulse to pump the crystal below its Néel temperature, where the magnetic field of the pulse resonantly excited a magnon mode (Fig.
-8a) 206.
-After approximately 10 ps, the amplitude of the magnon decreases, followed by a persistent phonon oscillation over the next tens of picoseconds.
+8a) 210.
+After approximately 10 ps, the amplitude of the magnon decreased, followed by a persistent phonon oscillation over the next tens of picoseconds.
 By contrast, the magnon and phonon oscillations both disappeared above the Néel temperature (Fig.
 8b), strongly suggesting that the phonons are induced by the magnons and not vice versa.
-The researchers further performed 2D terahertz spectroscopy to demonstrate the causal relation that the nonlinear excitation of phonons was mediated by a coherent magnon, as shown by the Fourier spectrum of the nonlinear amplitude in Fig.
+The researchers further performed 2D terahertz spectroscopy to demonstrate the causal relation that a coherent magnon mediated the nonlinear excitation of phonons, as shown by the Fourier spectrum of the nonlinear amplitude in Fig.
 8c.
@@
 The magnon-mediated excitation pathway for the coherent phonon is summarized in Fig.
-8e.
+8d.
 In the first step, an intense terahertz field resonantly populates the coherent magnonic state at frequency , creating an intermediate state.
@@
 When a coherent phonon mode is driven out of equilibrium following an optical pump, the resulting atomic motions can in turn modulate the optical properties in a coherent manner.
-In materials whose optical response is strongly modified by the presence of excitons — or bound electron-hole pairs — the exciton binding energy 211, 212, 213, lifetime 214, 215, and transport behavior 216 can also be changed.
-Such phonon-exciton coupling has been widely studied in quantum dots and quantum wells 217, 218, 219, and more recently, in complex bulk and atomically-thin materials 211, 212, 213.
+In materials whose optical response is strongly modified by the presence of excitons — or bound electron-hole pairs — the exciton binding energy 215, 216, 217, lifetime 108, 218, and transport behavior 219 can also be changed.
+Such phonon-exciton coupling has been widely studied in quantum dots and quantum wells 220, 221, 222, and more recently, in complex bulk and atomically-thin materials 215, 216, 217.
 By comparing the exciton lineshape in broadband absorption or photoluminescence spectra before and after the photoexcitation event, one can observe how excitons respond to phonons in the time domain, yielding insights for tuning exciton properties via ultrafast phonon engineering.
 For example, in anatase TiO2, a bulk semiconductor, Baldini et al.
-used photoinduced coherent acoustic phonons to control its excitons at room temperature, observing an exciton energy shift up to 30 meV, which is one of the largest reported in solids under a time-dependent perturbation 211.
-In a prototypical two-dimensional lead iodide perovskite, (PEA)2PbI4 (PEA = phenylethylammonium), it was further discovered that different excitons have varying coupling strengths with different coherent phonons, suggesting distinct polaronic characters of each exciton in this system and possibly in a broader class of hybrid semiconductors 212.
+used photoinduced coherent acoustic phonons to control its excitons at room temperature, observing an exciton energy shift up to 30 meV, which is one of the largest reported in solids under a time-dependent perturbation 215.
+In a prototypical two-dimensional lead iodide perovskite, (PEA)2PbI4 (PEA = phenylethylammonium), it was further discovered that different excitons have varying coupling strengths with different coherent phonons, suggesting distinct polaronic characters of each exciton in this system and possibly in a broader class of hybrid semiconductors 216.
 In the atomically thin limit, Trovatello et al.
-also showed a pronounced phonon-exciton coupling effect in a monolayer MoS2, whose excitons are strongly affected by a photoinduced coherent phonon 213.
+also showed a pronounced phonon-exciton coupling effect in a monolayer MoS2, whose excitons are strongly affected by a photoinduced coherent phonon 217.
 Using transient absorption spectroscopy, the authors studied the interaction between out-of-plane phonon modes and excitons, noting a marked enhancement in the phonon oscillation amplitude at resonant exciton frequencies.
 The aforementioned studies of phonon-exciton coupling, however, lack spatial information, which is important for learning the transport behavior of excitons under the influence of phonons.
-Recently, Peng and coworkers demonstrated microscopic imaging of directional transport of interlayer excitons in bilayer WSe2 with the assistance of surface acoustic waves (ref.
-216), and their experimental scheme is summarized in Fig.
+Recently, Peng and coworkers demonstrated microscopic imaging of directional transport of interlayer excitons in bilayer WSe2 with the assistance of surface acoustic waves 219, and their experimental scheme is summarized in Fig.
 9a.
@@
 Despite the lack of temporal resolution in this measurement, one can imagine a spatiotemporal experimental scheme by using a pulsed laser to launch the excitons.
-Crucially, this experiment introduces a new way for controlling exciton transport via electrically-excited or photoinduced acoustic waves, a versatile approach that can be adapted to many other 2D semiconductors and heterostructures.
+Crucially, this experiment introduces a new way to control exciton transport via electrically-excited or photoinduced acoustic waves, a versatile approach that can be adapted to many other 2D semiconductors and heterostructures.
 These studies underscore the diversity and tunability of phonon-exciton interactions across material systems, providing insights into phononic control over excitonic properties, which has far-reaching implications for the next-generation design of optoelectronic devices.
 IV.5 Other collective excitation couplings
-Despite the special focus on phonons as one of the coupled modes in the preceding sections, the full combinatorics of collective excitation coupling has been applied to many other scenarios such as magnon-magnon 222, 221, 220, 223, exciton-magnon 224, plasmon-polariton 225, 144 among other interactions.
+Despite the special focus on phonons as one of the coupled modes in the preceding sections, the full combinatorics of collective excitation coupling has been applied to many other scenarios such as magnon-magnon 225, 224, 223, 226, 227, exciton-magnon 228, plasmon-polariton 229, 148 among other interactions.
 In this section, we discuss a few examples that showcase the diversity of collective excitation couplings and their role in determining the ultrafast dynamics of quantum materials.
-As an analogy of nonlinear phononics that opens a new pathway for coherent manipulation of the crystalline structure, magnon-magnon coupling provides an avenue for controlling the spin dynamics.
+Similar to nonlinear phononics, which opens a new pathway for the coherent manipulation of the crystalline structure, magnon-magnon coupling provides an avenue for controlling the spin dynamics.
 Utilizing 2D THz spectroscopy, Blank and coworkers revealed the down-conversion coupling of two magnon modes in the canted antiferromagnet FeBO3 (ref.
-222).
-In another canted antiferromagnet, ErFeO3, Zhang and coworkers also demonstrated a unidirectional up-conversion coupling between distinct magnon modes (ref.
-221).
+225).
+In another canted antiferromagnet, ErFeO3, Zhang and coworkers also demonstrated a unidirectional up-conversion coupling between distinct magnon modes 224.
 The canted spins in these materials result from the competition between antiferromagnetic interaction and Dzyaloshinskii-Moriya interaction, yielding a small net magnetic moment .
@@
 This observation is expected as qFM involves a precession in the - plane while qAFM involves an amplitude oscillation along .
-Surprisingly, when , qAFM mode was still observed with (lower panel in Fig.
-10b), which was not expected if there were only linear interactions between the driving THz magnetic field and magnons.
+Surprisingly, when , qAFM mode is still observed with (lower panel in Fig.
+10b), which is not expected if there were only linear interactions between the driving THz magnetic field and magnons.
 To elicit the nonlinear interactions, 2D THz spectroscopy measurements were employed, where the nonlinear spectrum reveals a unidirectional up-conversion process from qFM to qAFM.
@@
 5c, ref.
-163), the sum and difference frequencies between these two magnon modes have been demonstrated in a cousin rare-earth ferrite, YFeO3, which are featured by a plethora of off-diagonal peaks in the nonlinear 2D terahertz spectrum in Fig.
+187), the sum and difference frequencies between these two magnon modes have been demonstrated in a cousin rare-earth ferrite, YFeO3, which are featured by a plethora of off-diagonal peaks in the nonlinear 2D terahertz spectrum in Fig.
 10d (ref.
-220).
+223).
 Like nonlinear phononics, these nonlinear effects of magnons demonstrated in the time domain help usher in new ways to coherently manipulate the spin degree of freedom in magnetic materials.
@@
 When coupled with other collective excitations, magnon propagation can be modified.
-For instance, for magnons that are coupled with acoustic phonons, theoretical 226 and experimental 227, 228 works have shown that magnon polarons travel over long distances with a faster group velocity than magnons alone.
-In another study of a van der Waals antiferromagnetic semiconductor CrSBr, magnons were observed to coherently travel beyond several micrometers with a very fast group velocity, an effect that Bae and coworkers interpreted as the coupling between the magnon and acoustic phonon branches 224, 229.
-What is interesting about this experiment is that coherent magnons were probed by laser pulses tuned to the exciton resonances in CrSBr, demonstrating an efficient optical access to spintronic information due to strong exciton-magnon couplings in this material.
-A subsequent experiment suggested that optically launched magnon propagation in CrSBr and likely other van der Waals magnet is instead mediated by long-range magnetic-dipole interactions 230, calling for additional spatiotemporally-resolved studies to understand magnon dynamics in this important class of materials.
+For instance, for magnons that are coupled with acoustic phonons, theoretical 230 and experimental 231, 232 works have shown that magnon polarons travel over long distances with a faster group velocity than magnons alone.
+In a study of a van der Waals antiferromagnetic semiconductor CrSBr, magnons were observed to coherently travel beyond several micrometers with a very fast group velocity, an effect that Bae and coworkers interpreted as the coupling between the magnon and acoustic phonon branches 228, 233.
+What is interesting about this experiment is that coherent magnons were probed by laser pulses tuned to the exciton resonances in CrSBr, demonstrating efficient optical access to spintronic information due to strong exciton-magnon couplings in this material.
+A subsequent experiment suggested that optically launched magnon propagation in CrSBr and likely other van der Waals magnets are instead mediated by long-range magnetic-dipole interactions 234, calling for additional spatiotemporally-resolved studies to understand magnon dynamics in this important class of materials.
 Non-local propagation effects mediated by the coupling between collective excitations are not restricted to magnons.
-Other examples include exciton propagation with the aid of surface acoustic waves in a GaAs quantum well 231, 232 or in a transition metal dichalcogenide 216, and surface-plasmon-polariton propagation at a metal-dielectric interface 233.
+Other examples include exciton propagation with the aid of surface acoustic waves in a GaAs quantum well 235, 236 or in a transition metal dichalcogenide 219, and surface-plasmon-polariton propagation at a metal-dielectric interface 237.
 Similar to a phonon-polariton, a plasmon-polariton is a hybridized excitation between charge oscillations and light.
@@
 As a result, a hybridized wave, namely surface-plasmon-polariton that involves both the charge motion on the metal surface and electromagnetic field in the dielectric, propagates along the interface.
-To observe and control its propagation dynamics, PINEM is one of the leading techniques 146, 144.
-Figure 11a shows the spatiotemporally-resolved imaging of surface-plasmon-polaritons induced by near-infrared laser pulses on an isolated silver nanowire suspended on a graphene film 144, where surface-plasmon-polaritons are evidenced by the additional peaks that appear in the electron energy loss spectra when the light and electron pulses are temporally overlapped.
+To observe and control its propagation dynamics, PINEM is one of the leading techniques 150, 148.
+Figure 11a shows the spatiotemporally-resolved imaging of surface-plasmon-polaritons induced by near-infrared laser pulses on an isolated silver nanowire suspended on a graphene film 148, where surface-plasmon-polaritons are evidenced by the additional peaks that appear in the electron energy loss spectra when the light and electron pulses are temporally overlapped.
 These peaks result from the quantized gain and loss of energy when electrons interact with the surface-plasmon-polariton field.
-Using energy-filtered electrons, the temporal dynamics of the surface-plasmon-polariton field was imaged with high spatial resolution as it forms an evolving interferometric standing wave pattern (Fig.
+Using energy-filtered electrons, the temporal dynamics of the surface-plasmon-polariton field was imaged with high spatial resolution as it formed an evolving interferometric standing wave pattern (Fig.
 11b–f).
-Moreover, with different light polarizations, different surface-plasmon-polariton field distributions were obtained, demonstrating the potential to control surface-plasmon-polariton using light.
+Moreover, with different light polarizations, different surface-plasmon-polariton field distributions were obtained, demonstrating the potential to control surface-plasmon-polaritons using light.
 V Summary and outlook
-The study and manipulation of coupled collective excitations by ultrafast light pulses offer a new perspective to disentangle intertwined degrees of freedom that are difficult to isolate in equilibrium.
-In this review, we have highlighted several couplings that both yield an in-depth understanding of the equilibrium properties and help one to coherently control the nonequilibrium dynamics in the host material.
+The study and manipulation of coupled collective excitations by ultrafast light pulses offer a new perspective on disentangling intertwined degrees of freedom that are difficult to isolate in equilibrium.
+In this review, we have highlighted several couplings that both yield an in-depth understanding of the equilibrium properties and enable the coherent control of the nonequilibrium dynamics in the host material.
 While examples mentioned so far only include two participating collective excitations, multipartite couplings have been theoretically pondered, such as a hybrid exciton-phonon-photon that gives rise to a “phonoriton” in an optical cavity (Fig.
-12a) 236, 234, and experimentally realized, such as the hybrid plasmon-exciton-photon — known as “plexcitons” — in WS2 (Fig.
-12b), where cavity effects were achieved through nano-patterning (refs.
-235, 237).
-These studies fall into a larger body of works that aims to manipulate quantum materials via cavity control, in which the light-matter coupling strength can be greatly enhanced by replacing the classical laser fields with quantum-mechanical photon modes.
-One of the most celebrated examples is microcavity exciton–polaritons in GaAs quantum wells, which arise from the strong coupling between photons and excitons and exhibit remarkable effects such as Bose-Einstein condensation 238, 239, 240.
-More recently, a multitude of theoretical proposals have been put forward for tailoring the unique properties of quantum materials via cavity design 241, 242, with experimental progress in the study of topological state 243, superconductivity 244, and phase transition 245 in different material platforms.
+12a) 240, 238, and experimentally realized, such as the hybrid plasmon-exciton-photon — known as a “plexciton” — in WS2 (Fig.
+12b), where cavity effects were achieved through nano-patterning 239, 241.
+These studies fall into a larger body of works that aims to manipulate quantum materials via cavity control, in which the light-matter coupling strength can be greatly enhanced by replacing classical laser fields with quantum-mechanical photon modes.
+One of the most celebrated examples is microcavity exciton–polaritons in GaAs quantum wells, which arise from the strong coupling between photons and excitons and exhibit remarkable effects such as Bose-Einstein condensation 242, 243, 244.
+More recently, a multitude of theoretical proposals have been put forward for tailoring the unique properties of quantum materials via cavity design 245, 246, with experimental progress in the study of topological states 247, superconductivity 248, and phase transitions 249 in different material platforms.
 Besides the potential in the experimental realization of such multi-way couplings in the context of cavity quantum materials, we comment on a few other directions in the study of coupled excitations from the viewpoint of ultrafast science.
-First, advances in engineering the waveform of the excitation light pulse — such as its polarization, carrier envelope phase, and orbital angular momentum — can greatly expand the tool kits for resonant excitation of collective modes and hence a deterministic control over material properties.
+First, advances in engineering the waveform of excitation light pulses — such as their polarization, carrier envelope phase, and orbital angular momentum — can greatly expand the toolkits for resonant excitation of collective modes and hence enable a deterministic control over material properties.
 A recent example is the synthesis of an intense, circularly-polarized THz pulse used to induce circular motions of atoms, otherwise known as chiral phonons, which can in turn induce a large effective magnetic field 76, 71, 70, 69.
 In particular, Davies et al.
-demonstrated that the induced field can significantly alter the magnetic order even at considerable distances from the phonon creation point 70, underscoring the need for a deeper insight into chiral phonon transport.
+demonstrated that the induced field can significantly alter the magnetic order even at considerable distances from the phonon creation point 70, underscoring the need for deeper insight into chiral phonon transport.
 Equipped with the capability to deterministically excite particular phonon and magnon modes using the mechanisms introduced in this review, one can envision future experiments where chiral phonons and magnons are created at distinct sample locations, and their mutual interaction over space and time can be used to gain a mechanistic understanding of the propagation mechanism for realizing ultrafast non-local control of magnetic ordering.
@@
 An alternative route is to look for complementary probes with a similar probing depth, such as high-energy electron diffraction and extreme ultraviolet absorption spectroscopy, both operated under a transmission geometry.
-While the former is sensitive to the lattice degree of freedom, the latter excels in capturing carrier, magnetic, and excitonic excitations with an exceptional time resolution 25, 246.
+While the former is sensitive to the lattice degree of freedom, the latter excels in capturing carrier, magnetic, and excitonic excitations with exceptional time resolution 25, 250.
 One can envision a chamber design that integrates both probes in a single setup, making it possible to concurrently access lattice and electronic information following the same excitation pulse.
-Third, while the review has largely focused on coherent collective dynamics with a well-defined phase, not all light-induced collective excitations are coherent or fully reproducible over multiple pump pulses, particularly in systems with spatial inhomogeneities where defects and disorders are prevalent 247, 248, 249, 250.
-In this aspect, techniques with both spatial and temporal resolutions that operate in the single-shot mode are important for revealing the local, incoherent, and stochastic dynamics that vary from pulse to pulse, as recently demonstrated in the study of magnetic fluctuations in magnetic multilayer thin films 138, 139 and the heterogeneous insulator-to-metal transition in VO2 (ref.
-140), both of which utilized a coherent X-ray source in large-scale user facilities.
-It is worth noting that stochastic dynamics can also be accessed without relying on single-shot detection 251, which makes it possible to study fluctuating systems with the stroboscopic scheme employed in the vast majority of ultrafast measurements.
+Third, while the review has largely focused on coherent collective dynamics with a well-defined phase, not all light-induced collective excitations are coherent or fully reproducible over multiple pump pulses, particularly in systems with spatial inhomogeneities where defects and disorders are prevalent 251, 252, 253, 254.
+In this aspect, techniques with both spatial and temporal resolutions that operate in the single-shot mode are important for revealing the local, incoherent, and stochastic dynamics that vary from pulse to pulse, as recently demonstrated in the study of magnetic fluctuations in magnetic multilayer thin films 142, 143 and the heterogeneous insulator-to-metal transition in VO2 (ref.
+144), both of which utilized a coherent X-ray source in large-scale user facilities.
+It is worth noting that stochastic dynamics can also be accessed without relying on single-shot detection 255, which makes it possible to study fluctuating systems with the stroboscopic scheme employed in the vast majority of ultrafast measurements.
 Along with rapid progress in facility-based ultrashort coherent X-ray sources, we also envision two key developments in table-top experiments for probing local and incoherent collective modes.
-In the extreme spatial scale, innovations in scanning tunneling microscopy has enabled new ways to image collective excitations at the atomic scale 141, 142, paving the way for controlling these local dynamics through creative engineering of nanoscopic defects.
-In the extreme temporal scale, ultrafast electron diffraction and microscopy that have witnessed a rapid improvement of their temporal resolution 252, 109, 253, 254 would provide a useful platform to study the initial emergence of incoherent phonons following photoexcitation 255, 256, which play an important role in determining the ensuing defect structure 53 and energy flow in the host material system 257.
-As we conclude this review that has touched a diverse set of experimental techniques and observables, it is an apt place to be reminded that the overarching goal of studying coupled collective excitations is to uncover a set of unifying principles that govern the interaction among the different degrees of freedom in solids.
+At the extreme spatial scale, innovations in scanning tunneling microscopy have enabled new ways to image collective excitations at the atomic scale 145, 146, paving the way for controlling these local dynamics through creative engineering of nanoscopic defects.
+At the extreme temporal scale, ultrafast electron diffraction and microscopy, which have witnessed a rapid improvement of their temporal resolution 256, 109, 257, 258, 259, would provide a useful platform to study the initial emergence of incoherent phonons following photoexcitation 260, 261, which play an important role in determining the ensuing defect structure 53 and energy flow in the host material system 262.
+As we conclude this review that has touched a diverse set of experimental techniques and observables, it is apt to be reminded that the overarching goal of studying coupled collective excitations is to uncover a set of unifying principles that govern the interaction among the different degrees of freedom in solids.
 While studying collective excitations is certainly not the only way to unravel the intricate microscopic interactions, they offer us a way to describe a many-body system without the need to keep track of individual dynamics of an astronomical number of particles, making it tractable to study and engineer material properties both in and out of equilibrium.
 Acknowledgements.
-We thank helpful input from Edoardo Baldini, Henry Bell, Thomas Blank, Yu-Che Chien, Shaofeng Duan, Nuh Gedik, Patrick Liu, Tianchuang Luo, Mariano Trigo, and Xunqing Yin during the preparation of this manuscript.References
+We thank helpful input from Edoardo Baldini, Henry Bell, Thomas Blank, Yu-Che Chien, Shaofeng Duan, Nuh Gedik, Patrick Liu, Megan Loh, Tianchuang Luo, Enrico Perfetto, Ruixin Qiu, Gianluca Stefanucci, Mariano Trigo, and Xunqing Yin during the preparation of this manuscript.References
 - Lee et al. [2006] P. A. Lee, N. Nagaosa, and X.-G. Wen, Doping a Mott insulator: Physics of high-temperature superconductivity, Rev. Mod. Phys. 78, 17 (2006).
@@
 Rubio, and N.
-Gedik, Terahertz field-induced metastable magnetization near criticality in FePS3, Nature 636, 609 (2024).
+Gedik, Terahertz field-induced metastable magnetization near criticality in FePS3, Nature 636, 609–614 (2024).
 - Yoshikawa et al.
@@
 - Robert et al.
-[2016a] C.
+[2016] C.
 Robert, D.
@@
 Rev.
-B 93, 205423 (2016a).
+B 93, 205423 (2016).
 - Qi et al.
@@
 Demler, and N.
-Gedik, Role of equilibrium fluctuations in light-induced order, Physical Review Letters 127, 227401 (2021).
+Gedik, Role of equilibrium fluctuations in light-induced order, Phys.
+Rev.
+Lett.
+127, 227401 (2021).
 - Pan and Caruso [2023] Y. Pan and F. Caruso, Vibrational dichroism of chiral valley phonons, Nano Lett. 23, 7463 (2023).
@@
 Dani, Structure of the moiré exciton captured by imaging its electron and hole, Nature 603, 247–252 (2022).
+- Perfetto et al.
+[2016] E.
+Perfetto, D.
+Sangalli, A.
+Marini, and G.
+Stefanucci, First-principles approach to excitons in time-resolved and angle-resolved photoemission spectra, Phys.
+Rev.
+B 94, 245303 (2016).
+- Steinhoff et al.
+[2017] A.
+Steinhoff, M.
+Florian, M.
+Rösner, G.
+Schönhoff, T.
+O.
+Wehling, and F.
+Jahnke, Exciton fission in monolayer transition metal dichalcogenide semiconductors, Nat.
+Commun.
+8, 1166 (2017).
 - Rustagi and Kemper [2018] A. Rustagi and A. F. Kemper, Photoemission signature of excitons, Phys. Rev. B 97, 235310 (2018).
+- Rustagi and Kemper [2019] A. Rustagi and A. F. Kemper, Coherent excitonic quantum beats in time-resolved photoemission measurements, Phys. Rev. B 99, 125303 (2019).
+- Christiansen et al.
+[2019] D.
+Christiansen, M.
+Selig, E.
+Malic, R.
+Ernstorfer, and A.
+Knorr, Theory of exciton dynamics in time-resolved ARPES: Intra- and intervalley scattering in two-dimensional semiconductors, Phys.
+Rev.
+B 100, 205401 (2019).
 - Zewail and Thomas [2009] A. H. Zewail and J. M. Thomas, 4D electron microscopy: imaging in space and time (World Scientific, 2009).
@@
 Phys.
-10.1038/s41567-024-02552-7 (2024).
+20, 1603–1608 (2024).
 - Barwick and Zewail [2015] B. Barwick and A. H. Zewail, Photonics and plasmonics in 4D ultrafast electron microscopy, ACS Photonics 2, 1391 (2015).
@@
 - Burkel [2000] E. Burkel, Phonon spectroscopy by inelastic x-ray scattering, Rep. Prog. Phys. 63, 171 (2000).
-- Kozina et al.
-[2019] M.
-Kozina, M.
-Fechner, P.
-Marsik, T.
-van Driel, J.
-M.
-Glownia, C.
-Bernhard, M.
-Radovic, D.
-Zhu, S.
-Bonetti, U.
-Staub, and M.
-C.
-Hoffmann, Terahertz-driven phonon upconversion in SrTiO3, Nat.
-Phys.
-15, 387 (2019).
 - Först et al. [2011] M. Först, C. Manzoni, S. Kaiser, Y. Tomioka, Y. Tokura, R. Merlin, and A. Cavalleri, Nonlinear phononics as an ultrafast route to lattice control, Nat. Phys. 7, 854 (2011).
@@
 Cavalleri, Comment on arXiv:2210.01114: Optical Saturation Produces Spurious Evidence for Photoinduced Superconductivity in K3C60 (2023), arXiv:2303.10169 .
+- Kozina et al.
+[2019] M.
+Kozina, M.
+Fechner, P.
+Marsik, T.
+van Driel, J.
+M.
+Glownia, C.
+Bernhard, M.
+Radovic, D.
+Zhu, S.
+Bonetti, U.
+Staub, and M.
+C.
+Hoffmann, Terahertz-driven phonon upconversion in SrTiO3, Nat.
+Phys.
+15, 387 (2019).
 - Blank et al.
@@
 Dal Conte, Strongly coupled coherent phonons in single-layer MoS2, ACS nano 14, 5700 (2020a).
-- Robert et al.
-[2016b] C.
-Robert, D.
-Lagarde, F.
-Cadiz, G.
-Wang, B.
-Lassagne, T.
-Amand, A.
-Balocchi, P.
-Renucci, S.
-Tongay, B.
-Urbaszek, and X.
-Marie, Exciton radiative lifetime in transition metal dichalcogenide monolayers, Phys.
-Rev.
-B 93, 205423 (2016b).
 - Trovatello et al.
@@
 V.
-Kimel, Empowering control of antiferromagnets by THz-induced spin coherence, Physical Review Letters 131, 096701 (2023).
+Kimel, Empowering control of antiferromagnets by THz-induced spin coherence, Phys.
+Rev.
+Lett.
+131, 096701 (2023).
 - Leenders et al. [2024] R. A. Leenders, D. Afanasiev, A. V. Kimel, and R. V. Mikhaylovskiy, Canted spin order as a platform for ultrafast conversion of magnons, Nature 630, 335–339 (2024).
+- Zhang et al.
+[2024d] Z.
+Zhang, Y.-C.
+Chien, M.
+Tou Wong, F.
+Y.
+Gao, Z.-J.
+Liu, X.
+Ma, S.
+Cao, E.
+Baldini, and K.
+A.
+Nelson, Terahertz stimulated parametric downconversion of a magnon mode in an antiferromagnet (2024d), arXiv:2412.01989 .
 - Bae et al.
@@
 - Hui et al. [2024] D. Hui, H. Alqattan, M. Sennary, N. V. Golubev, and M. T. Hassan, Attosecond electron microscopy and diffraction, Sci. Adv. 10, eadp5805 (2024).
+- Baum and Ropers [2024] P. Baum and C. Ropers, Comment on ’Attosecond electron microscopy and diffraction’ (2024), arXiv:2411.14518 .
 - René de Cotret et al.